### v1

# Robust Survival Analysis with Adversarial Regularization

Abstract
Survival Analysis (SA) is about modeling the time for an event of interest to occur, which has important applications in many fields, including medicine, defense, finance, and aerospace. Recent work has demonstrated the benefits of using Neural Networks (NNs) to capture complicated relationships in SA. However, the datasets used to train these models are often subject to uncertainty (e.g., noisy measurements, human error), which we show can substantially degrade the performance of existing techniques. To address this issue, this work leverages recent advances in NN verification to provide new algorithms for generating fully-parametric survival models that are robust to such uncertainties. In particular, we introduce a robust loss function for training the models and use CROWN-IBP regularization to address the computational challenges with solving the resulting Min-Max problem. To evaluate the proposed approach, we apply relevant perturbations to publicly available datasets in the SurvSet repository and compare survival models against several baselines. We empirically show that Survival Analysis with Adversarial Regularization (SAWAR) method on average ranks best for dataset perturbations of varying magnitudes on metrics such as Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI), concluding that adversarial regularization enhances performance in SA.
Code: https://github.com/mlpotter/SAWAR
Index Terms:
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
Survival Analysis (SA) is a highly active research field [1, 2] used to model the time before an event happens. For example, SA has been used extensively in medicine to quantify the effectiveness of various treatments as part of studies for cancer patients and drug trials [3, 4, 5]. While early approaches (dating back to the 1950s) predominantly used linear parametric models [6, 7], recent methods often leverage Neural Networks (NNs) to model complicated time-to-event data [8].
A key open challenge in this field is in developing models that are robust to uncertainty and variability in the datasets. Successful generalization of machine learning models over unseen datasets is crucial for clinical applications, as efficacy of developed drugs or treatments should be demonstrated over diverse patient demographics, institutions, and data acquisition settings [9, 10, 11]. Data uncertainty could be caused by human errors (e.g., notoriously poor handwriting of physicians [12], ineffective medical record keeping [13]), or inherent noise during transmissions or data collection [14]. However, many approaches assume perfect measurement of the covariates, ignoring aleatoric uncertainty [15]. In SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event [16, 6]. Moreover, several studies have shown that adversaries can influence the final model to be over-/under-confident by purposefully poisoning studies or otherwise disrupting the data collection activities [17, 18]. Despite these known failure modes, and coupled with the sensitivity of NNs to input uncertainties, addressing the robustness of NNs against $\epsilon$-perturbations trained specifically for fully-parametric SA remains unexplored.
To address these issues, this paper’s contributions include:
  • identification of the sensitivity of existing SA methods, such as Deep Regularized Accelerated Failure Time (DRAFT) [19], to covariate perturbations via evaluating open-source datasets with a wide range of metrics related to predictive accuracy, calibration, and population curve estimation.
  • our approach, Survival Analysis with Adversarial Regularization (SAWAR), which is the first method for enabling adversarial robustness to covariate perturbations within a bounded uncertainty set for fully-parametric NN-based SA. We improve calibration and stability on unexpected or noisy data by introducing a loss function based on a Min-Max formulation and Convex Relaxation Based Backward Bounding Pass with Interval Bound Propagation (CROWN-IBP) [20, 21].
The paper is structured as follows: Section II describes the State of the Art; Section III provides background to SA terminology and dataset structure; Section IV provides background on the adversarial and perturbation attacks used; Section V discusses the custom loss employed to deal with those attacks; Section VI provides the results of the experimental activity; finally, Section VII derives conclusions on our findings.
II Related Work
This section focuses on the literature about NNs for SA, with further detail about NN implementations available in a recent survey paper [22]. The first use of NNs in this field was a Feedforward Neural Network (FFNN) to approximate the survival function [23], defined as the probability that no event occurs up to a specific time $t$. More recently, [24] proposed Cox-nnet, a NN that outputs a proportional-hazard [25, 6], which is the assumption that the difference in effect between different instances is constant as a function of time. Building on this idea, [26] developed DeepSurv, a Cox proportional-hazard NN for personalized treatment, which uses dropout layers to improve regularization and reduce overfitting. Instead of proportional hazard estimation, DeepWeiSurv [27] regresses the parameters of a Weibull mixture distribution with NNs.
Several recent works in SA aim to improve model calibration (e.g., by using new terms in the objective function [28, 29, 30]). A calibrated model provides a predicted confidence that is indicative of the true likelihood that the model is correct [31]. Meanwhile, few works have addressed adversarial robustness in SA. In general, a model is adversarially robust if small perturbations to the input do not cause the model to make incorrect predictions [32]. Adversarial robustness and model calibration are loosely coupled: while [33] showed that better calibration does not necessarily improve generalization, [34] showed that incorporating adversarial robustness during training improves trustworthiness in calibration and stability on unexpected data.
Existing adversarial training methods for SA face important tradeoffs between model complexity, performance, and ease of embedding of physical assumptions. Most of these techniques use Conditional Generative Adversarial Networks (CGANs) to generate samples of the time-to-event from a survival distribution in a non-parametric manner [19, 35, 36]. These models (e.g., CGANs) do not assume a functional form and define the survival distribution implicitly [37, 38, 39]. Therefore, there is no mathematical expression that describes the time-to-event Probability Density Function (PDF) value given covariates. One benefit of implicit models is that increased model complexity leads to increased flexibility, which is the ability to fit many other potential data realizations [40]. However, the higher model complexity generally violates the Bayesian Occam’s Razor, which penalizes models for having too much flexibility [40]. To satisfy the Bayesian Occam’s Razor’s “just right“ model complexity, modeling assumptions known as inductive biases are employed in the form of an explicit parametric model [39, 37]. For example, a common inductive bias for SA models is the memoryless property [38], which the exponential distribution satisfies. For this reason, this paper uses the exponential distribution, which is commonly used in SA [41, 42, 43] (and more generally in queuing theory [44, 45], reliability analysis [46, 8], and telecommunications [45]).
III Preliminaries on Survival Analysis
SA [25, 6] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [47]. Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded. Upon entry time to a study, the instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured. In summary, the dataset $\mathcal{D}$ has the following structure:
$\mathcal{D}=\{(\boldsymbol{x_{i}},t_{i},e_{i})\}_{i=1}^{N}.$
Being a statistical approach, we define the required probabilistic functions to model time-to-event data next.
III-A Proportional Hazard Model
Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with PDF $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$. The survival function $S(t):=1-F(t)$ is the complement of this CDF [48, 6].
An exponential parametric survival model is expressed as
$S(t)=e^{-\Lambda(t)},$ (1)
[…]
$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta}$ (2)
is the instantaneous failure rate at time $t$.
To incorporate the instance’s covariates, we use the proportional hazard model [49, 7], This model assumes that the covariates increase or decrease the hazard by a proportional amount at all times [48, 6],
$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=\lambda_{0}(t)e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})},$ (3)
where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative risk associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
The exponential proportional hazard model characterizes the baseline hazard rate as a constant function of time, and therefore $\lambda_{0}(t)=\lambda_{0}$. For simplicity, we assume an exponential baseline hazard, and we may “absorb” the $\lambda$ term into the bias of the NN:
$\lambda_{0}e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (4)
The proportional hazard model PDF, CDF, and complement CDF for an exponential proportional hazard model are then given by:
$f_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
$F_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}.$
$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
The population survival curve is the marginalization of the instance survival function over the covariates,
[…]
Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown [50], the integral in Eq. 5 is often replaced with a Monte Carlo estimate [28]:
$S(t)$ $\approx\frac{1}{N}\sum_{i=1}^{N}S_{\boldsymbol{\theta}}(t|\boldsymbol{x_{i}}).$ (6)
III-B Baseline Objective Functions
Given those definitions, the following loss functions enable finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$. In particular, we follow [51, 52, 19] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the proportional hazard model [6]:
$\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})=$ $\sum_{i=1}^{N}e_{i}\cdot\log f_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}})\ +$
$\sum_{i=1}^{N}(1-e_{i})\cdot\log S_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}}).$ (7)
The first summation in Eq. 7 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time of occurrence is observed. The second summation in Eq. 7 is the log likelihood when right censoring occurs.
III-B2 Ranking Objective
The $\mathcal{L}_{rank}$ objective is a ranking loss that penalizes an incorrect ordering of pairs $\{F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\}$, where instance $i$ with event time $t_{i}<t_{j}$ should have a higher probability of failure than instance $j$ at time $t_{i}$:
$\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ $=\sum_{i\neq j}A_{i,j}\eta\bigg{(}F_{\theta}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\bigg{)},$ (8)
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [51, 52, 53]. We set $\sigma=1$ for all experiments.
III-B3 Combined Objective
The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$ is
$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (9)
This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times. We will denote the combined loss as the baseline objective. Our baseline model’s objective is similar to DRAFT [54] but with a different ranking objective function.
Thus far, we have described an objective function that improves predictive performance:
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ (10)
In the next section, we address how to develop NN models that are robust against adversarial attacks.
IV Preliminaries on Adversarial Training
Adversarial training [32, 55, 21, 56] is a technique to improve robustness and generalization of NNs to uncertain data. Rather than training on pristine data, adversarial training perturbs the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [56] formulated adversarial training as solving a Min-Max robust optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}),$ (11)
where $\tilde{\boldsymbol{X}}=\{\tilde{\boldsymbol{x}}_{i}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)\}_{i=1}^{N}$, and $\mathcal{B}(\boldsymbol{x_{i}},\epsilon)$ is the $\ell_{\infty}$-ball of radius $\epsilon$ around $\boldsymbol{x_{i}}$.
In practice, the Min-Max problem posed in [56] is intractable to compute exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 11, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Certifiable Robustness
One such approximation is to use techniques from NN verification during training. For example, CROWN-IBP combines Convex Relaxation Based Backward Bounding Pass (CROWN) and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [55, 57]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [58, 55] showed the linear relaxation may be extended to the entire objective function:
[…]
This relaxation has several benefits: it is convex with respect to the NN weights, which enables efficient optimization, and by optimizing over an upper bound on the inner maximization, this relaxation represents accounting for a strong adversary during training.
IV-B Adversarial Attacks
Another way to approximate Eq. 19 is to use adversarial attacks. Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial attack algorithms typically provide a lower bound:
$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\approx\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}_{FGSM}})$ (13)
Fast Gradient Sign Method (FGSM) [59] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [56] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation to $\boldsymbol{x}$ within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
Considering the structure of the datasets as described in Section III, we perturb only the covariates $\boldsymbol{x}$ and not the time-to-event $t$. The perturbed covariates are calculated as:
$\boldsymbol{x}^{(0)}$ $=\boldsymbol{x}$ (14)
$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),$ (15)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)},$ (16)
where $k$ is the iteration step and $\alpha$ is a step size parameter.
IV-C Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [60]. Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$ ball:
$\boldsymbol{z}$ $\sim\mathcal{N}(0,\boldsymbol{I}),$ (17)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}+\Pi_{\mathcal{B}(0,\epsilon)}(\sqrt{\epsilon}\boldsymbol{I}\boldsymbol{z}).$ (18)
We next introduce our approach, SAWAR, to incorporate adversarial regularization with SA models for improved calibration, predictive generalization, and adversarial robustness.
V Survival Analysis with Adversarial Regularization
As discussed in Section I, a key open challenge in SA is in developing models that are robust to uncertainty and variability in the datasets. This paper aims to improve the adversarial robustness of an NN against covariate perturbations within the SA framework. In doing so, this paper proposes the first method for enabling adversarial robustness for fully-parametric NN-based SA, which improves calibration and stability on unexpected or noisy data.
The following sections detail our approach, which extends the baseline objective function in Section III-B3 to include adversarial regularization posed as a Min-Max optimization problem in Section Section IV [21, 55].
V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [15]. Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [16, 6]. Instead, SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of the baseline objective in Section III-B3:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e}).$ (19)
As NNs are highly non-convex, solving this optimization problem is difficult [61]. Therefore, we use the adversarial training methods discussed in Section IV to approximate/relax the inner maximization term of the Min-Max formulation.
V-B Adversarial Robustness Regularization
[21] shows that tight relaxations to the inner maximization of a Min-Max problem may over-regularize the NN, leading to poor predictive performance. Therefore, we extend the baseline objective in Eq. 9 with the adversarial robustness objective in Eq. 19 [55, 57] to balance predictive performance and adversarial robustness as:
$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}\right],$ (20)
where $\kappa\in(0,1)$.
We propose to use the open-source library auto_LiRPA [58] to relax the inner maximization term of the adversarial regularization with CROWN-IBP. auto_LiRPA seamlessly integrates with PyTorch for automatic backpropagation, computing a fully differentiable upper bound on the inner maximization term of Eq. 20 in a computationally feasible manner.
V-C SAWAR Training
We show the pseudo-code in Algorithm 1. The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [62] (12), where we use early stopping [63] with respect to the validation set to prevent overfitting (25, 26 and 27). We introduce the covariate perturbation attacks after 100 epochs (4, 5, 6 and 7). A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window (28) after 100 epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
Algorithm 1 Robust Training with CROWN-IBP
1:procedure train_step(ModelLoss, Data, $\epsilon$, $\kappa$, Train)
2:   $loss_{epoch}\leftarrow 0$
3:   for $x_{B},t_{B},y_{B}$ in Data do
4:      if $\epsilon>1e^{-20}$ then
5:         lossnatural $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
6:         $\tilde{x}_{B}\leftarrow\text{PerturbationLpNorm}(x_{B},\epsilon)$
[…]
8:         loss $\leftarrow\kappa\times loss_{natural}+(1-\kappa)\times loss_{robust}$
9:      else
10:         loss $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
11:      if Train then
12:         ADAM Optimizer Step with loss.backward()
13:      lossepoch $\leftarrow$ lossepoch + loss
14:   return lossepoch
15:
16:procedure train_robust(Model, Datatr, Dataval, $\kappa$)
17:   Init random seeds, ADAM optimizer, lr-scheduler, $\epsilon$-scheduler
18:   ModelLoss = Losswrapper(Model)
[…]
20:   for $t$ in Epochs do
21:      if $\epsilon$-scheduler reached $\epsilon_{max}$ then
22:         lr-scheduler Step
[…]
26:         Save Best Model
27:         lossbest $\leftarrow$ lossval
28:      $\epsilon$-scheduler Step
29:   return Model
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [34] via the adversarial regularization term in Eq. 20 and illustrate SAWAR hypothesis via extensive experiments in the next section.
VI Experiments
We conduct experiments over 10 benchmark SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete. As evaluation metrics, we consider the Concordance Index (CI) [64], the Integrated Brier Score (IBS) [65] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1. We show that SAWAR improves predictive accuracy, calibration, and population curve estimation.
VI-A Configuration
The experiments are conducted on a 12th Gen Intel(R) Core (TM) 16 GB RAM i9-12900 processor running a Windows operating system. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively. The NN $G_{\boldsymbol{\theta}}$ has two hidden layers of 50 neurons each and Leaky ReLU activation functions. lifelines [66], and sksurv [53] software packages are employed to ensure standardization and reproducibility in computing the SA metrics.
VI-B Datasets
We use the open-source API SurvSet [67], to download the benchmark SA datasets shown in Table I.
Dataset
$N$
$n_{fac}$
$n_{ohe}$
$n_{num}$
TRACE
1878
4
4
2
stagec
146
4
15
3
flchain
7874
4
26
6
Aids2
2839
3
11
1
Framingham
4699
2
12
5
dataDIVAT1
5943
3
14
2
prostate
502
6
16
9
zinc
431
11
18
2
retinopathy
394
5
9
2
LeukSurv
1043
2
24
5
TABLE I: Studied datasets from SurvSet. $n_{fac}$ is the number of categorical features, $n_{ohe}$ is number of binary features (one-hot-encoded) and $n_{num}$ is number of numerical features.
  1. 1.
    The TRACE [68] dataset is from a study on a subset of patients admitted after myocardial infarction to examine various risk factors.
  2. 2.
    The stagec [69] dataset is from a study exploring the prognostic value of flow cytometry for patients with stage C prostate cancer.
  3. 3.
    The flchain [70] dataset is from a study on the relationship between serum free light chain (FLC) and mortality of Olmsted Country residents aged 50 years or more.
  4. 4.
    The Aids2 [71] dataset is from patients diagnosed with AIDS in Australia before 1 July 1991.
  5. 5.
    The Framingham [72] dataset is from the first prospective study of cardiovascular disease with identification of risk factors and joint effects.
  6. 6.
    The dataDIVAT1 [73] dataset is from the first sample from the DIVAT Data Bank for French kidney transplant recipients from DIVAT cohort.
  7. 7.
    The prostate [74] dataset is from a randomised clinical trial comparing treatment for patients with stage 3 and stage 4 prostate cancer.
  8. 8.
    The zinc [75] is from the first study to examine association between tissue elemental zinc levels and the esophageal squamous cell carcinoma.
  9. 9.
    The retinopathy [76] is from the trial of laser coagulation as treatment to delay diabetic retinopathy.
  10. 10.
    The LeukSurv [77] is from the study on survival of acute myeloid leukemia and connection to spatial variation in survival.
We preprocess each dataset by one-hot encoding categorical covariates, followed by standard normalizing all covariates.
Figure 1: Distribution of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP. As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows. However, the credible interval of SAWAR does not drastically differ from the baseline.
VI-C Adversary Setup
A detailed insight into the robust evaluation procedure is shown in Algorithm 2. We use two adversarial attack methods when benchmarking the robustness of the adversarial regularization method (14, 15 and 19): FGSM attack (Section IV-B) and a novel worst-case attack. We define the worst-case attack as the maximum failure rate with respect to the covariate uncertainty:
$\tilde{\lambda}_{\theta}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (21)
and define the $\epsilon$-worst-case survival curve as:
$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\theta}t},$ (22)
which is the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates. For the exponential distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq. 21. We evaluate the SA metrics for $\epsilon\in[0,1]$ for the two attacks methods to study how sensitive each adversarial training method is to covariate perturbation (1).
Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP). However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
Algorithm 2 Adversarial Robustness Evaluation
1:procedure eval_robust(Model, Data, Metric, Method, $\epsilon$)
2:   $X,t,e$ = Data
3:   MetricValues=[]
4:   for $\epsilon$ in $0:0.1:1$ do
5:      if $\epsilon=0$ then
6:         $\lambda_{\theta}\leftarrow$ Model(X)
7:         MetricValues = [MetricValues, $\text{Metric}(\lambda_{\theta},t,e)$]
8:      else
9:         $\tilde{\lambda}_{\theta}\leftarrow$ ATTACK(Model,Data,Method,$\epsilon$)
10:         MetricValues = [MetricValues, $\text{Metric}(\tilde{\lambda}_{\theta},t,e)$]
11:   return MetricValues
12:
13:procedure attack(Model, Data, Method, $\epsilon$)
14:   $X,t,e$ = Data
15:   if Method = FGSM then
16:      ModelLoss = Losswrapper(Model)
17:      $\tilde{X}\leftarrow$ FGSM(ModelLoss, $X,t,e,\epsilon$)
18:      $\tilde{\lambda}_{\theta}\leftarrow$ Model($\tilde{X}$)
19:   if Method = worst-case then
20:      $\tilde{X}\leftarrow\text{PerturbationLpNorm}(X,\epsilon)$
21:      $\tilde{\lambda}_{\theta}\leftarrow\text{CROWN-IBP}(\text{Model})(\tilde{X},t,e)$
22:   return $\tilde{\lambda}_{\theta}$
VI-D Results
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
1.00 2.9 3.35 3.35 2.5 2.9 4.6 4.4 2.4 1.9 1.7 4.0 4.7 2.7 2.1 1.5
0.90 3.1 3.5 3.45 2.45 2.5 4.5 4.5 2.5 1.8 1.7 4.0 4.7 2.9 1.9 1.5
0.80 3.4 3.4 3.2 2.35 2.65 4.6 4.4 2.6 1.8 1.6 4.1 4.6 3.0 1.9 1.4
0.70 3.45 3.35 3.1 2.4 2.7 4.7 4.3 2.6 1.8 1.6 4.2 4.5 3.0 2.0 1.3
0.60 3.45 3.45 3.0 2.5 2.6 4.7 4.3 2.7 1.8 1.5 4.2 4.5 3.1 2.1 1.1
0.50 3.85 3.65 3.1 2.1 2.3 4.4 4.6 2.9 1.8 1.3 4.1 4.5 3.2 2.1 1.1
0.40 3.9 3.8 3.05 2.1 2.15 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.30 4.4 4.2 3.0 1.9 1.5 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.20 4.35 4.0 3.3 2.25 1.1 4.2 4.6 3.2 2.0 1.0 3.9 4.7 3.2 2.1 1.1
0.10 4.25 4.0 3.3 2.35 1.1 4.0 4.6 3.2 2.2 1.0 3.8 4.8 3.1 2.3 1.0
0.05 3.85 4.0 3.45 2.6 1.1 3.8 4.7 2.9 2.2 1.4 3.9 4.7 3.1 2.3 1.0
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
1.00 2.8 3.4 3.5 3.6 1.7 2.6 4.3 3.8 3.3 1.0 3.1 4.5 3.4 3.0 1.0
0.90 3.1 3.3 3.5 3.4 1.7 3.0 4.3 3.6 3.1 1.0 3.1 4.5 3.5 2.9 1.0
0.80 3.0 3.3 3.5 3.7 1.5 3.0 4.5 3.5 3.0 1.0 3.0 4.6 3.5 2.9 1.0
0.70 3.05 3.3 3.6 3.65 1.4 3.0 4.5 3.5 3.0 1.0 3.1 4.5 3.5 2.9 1.0
0.60 2.85 3.55 3.55 3.75 1.3 3.0 4.5 3.5 3.0 1.0 3.0 4.7 3.5 2.8 1.0
0.50 3.2 3.9 3.35 3.45 1.1 3.2 4.4 3.5 2.9 1.0 3.0 4.7 3.5 2.8 1.0
0.40 3.25 3.95 3.2 3.5 1.1 3.3 4.8 3.2 2.7 1.0 3.1 4.7 3.4 2.8 1.0
0.30 3.0 3.05 3.95 3.6 1.4 3.5 4.7 3.3 2.5 1.0 3.3 4.9 3.2 2.6 1.0
0.20 2.05 2.95 4.2 3.8 2.0 3.7 4.7 3.4 2.2 1.0 3.4 4.9 3.4 2.3 1.0
0.10 2.85 3.4 3.4 3.35 2.0 3.3 4.5 3.0 2.3 1.9 3.5 4.5 3.2 2.2 1.6
0.05 3.05 4.15 3.4 2.9 1.5 2.6 4.4 3.0 2.6 2.4 3.3 4.4 3.0 2.4 1.9
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
Furthermore, we analyze how these adversarial attacks impact the characteristics of the population survival curves.
VI-D1 Failure Rate Distribution
Let us assume that the input covariates are random variables. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates. Moreover, we may may treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process:
$\boldsymbol{x}\sim p(\boldsymbol{x});\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$ (23)
$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}$ (24)
Furthermore, we compute statistical quantities such as expectation, lower 5th quantile, and upper 95th quantile:
$LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})),$ (25)
$UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})).$ (26)
Fig. 1 visualizes the credible interval and mean of the survival function stochastic process in Eq. 24. The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 1). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 1). The variance of the failure rate decreasing leads to a more robust model against adversarial attacks. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 20, we trade-off individualized survival curves with vulnerability to adversarial attacks and a population survival curve which is completely resilient to adversarial attacks.
Figure 2: Adversarial robustness of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case attack. As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
However, despite exhibiting significantly higher adversarial robustness, the credible interval remains relatively unchanged compared to the baseline method.
VI-D2 Population Survival Curve
The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the ”closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig. 2). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 2 demonstrates the strong calibration of SAWAR model [29].
We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case attacks (Tables III and II). We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes.
The exact corresponding metric values for each $\epsilon$ of the worst-case attack are provided in Appendix A.
VII Conclusions
This paper is the first to introduce CROWN-IBP regularization for adversarial robustness training in SA with fully-parametric NN models. We benchmarked SAWAR method on 10 time-to-event datasets with standard adversarial regularization methods such as Gaussian Noise, FGSM, and PGD against the FGSM and worst-case attack. SAWAR’s empirical results indicate that CROWN-IBP regularization consistently enhances performance in SA with respect to CI and IBS, and NegLL for all $\epsilon$-perturbation magnitudes from 0 to 1. Moreover, compared to conventional adversarial regularization methods, SAWAR empirical evidence demonstrates that CROWN-IBP regularization is more resilient to input covariate perturbation attacks, where we see relative improvements to the baseline in CI and IBS of 1%-150% across all $\epsilon$-perturbation magnitudes (Appendix A). Thus adversarial regularization improves predictive performance, adversarial robustness, and calibration.
Future work will include the use the Weibull Proportional Hazard model to enable time varying hazard rates. This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller. The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is 29.46, 48.18, 56.27, 72.23, 102.44, 133.35, 152.17, 140.33, 91.78, 21.17, 9.63, 1.07 respectively. We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 baseline 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
noise 0.492 0.494 0.496 0.5 0.5 0.498 0.507 0.526 0.556 0.556 0.554 0.554
fgsm 0.498 0.495 0.498 0.497 0.5 0.502 0.508 0.529 0.554 0.557 0.555 0.553
pgd 0.498 0.497 0.497 0.5 0.501 0.501 0.507 0.531 0.557 0.558 0.556 0.554
SAWAR 0.567 0.575 0.577 0.579 0.58 0.58 0.579 0.579 0.579 0.579 0.579 0.579
Framingham baseline 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
noise 0.606 0.616 0.626 0.635 0.642 0.649 0.659 0.674 0.683 0.685 0.685 0.685
fgsm 0.607 0.611 0.617 0.622 0.626 0.635 0.652 0.669 0.682 0.686 0.687 0.686
pgd 0.567 0.578 0.588 0.599 0.61 0.624 0.642 0.659 0.68 0.688 0.688 0.687
SAWAR 0.694 0.697 0.699 0.701 0.702 0.704 0.704 0.705 0.705 0.704 0.704 0.704
LeukSurv baseline 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
noise 0.552 0.557 0.562 0.566 0.569 0.574 0.587 0.61 0.628 0.627 0.621 0.617
fgsm 0.546 0.549 0.554 0.563 0.561 0.566 0.583 0.602 0.624 0.624 0.621 0.62
pgd 0.537 0.536 0.538 0.549 0.546 0.558 0.577 0.598 0.62 0.622 0.619 0.617
SAWAR 0.483 0.498 0.51 0.532 0.553 0.583 0.589 0.607 0.617 0.623 0.632 0.636
TRACE baseline 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
noise 0.565 0.588 0.613 0.64 0.668 0.692 0.716 0.73 0.733 0.735 0.734 0.735
fgsm 0.553 0.577 0.603 0.63 0.658 0.688 0.713 0.722 0.727 0.73 0.733 0.734
pgd 0.557 0.582 0.609 0.636 0.664 0.696 0.723 0.731 0.733 0.732 0.733 0.733
SAWAR 0.73 0.733 0.736 0.738 0.739 0.739 0.74 0.74 0.739 0.738 0.738 0.737
dataDIVAT1 baseline 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
noise 0.534 0.539 0.551 0.571 0.587 0.612 0.636 0.656 0.654 0.64 0.635 0.631
fgsm 0.575 0.585 0.594 0.606 0.617 0.628 0.638 0.644 0.651 0.643 0.638 0.635
pgd 0.517 0.529 0.548 0.568 0.591 0.618 0.636 0.654 0.657 0.651 0.647 0.644
SAWAR 0.627 0.642 0.652 0.658 0.659 0.659 0.659 0.658 0.657 0.656 0.656 0.655
flchain baseline 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
noise 0.096 0.099 0.104 0.112 0.127 0.164 0.335 0.768 0.912 0.921 0.923 0.923
fgsm 0.102 0.116 0.159 0.288 0.599 0.852 0.91 0.917 0.923 0.927 0.927 0.925
pgd 0.102 0.127 0.235 0.538 0.824 0.904 0.917 0.921 0.925 0.929 0.929 0.929
SAWAR 0.84 0.925 0.927 0.929 0.93 0.93 0.929 0.929 0.93 0.93 0.931 0.931
prostate baseline 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
noise 0.49 0.503 0.514 0.524 0.533 0.541 0.553 0.566 0.576 0.572 0.577 0.576
fgsm 0.514 0.522 0.526 0.536 0.541 0.549 0.558 0.573 0.582 0.596 0.596 0.596
pgd 0.522 0.526 0.532 0.535 0.54 0.548 0.557 0.567 0.576 0.593 0.591 0.593
SAWAR 0.575 0.584 0.598 0.6 0.608 0.632 0.638 0.638 0.639 0.641 0.642 0.644
retinopathy baseline 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
noise 0.526 0.545 0.561 0.57 0.581 0.588 0.588 0.6 0.603 0.6 0.604 0.606
fgsm 0.47 0.485 0.51 0.536 0.569 0.578 0.591 0.6 0.599 0.602 0.606 0.607
pgd 0.471 0.481 0.509 0.53 0.567 0.577 0.591 0.599 0.598 0.601 0.606 0.607
SAWAR 0.588 0.597 0.611 0.616 0.619 0.615 0.621 0.62 0.62 0.619 0.617 0.617
stagec baseline 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
noise 0.471 0.481 0.505 0.49 0.514 0.534 0.582 0.697 0.721 0.712 0.688 0.697
fgsm 0.423 0.438 0.447 0.471 0.51 0.567 0.62 0.663 0.688 0.678 0.688 0.688
pgd 0.462 0.486 0.471 0.495 0.51 0.567 0.611 0.668 0.692 0.688 0.697 0.697
SAWAR 0.428 0.447 0.481 0.514 0.567 0.639 0.673 0.692 0.702 0.692 0.692 0.707
zinc baseline 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
noise 0.315 0.329 0.339 0.376 0.424 0.545 0.704 0.778 0.791 0.805 0.806 0.81
fgsm 0.339 0.357 0.384 0.439 0.55 0.657 0.738 0.776 0.787 0.804 0.809 0.812
pgd 0.333 0.351 0.384 0.439 0.55 0.671 0.738 0.776 0.79 0.8 0.809 0.808
SAWAR 0.383 0.464 0.533 0.649 0.707 0.746 0.766 0.779 0.789 0.787 0.795 0.8
TABLE IV: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller. The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is -44.13, -45.46, -46.68, -47.57, -47.86, -47.10, -45.03, -40.89, -33.06, -20.29, -11.34, -1.68 respectively (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 baseline 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
noise 0.254 0.253 0.25 0.245 0.236 0.22 0.196 0.161 0.13 0.121 0.121 0.122
fgsm 0.252 0.249 0.244 0.236 0.223 0.203 0.176 0.145 0.125 0.121 0.121 0.121
pgd 0.252 0.248 0.243 0.234 0.221 0.2 0.172 0.143 0.124 0.121 0.121 0.121
SAWAR 0.126 0.124 0.124 0.124 0.123 0.123 0.123 0.123 0.123 0.123 0.123 0.124
Framingham baseline 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
noise 0.818 0.818 0.816 0.809 0.778 0.681 0.495 0.294 0.174 0.128 0.119 0.116
fgsm 0.818 0.816 0.81 0.791 0.728 0.58 0.369 0.208 0.14 0.12 0.116 0.115
pgd 0.818 0.818 0.817 0.811 0.784 0.683 0.461 0.235 0.138 0.117 0.115 0.115
SAWAR 0.165 0.152 0.141 0.133 0.127 0.122 0.118 0.116 0.114 0.114 0.113 0.113
LeukSurv baseline 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
noise 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.157 0.134 0.12 0.114
fgsm 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.161 0.148 0.122 0.114 0.111
pgd 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.16 0.145 0.12 0.114 0.111
SAWAR 0.154 0.15 0.144 0.136 0.128 0.12 0.115 0.112 0.11 0.107 0.107 0.106
TRACE baseline 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
noise 0.634 0.634 0.632 0.622 0.596 0.536 0.435 0.319 0.23 0.183 0.172 0.167
fgsm 0.634 0.634 0.634 0.631 0.613 0.555 0.434 0.294 0.212 0.177 0.171 0.168
pgd 0.634 0.634 0.632 0.621 0.58 0.491 0.362 0.245 0.189 0.171 0.167 0.166
SAWAR 0.282 0.254 0.231 0.213 0.198 0.186 0.177 0.17 0.166 0.163 0.163 0.163
dataDIVAT1 baseline 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
noise 0.749 0.749 0.748 0.748 0.743 0.72 0.626 0.413 0.248 0.198 0.192 0.192
fgsm 0.749 0.749 0.748 0.747 0.74 0.699 0.547 0.322 0.215 0.193 0.192 0.192
pgd 0.749 0.749 0.749 0.748 0.746 0.724 0.582 0.305 0.192 0.184 0.186 0.187
SAWAR 0.266 0.241 0.224 0.212 0.202 0.193 0.187 0.183 0.18 0.178 0.178 0.178
flchain baseline 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
noise 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.807 0.25 0.087 0.063 0.052
fgsm 0.831 0.831 0.831 0.83 0.83 0.828 0.707 0.146 0.094 0.06 0.054 0.051
pgd 0.831 0.831 0.831 0.83 0.83 0.797 0.298 0.114 0.079 0.055 0.051 0.05
SAWAR 0.522 0.161 0.076 0.063 0.057 0.054 0.052 0.051 0.05 0.05 0.049 0.049
prostate baseline 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
noise 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.477 0.371 0.325 0.296
fgsm 0.53 0.53 0.53 0.53 0.53 0.53 0.529 0.509 0.413 0.29 0.257 0.246
pgd 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.502 0.396 0.28 0.254 0.244
SAWAR 0.508 0.482 0.441 0.393 0.348 0.309 0.278 0.251 0.229 0.214 0.209 0.205
retinopathy baseline 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
noise 0.662 0.657 0.645 0.622 0.583 0.523 0.432 0.326 0.245 0.204 0.196 0.195
fgsm 0.656 0.646 0.626 0.597 0.556 0.493 0.401 0.301 0.232 0.201 0.196 0.197
pgd 0.656 0.645 0.625 0.596 0.554 0.49 0.397 0.299 0.231 0.201 0.197 0.198
SAWAR 0.526 0.484 0.436 0.385 0.334 0.29 0.254 0.228 0.211 0.203 0.202 0.202
stagec baseline 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
noise 0.505 0.505 0.505 0.504 0.502 0.485 0.432 0.338 0.216 0.175 0.181 0.192
fgsm 0.505 0.505 0.505 0.503 0.494 0.458 0.382 0.268 0.174 0.161 0.169 0.178
pgd 0.505 0.505 0.505 0.503 0.493 0.454 0.377 0.261 0.169 0.159 0.167 0.176
SAWAR 0.491 0.476 0.45 0.408 0.354 0.296 0.24 0.195 0.168 0.16 0.162 0.167
zinc baseline 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
noise 0.908 0.907 0.907 0.904 0.893 0.838 0.625 0.32 0.168 0.109 0.096 0.089
fgsm 0.907 0.907 0.904 0.892 0.839 0.659 0.365 0.194 0.124 0.095 0.089 0.086
pgd 0.907 0.907 0.903 0.89 0.832 0.64 0.348 0.187 0.122 0.094 0.088 0.087
SAWAR 0.72 0.609 0.476 0.352 0.26 0.198 0.159 0.132 0.113 0.1 0.096 0.094
TABLE V: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 baseline 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
noise 3.84e+05 1.61e+05 6.68e+04 2.78e+04 1.14e+04 4.42e+03 1.83e+03 8.54e+02 5.64e+02 5.20e+02 5.19e+02 5.21e+02
fgsm 6.77e+04 3.50e+04 1.78e+04 8.89e+03 4.35e+03 2.14e+03 1.10e+03 6.66e+02 5.35e+02 5.19e+02 5.19e+02 5.20e+02
pgd 5.49e+04 2.86e+04 1.49e+04 7.58e+03 3.79e+03 1.92e+03 1.02e+03 6.42e+02 5.31e+02 5.19e+02 5.19e+02 5.19e+02
SAWAR 5.29e+02 5.24e+02 5.22e+02 5.22e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.22e+02 5.22e+02 5.22e+02
Framingham baseline 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
noise 1.27e+07 3.17e+06 8.03e+05 2.06e+05 5.38e+04 1.48e+04 4.92e+03 2.39e+03 1.72e+03 1.54e+03 1.51e+03 1.50e+03
fgsm 1.95e+06 5.96e+05 1.84e+05 5.70e+04 1.81e+04 6.33e+03 2.84e+03 1.85e+03 1.58e+03 1.51e+03 1.50e+03 1.49e+03
pgd 2.01e+07 4.36e+06 9.57e+05 2.13e+05 4.88e+04 1.21e+04 3.85e+03 1.97e+03 1.57e+03 1.50e+03 1.49e+03 1.49e+03
SAWAR 1.66e+03 1.62e+03 1.58e+03 1.56e+03 1.53e+03 1.52e+03 1.50e+03 1.50e+03 1.49e+03 1.49e+03 1.49e+03 1.49e+03
LeukSurv baseline 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
noise 5.46e+23 1.79e+21 7.98e+18 2.52e+16 8.53e+13 1.88e+11 5.25e+08 2.35e+06 3.37e+04 2.60e+03 1.21e+03 7.21e+02
fgsm 8.34e+16 1.79e+15 3.10e+13 4.86e+11 8.13e+09 1.49e+08 2.72e+06 7.03e+04 3.45e+03 6.20e+02 4.31e+02 3.59e+02
pgd 1.28e+15 3.89e+13 1.03e+12 2.67e+10 7.96e+08 2.30e+07 7.01e+05 2.80e+04 1.98e+03 4.65e+02 3.58e+02 3.17e+02
SAWAR 2.07e+04 8.38e+03 3.50e+03 1.45e+03 6.62e+02 3.45e+02 2.91e+02 2.70e+02 2.60e+02 2.54e+02 2.52e+02 2.50e+02
TRACE baseline 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
noise 9.11e+08 1.07e+08 1.34e+07 1.74e+06 2.43e+05 3.71e+04 6.63e+03 1.78e+03 8.35e+02 5.93e+02 5.51e+02 5.32e+02
fgsm 2.01e+11 9.10e+09 4.31e+08 2.24e+07 1.26e+06 8.26e+04 7.79e+03 1.61e+03 7.45e+02 5.74e+02 5.47e+02 5.35e+02
pgd 4.53e+08 4.81e+07 5.35e+06 6.12e+05 7.63e+04 1.07e+04 2.09e+03 8.37e+02 5.94e+02 5.40e+02 5.31e+02 5.26e+02
SAWAR 8.79e+02 7.60e+02 6.82e+02 6.29e+02 5.94e+02 5.69e+02 5.53e+02 5.41e+02 5.33e+02 5.28e+02 5.27e+02 5.27e+02
dataDIVAT1 baseline 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
noise 5.69e+08 7.64e+07 1.05e+07 1.44e+06 2.07e+05 3.17e+04 5.75e+03 1.61e+03 9.08e+02 7.81e+02 7.67e+02 7.64e+02
fgsm 2.48e+08 3.36e+07 4.65e+06 6.56e+05 9.64e+04 1.54e+04 3.12e+03 1.14e+03 8.16e+02 7.65e+02 7.61e+02 7.60e+02
pgd 1.85e+11 8.02e+09 3.47e+08 1.75e+07 9.39e+05 6.14e+04 5.80e+03 1.27e+03 8.03e+02 7.57e+02 7.56e+02 7.57e+02
SAWAR 9.18e+02 8.55e+02 8.21e+02 8.00e+02 7.86e+02 7.75e+02 7.66e+02 7.59e+02 7.55e+02 7.52e+02 7.51e+02 7.51e+02
flchain baseline 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
noise 2.18e+35 1.20e+31 7.58e+26 4.68e+22 3.48e+18 2.65e+14 2.21e+10 6.77e+06 4.45e+04 1.96e+03 1.22e+03 1.11e+03
fgsm 1.61e+20 3.47e+17 7.79e+14 1.86e+12 5.49e+09 4.02e+07 6.37e+05 1.86e+04 2.15e+03 1.15e+03 1.09e+03 1.08e+03
pgd 1.30e+16 7.91e+13 4.93e+11 3.47e+09 4.60e+07 1.33e+06 4.93e+04 4.99e+03 1.45e+03 1.10e+03 1.08e+03 1.07e+03
SAWAR 6.31e+03 2.22e+03 1.43e+03 1.19e+03 1.12e+03 1.09e+03 1.08e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03
prostate baseline 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
noise 7.00e+20 5.05e+18 3.75e+16 2.85e+14 2.28e+12 2.03e+10 2.59e+08 9.04e+06 4.85e+05 6.66e+04 2.94e+04 1.37e+04
fgsm 3.16e+14 1.04e+13 3.40e+11 1.16e+10 4.07e+08 1.52e+07 6.47e+05 3.56e+04 3.75e+03 1.03e+03 7.20e+02 5.82e+02
pgd 8.84e+12 4.29e+11 2.18e+10 1.14e+09 5.94e+07 3.27e+06 1.92e+05 1.44e+04 1.91e+03 6.40e+02 5.00e+02 4.39e+02
SAWAR 9.65e+03 3.88e+03 1.66e+03 8.48e+02 5.43e+02 4.30e+02 3.83e+02 3.59e+02 3.45e+02 3.37e+02 3.35e+02 3.34e+02
retinopathy baseline 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
noise 1.11e+04 5.61e+03 2.88e+03 1.51e+03 8.13e+02 4.68e+02 2.99e+02 2.19e+02 1.84e+02 1.70e+02 1.68e+02 1.68e+02
fgsm 7.43e+03 3.89e+03 2.07e+03 1.12e+03 6.33e+02 3.86e+02 2.63e+02 2.03e+02 1.78e+02 1.68e+02 1.67e+02 1.68e+02
pgd 7.30e+03 3.83e+03 2.04e+03 1.11e+03 6.25e+02 3.81e+02 2.60e+02 2.02e+02 1.77e+02 1.68e+02 1.67e+02 1.68e+02
SAWAR 4.02e+02 3.31e+02 2.79e+02 2.41e+02 2.14e+02 1.95e+02 1.82e+02 1.74e+02 1.69e+02 1.67e+02 1.67e+02 1.67e+02
stagec baseline 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
noise 2.07e+06 4.93e+05 1.14e+05 2.73e+04 6.49e+03 1.61e+03 4.19e+02 1.25e+02 5.50e+01 4.10e+01 3.96e+01 3.95e+01
fgsm 1.80e+05 5.45e+04 1.67e+04 5.23e+03 1.64e+03 5.30e+02 1.77e+02 7.43e+01 4.54e+01 3.96e+01 3.91e+01 3.92e+01
pgd 1.54e+05 4.77e+04 1.47e+04 4.63e+03 1.47e+03 4.84e+02 1.69e+02 7.23e+01 4.48e+01 3.93e+01 3.89e+01 3.90e+01
SAWAR 1.03e+03 5.79e+02 3.30e+02 1.94e+02 1.21e+02 8.21e+01 6.10e+01 4.88e+01 4.22e+01 3.89e+01 3.83e+01 3.81e+01
zinc baseline 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
noise 2.93e+07 4.72e+06 7.76e+05 1.32e+05 2.34e+04 4.45e+03 9.48e+02 2.56e+02 1.14e+02 8.17e+01 7.69e+01 7.53e+01
fgsm 2.50e+06 4.44e+05 9.49e+04 2.11e+04 4.89e+03 1.24e+03 3.66e+02 1.44e+02 9.01e+01 7.66e+01 7.53e+01 7.57e+01
pgd 2.55e+06 4.38e+05 8.27e+04 1.87e+04 4.43e+03 1.13e+03 3.40e+02 1.39e+02 8.88e+01 7.66e+01 7.55e+01 7.60e+01
SAWAR 1.31e+03 7.77e+02 4.72e+02 2.98e+02 1.98e+02 1.42e+02 1.10e+02 9.21e+01 8.28e+01 7.87e+01 7.80e+01 7.78e+01
TABLE VI: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.