### v3

# Survival Analysis with Adversarial Regularization

Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent work shows that Neural Networks (NNs) can capture complex relationships in SA. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade model performance. To address this, we leverage NN verification advances to create algorithms for robust, fully-parametric survival models. We introduce a robust loss function and use CROWN-IBP regularization to handle computational challenges in the Min-Max problem. Evaluating our approach on SurvSet datasets, we find that our Survival Analysis with Adversarial Regularization (SAWAR) method consistently outperforms baselines under various perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI). This demonstrates that adversarial regularization enhances SA performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets up to 150% across all perturbation magnitudes.
Code: https://github.com/mlpotter/SAWAR
Index Terms:
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP, Cox-Weibull
I Introduction
Survival Analysis (SA) models the time until an event occurs and has extensive applications. For example, SA has been used extensively in medicine for evaluating treatments in cancer patients and drug trials [1]. Early methods relied on linear parametric models [2], but recent advancements utilize Neural Networks (NNs) to handle complex time-to-event data [3].
A major challenge in Survival Analysis (SA) is developing models robust to dataset uncertainty and variability. This is particularly essential for clinical applications where treatment efficacy must be demonstrated across diverse patient demographics, institutions, and data acquisition settings [4, 5, 6]. Data uncertainty can arise from human errors, such as poor handwriting by physicians [7] and ineffective medical record-keeping [8], as well as inherent noise during data collection [9]. Many approaches assume perfect measurement of covariates, neglecting aleatoric uncertainty [10], which in SA is associated with noise in the log of measured time-to-event [11]. Further, adversaries can manipulate data to make models over- or under-confident [12]. Despite these issues and the sensitivity of NNs to input uncertainties, robustness techniques against data perturbations in SA remains underexplored.
To address these issues, this paper’s contributions include:
  • Survival Analysis with Adversarial Regularization (SAWAR) for training robust, fully-parametric, NN-based SA models using adversarial covariate perturbations and convex relaxations, enhancing model calibration and stability against noisy data.
  • Experiments demonstrating the sensitivity of existing NN-based SA methods to covariate perturbations across 10 publicly available medical SA datasets, highlighting an area for improvement in the current state of the art.
II Related Work
This section reviews NN literature for SA, with more details in [13]. The first use of NNs in SA was a shallow Feedforward Neural Network (FFNN) to approximate the survival function [14]. Later, deeper models were developed, such as Cox-nnet [15], a NN for Cox Proportional Hazard (CPH) modeling [11]. Building on this, DeepSurv [16] introduced a CPH NN with dropout layers for better generalization in personalized treatment. Then, mixture models such as DeepWeiSurv [17] further advanced the field by using NNs to regress Weibull mixture distribution parameters. These methods enhance modeling capability but do not always improve model calibration and robustness.
Recent works focus on improving model calibration [18, 19, 20], which ensures predicted confidence matches true likelihood [21]. However, few works address adversarial robustness in SA, where small input perturbations should not mislead the model [22]. Although better calibration does not always enhance generalization [23], integrating adversarial robustness improves model trustworthiness in calibration and and stability on unexpected data [24].
Existing adversarial training methods for SA involve tradeoffs between model complexity, performance, and incorporating physical assumptions. Most techniques use Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples from a survival distribution non-parametrically [25, 26, 27]. These models implicitly define the survival distribution without assuming a functional/analytical form, offering increased flexibility to fit various data realizations [28, 29, 30]. However, this flexibility can violate Bayesian Occam’s Razor, which penalizes overly complex models [30]. To balance complexity and performance, inductive biases are used in explicit parametric models [29]. One common bias in SA is the memoryless property, satisfied by the exponential distribution [28]. Thus, this paper employs the exponential CPH model, widely used in SA [31, 32, 33], as well as in queuing theory [34], reliability analysis [3], and telecommunications [35].
III Preliminaries on Survival Analysis
SA [11] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [36]. Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded. Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured. In summary, the dataset $\mathcal{D}$ has the following structure:
$\mathcal{D}=\{(\boldsymbol{x}_{i},t_{i},e_{i})\}_{i=1}^{N}.$ (1)
Being a statistical approach, we define the required probabilistic functions to model time-to-event data next.
III-A Cox Proportional Hazard Model
Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$. The survival function $S(t):=1-F(t)$ is the complement of this CDF [37, 11].
An exponential parametric survival model is expressed as
$S(t)=e^{-\Lambda(t)},$ (2)
[…]
$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta}$ (3)
is the instantaneous failure rate at time $t$.
To incorporate the instance’s covariates, we use the prevalent CPH model [2]:
$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=\lambda_{0}(t)e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})},$ (4)
where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative risk associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
The exponential CPH model characterizes the baseline hazard rate as $\lambda_{0}(t)=\lambda_{0}$. For simplicity, we assume an exponential baseline hazard and “absorb” the $\lambda$ term into the bias of the NN:
$\lambda_{0}e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
The exponential CPH model’s PDF, CDF, and complement CDF are $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively. $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
The population survival curve is the marginalization of the instance survival function over the covariates,
[…]
Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate [18]:
$S(t)$ $\approx\frac{1}{N}\sum_{i=1}^{N}S_{\boldsymbol{\theta}}(t|\boldsymbol{x}_{i}).$ (7)
Given these definitions, the next section describes the baseline objective for finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
III-B Baseline Objective Functions
We follow [38, 25] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the CPH model [11]:
$\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})=$ $\sum_{i=1}^{N}e_{i}\cdot\log f_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}})\ +$
$\sum_{i=1}^{N}(1-e_{i})\cdot\log S_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}}).$ (8)
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time of occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs.
III-B2 Ranking Objective
The $\mathcal{L}_{rank}$ objective is a ranking loss that penalizes an incorrect ordering of pairs $\{F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\}$, where instance $i$ with event time $t_{i}<t_{j}$ should have a higher probability of failure than instance $j$ at time $t_{i}$:
$\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ $=\sum_{i\neq j}A_{i,j}\eta\bigg{(}F_{\theta}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\bigg{)},$ (9)
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [38, 39]. We set $\sigma=1$ for all experiments.
III-B3 Combined Objective
The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$, is
$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (10)
This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times. We will denote the combined loss as the baseline objective. Our baseline model’s objective is similar to Deep Regularized Accelerated Failure Time (DRAFT) [25] but with a different ranking objective function. Thus, we consider the baseline training method for robust SA baseline method as DRAFT.
Thus far, we have described an objective function that improves predictive performance:
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (11)
In the following sections, we address how to develop NN models that are robust against adversarial perturbations.
IV Preliminaries on Adversarial Training
Adversarial training [22, 40, 41] improves robustness and generalization of NNs to uncertain data. Rather than training on pristine data, most adversarial training approaches perturb the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [42] formulated adversarial training as solving a Min-Max robust optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}),$ (12)
where $\tilde{\boldsymbol{X}}=\{\tilde{\boldsymbol{x}}_{i}\in\mathcal{B}(\boldsymbol{x}_{i},\epsilon)\}_{i=1}^{N}$, and $\mathcal{B}(\boldsymbol{x}_{i},\epsilon)$ is the $\ell_{\infty}$-ball of radius $\epsilon$ around $\boldsymbol{x}_{i}$.
In practice, the Min-Max problem posed in [42] is intractable to compute exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Certifiable Robustness
One such approximation is to use techniques from NN verification during training. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [40, 43]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [44, 40] showed the linear relaxation can be extended to the entire objective function:
[…]
This relaxation is convex with respect to NN weights, enabling efficient optimization, and by optimizing over an upper bound on the inner maximization, it accounts for a strong adversary during training.
IV-B Adversarial Perturbations
Another way to approximate Eq. 20 is to use adversarial perturbations. Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:
$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\geq\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}_{FGSM}).$ (14)
Fast Gradient Sign Method (FGSM) [45] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [42] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
Considering the structure of the datasets as described in Section III, we perturb only the covariates $\boldsymbol{x}$ and not the time-to-event $t$. The perturbed covariates are calculated as:
$\boldsymbol{x}^{(0)}$ $=\boldsymbol{x},$ (15)
$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),$ (16)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)},$ (17)
where $k$ is the iteration step and $\alpha$ is a step size parameter.
IV-C Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [46]. Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$-ball:
$\boldsymbol{z}$ $\sim\mathcal{N}(0,\boldsymbol{I}),$ (18)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}+\Pi_{\mathcal{B}(0,\epsilon)}(\sqrt{\epsilon}\boldsymbol{I}\boldsymbol{z}).$ (19)
Building on this foundation, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing models robust to dataset uncertainty and variability, as motivated in Section I. We propose a method for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and stability on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with adversarial regularization as a Min-Max optimization problem (Eq. 12) [41, 40], as detailed in the following sections.
V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [11]. Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of Eq. 10:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e}).$ (20)
We use the open-source, PyTorch-based library auto_LiRPA [44] to compute the CROWN-IBP upper bound to the inner maximization term in Eq. 20, which is fully differentiable for backpropogation.
V-B Adversarial Robustness Regularization
[41] shows that tight relaxations to the inner maximization of a Min-Max problem can over-regularize the NN, leading to poor predictive performance. Therefore, we balance the objective in Eq. 10 with the adversarial robustness objective in Eq. 20 [40, 43] to combine predictive performance and adversarial robustness as:
$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e})\right],$ (21)
where $\kappa\in(0,1)$.
V-C SAWAR Training
The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [47], where we use early stopping [48] with respect to the validation set to prevent overfitting We introduce the covariate perturbations after 100 epochs. A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after 100 epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [24] via the adversarial regularization term in Eq. 21 and show SAWAR’s improved performance via extensive experiments in the next section.
VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete. As evaluation metrics, we consider the Concordance Index (CI) [49], the Integrated Brier Score (IBS) [50] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1. To demonstrate that SAWAR produces SA models that perform well and are robust to dataset uncertainty, this section quantifies the improvement in predictive accuracy, calibration, and population curve estimation.
VI-A Configuration
The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively. The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions. We use lifelines [51] and sksurv [39] for computing SA metrics.
VI-B Datasets
We use the API SurvSet [52] to download the benchmark medical SA datasets shown in Table I and described below:
Dataset $N$ $n_{fac}$ $n_{ohe}$ $n_{num}$
TRACE 1878 4 4 2
stagec 146 4 15 3
flchain 7874 4 26 6
Aids2 2839 3 11 1
Framingham 4699 2 12 5
dataDIVAT1 5943 3 14 2
prostate 502 6 16 9
zinc 431 11 18 2
retinopathy 394 5 9 2
LeukSurv 1043 2 24 5
TABLE I: Studied datasets from SurvSet. $n_{fac}$ is the number of categorical features, $n_{ohe}$ is number of binary features (one-hot-encoded) and $n_{num}$ is number of numerical features.
  1. 1.
    The TRACE [53] dataset is from a study on a subset of patients admitted after myocardial infarction to examine various risk factors.
  2. 2.
    The stagec [54] dataset is from a study exploring the prognostic value of flow cytometry for patients with stage C prostate cancer.
  3. 3.
    The flchain [55] dataset is from a study on the relationship between serum free light chain (FLC) and mortality of Olmsted Country residents aged 50 years or more.
  4. 4.
    The Aids2 [56] dataset is from patients diagnosed with AIDS in Australia before 1 July 1991.
  5. 5.
    The Framingham [57] dataset is from the first prospective study of cardiovascular disease with identification of risk factors and joint effects.
  6. 6.
    The dataDIVAT1 [58] dataset is from the first sample from the DIVAT Data Bank for French kidney transplant recipients from DIVAT cohort.
  7. 7.
    The prostate [59] dataset is from a randomised clinical trial comparing treatment for patients with stage 3 and stage 4 prostate cancer.
  8. 8.
    The zinc [60] is from the first study to examine association between tissue elemental zinc levels and the esophageal squamous cell carcinoma.
  9. 9.
    The retinopathy [61] is from the trial of laser coagulation as treatment to delay diabetic retinopathy.
  10. 10.
    The LeukSurv [62] is from the study on survival of acute myeloid leukemia and connection to spatial variation in survival.
We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates.
Figure 1: Distribution of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP. As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows. However, the credible interval of SAWAR does not drastically differ from the DRAFT.
Figure 2: Adversarial robustness of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case perturbation. As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
VI-C Adversary Setup
We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B) and a novel worst-case perturbation. FGSM is applied by using the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective Eq. 20, such that:
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
1.00 2.9 3.35 3.35 2.5 2.9 4.6 4.4 2.4 1.9 1.7 4.0 4.7 2.7 2.1 1.5
0.90 3.1 3.5 3.45 2.45 2.5 4.5 4.5 2.5 1.8 1.7 4.0 4.7 2.9 1.9 1.5
0.80 3.4 3.4 3.2 2.35 2.65 4.6 4.4 2.6 1.8 1.6 4.1 4.6 3.0 1.9 1.4
0.70 3.45 3.35 3.1 2.4 2.7 4.7 4.3 2.6 1.8 1.6 4.2 4.5 3.0 2.0 1.3
0.60 3.45 3.45 3.0 2.5 2.6 4.7 4.3 2.7 1.8 1.5 4.2 4.5 3.1 2.1 1.1
0.50 3.85 3.65 3.1 2.1 2.3 4.4 4.6 2.9 1.8 1.3 4.1 4.5 3.2 2.1 1.1
0.40 3.9 3.8 3.05 2.1 2.15 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.30 4.4 4.2 3.0 1.9 1.5 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.20 4.35 4.0 3.3 2.25 1.1 4.2 4.6 3.2 2.0 1.0 3.9 4.7 3.2 2.1 1.1
0.10 4.25 4.0 3.3 2.35 1.1 4.0 4.6 3.2 2.2 1.0 3.8 4.8 3.1 2.3 1.0
0.05 3.85 4.0 3.45 2.6 1.1 3.8 4.7 2.9 2.2 1.4 3.9 4.7 3.1 2.3 1.0
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
1.00 2.8 3.4 3.5 3.6 1.7 2.6 4.3 3.8 3.3 1.0 3.1 4.5 3.4 3.0 1.0
0.90 3.1 3.3 3.5 3.4 1.7 3.0 4.3 3.6 3.1 1.0 3.1 4.5 3.5 2.9 1.0
0.80 3.0 3.3 3.5 3.7 1.5 3.0 4.5 3.5 3.0 1.0 3.0 4.6 3.5 2.9 1.0
0.70 3.05 3.3 3.6 3.65 1.4 3.0 4.5 3.5 3.0 1.0 3.1 4.5 3.5 2.9 1.0
0.60 2.85 3.55 3.55 3.75 1.3 3.0 4.5 3.5 3.0 1.0 3.0 4.7 3.5 2.8 1.0
0.50 3.2 3.9 3.35 3.45 1.1 3.2 4.4 3.5 2.9 1.0 3.0 4.7 3.5 2.8 1.0
0.40 3.25 3.95 3.2 3.5 1.1 3.3 4.8 3.2 2.7 1.0 3.1 4.7 3.4 2.8 1.0
0.30 3.0 3.05 3.95 3.6 1.4 3.5 4.7 3.3 2.5 1.0 3.3 4.9 3.2 2.6 1.0
0.20 2.05 2.95 4.2 3.8 2.0 3.7 4.7 3.4 2.2 1.0 3.4 4.9 3.4 2.3 1.0
0.10 2.85 3.4 3.4 3.35 2.0 3.3 4.5 3.0 2.3 1.9 3.5 4.5 3.2 2.2 1.6
0.05 3.05 4.15 3.4 2.9 1.5 2.6 4.4 3.0 2.6 2.4 3.3 4.4 3.0 2.4 1.9
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\left(\text{FGSM}(\mathcal{L},\boldsymbol{x})\right)}.$ (22)
We define the worst-case perturbation as the maximum failure rate with respect to the covariate uncertainty:
$\tilde{\lambda}_{\boldsymbol{\theta}}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (23)
and the $\epsilon$-worst-case survival curve as the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (24)
For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq. 23. We evaluate the SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ is no perturbation to input covariates, for the two perturbations methods to study how sensitive each adversarial training method is to covariate perturbation.
Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP). However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
VI-D Results
Furthermore, we analyze how adversarial perturbations impact the characteristics of the population survival curves.
VI-D1 Failure Rate Distribution
Let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$. Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}$ (25)
Furthermore, we compute statistical quantities such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
Fig. 1 visualizes the credible interval and mean of the survival function stochastic process in Eq. 25. The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 1). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 1). The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 21, we trade-off individualized survival curves with vulnerability to adversarial perturbations and a population survival curve which is completely resilient to adversarial perturbations.
However, despite exhibiting significantly higher adversarial robustness, the credible interval usually remains relatively unchanged compared to the DRAFT method.
VI-D2 Population Survival Curve
The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the “closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig. 2). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 2 demonstrates the strong calibration of SAWAR model [19].
We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations (Tables III and II). We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes. More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case perturbation for each dataset are available in the code repository linked in the abstract.
VII Conclusions
This paper introduces CROWN-IBP regularization for enhancing adversarial robustness in fully-parametric SA using NN models. We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods such as Gaussian noise, FGSM, and PGD, as well as the novel worst-case perturbation.
Our empirical results show that CROWN-IBP regularization consistently improves performance in SA across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1. Compared to conventional adversarial regularization techniques, SAWAR demonstrates superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1% to 150% across all $\epsilon$-perturbation magnitudes.
In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration. This approach effectively mitigates data uncertainty in SA and improves generalization across unseen datasets.
Future work will include the use the Weibull CPH model to enable time varying hazard rates. This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
%$\%\Delta$ 29.46 48.18 56.27 72.23 102.44 133.35 152.17 140.33 91.78 21.17 9.63 1.07
TABLE IV: The relative percent change in Concordance Index metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets. A higher relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 DRAFT 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
Noise 0.492 0.494 0.496 0.5 0.5 0.498 0.507 0.526 0.556 0.556 0.554 0.554
FGSM 0.498 0.495 0.498 0.497 0.5 0.502 0.508 0.529 0.554 0.557 0.555 0.553
PGD 0.498 0.497 0.497 0.5 0.501 0.501 0.507 0.531 0.557 0.558 0.556 0.554
SAWAR 0.567 0.575 0.577 0.579 0.58 0.58 0.579 0.579 0.579 0.579 0.579 0.579
Framingham DRAFT 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
Noise 0.606 0.616 0.626 0.635 0.642 0.649 0.659 0.674 0.683 0.685 0.685 0.685
FGSM 0.607 0.611 0.617 0.622 0.626 0.635 0.652 0.669 0.682 0.686 0.687 0.686
PGD 0.567 0.578 0.588 0.599 0.61 0.624 0.642 0.659 0.68 0.688 0.688 0.687
SAWAR 0.694 0.697 0.699 0.701 0.702 0.704 0.704 0.705 0.705 0.704 0.704 0.704
LeukSurv DRAFT 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
Noise 0.552 0.557 0.562 0.566 0.569 0.574 0.587 0.61 0.628 0.627 0.621 0.617
FGSM 0.546 0.549 0.554 0.563 0.561 0.566 0.583 0.602 0.624 0.624 0.621 0.62
PGD 0.537 0.536 0.538 0.549 0.546 0.558 0.577 0.598 0.62 0.622 0.619 0.617
SAWAR 0.483 0.498 0.51 0.532 0.553 0.583 0.589 0.607 0.617 0.623 0.632 0.636
TRACE DRAFT 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
Noise 0.565 0.588 0.613 0.64 0.668 0.692 0.716 0.73 0.733 0.735 0.734 0.735
FGSM 0.553 0.577 0.603 0.63 0.658 0.688 0.713 0.722 0.727 0.73 0.733 0.734
PGD 0.557 0.582 0.609 0.636 0.664 0.696 0.723 0.731 0.733 0.732 0.733 0.733
SAWAR 0.73 0.733 0.736 0.738 0.739 0.739 0.74 0.74 0.739 0.738 0.738 0.737
dataDIVAT1 DRAFT 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
Noise 0.534 0.539 0.551 0.571 0.587 0.612 0.636 0.656 0.654 0.64 0.635 0.631
FGSM 0.575 0.585 0.594 0.606 0.617 0.628 0.638 0.644 0.651 0.643 0.638 0.635
PGD 0.517 0.529 0.548 0.568 0.591 0.618 0.636 0.654 0.657 0.651 0.647 0.644
SAWAR 0.627 0.642 0.652 0.658 0.659 0.659 0.659 0.658 0.657 0.656 0.656 0.655
flchain DRAFT 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
Noise 0.096 0.099 0.104 0.112 0.127 0.164 0.335 0.768 0.912 0.921 0.923 0.923
FGSM 0.102 0.116 0.159 0.288 0.599 0.852 0.91 0.917 0.923 0.927 0.927 0.925
PGD 0.102 0.127 0.235 0.538 0.824 0.904 0.917 0.921 0.925 0.929 0.929 0.929
SAWAR 0.84 0.925 0.927 0.929 0.93 0.93 0.929 0.929 0.93 0.93 0.931 0.931
prostate DRAFT 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
Noise 0.49 0.503 0.514 0.524 0.533 0.541 0.553 0.566 0.576 0.572 0.577 0.576
FGSM 0.514 0.522 0.526 0.536 0.541 0.549 0.558 0.573 0.582 0.596 0.596 0.596
PGD 0.522 0.526 0.532 0.535 0.54 0.548 0.557 0.567 0.576 0.593 0.591 0.593
SAWAR 0.575 0.584 0.598 0.6 0.608 0.632 0.638 0.638 0.639 0.641 0.642 0.644
retinopathy DRAFT 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
Noise 0.526 0.545 0.561 0.57 0.581 0.588 0.588 0.6 0.603 0.6 0.604 0.606
FGSM 0.47 0.485 0.51 0.536 0.569 0.578 0.591 0.6 0.599 0.602 0.606 0.607
PGD 0.471 0.481 0.509 0.53 0.567 0.577 0.591 0.599 0.598 0.601 0.606 0.607
SAWAR 0.588 0.597 0.611 0.616 0.619 0.615 0.621 0.62 0.62 0.619 0.617 0.617
stagec DRAFT 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
Noise 0.471 0.481 0.505 0.49 0.514 0.534 0.582 0.697 0.721 0.712 0.688 0.697
FGSM 0.423 0.438 0.447 0.471 0.51 0.567 0.62 0.663 0.688 0.678 0.688 0.688
PGD 0.462 0.486 0.471 0.495 0.51 0.567 0.611 0.668 0.692 0.688 0.697 0.697
SAWAR 0.428 0.447 0.481 0.514 0.567 0.639 0.673 0.692 0.702 0.692 0.692 0.707
zinc DRAFT 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
Noise 0.315 0.329 0.339 0.376 0.424 0.545 0.704 0.778 0.791 0.805 0.806 0.81
FGSM 0.339 0.357 0.384 0.439 0.55 0.657 0.738 0.776 0.787 0.804 0.809 0.812
PGD 0.333 0.351 0.384 0.439 0.55 0.671 0.738 0.776 0.79 0.8 0.809 0.808
SAWAR 0.383 0.464 0.533 0.649 0.707 0.746 0.766 0.779 0.789 0.787 0.795 0.8
TABLE V: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
%$\%\Delta$44.13 -45.46 -46.68 -47.57 -47.86 -47.10 -45.03 -40.89 -33.06 -20.29 -11.34 -1.68
TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets. A lower relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 DRAFT 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
Noise 0.254 0.253 0.25 0.245 0.236 0.22 0.196 0.161 0.13 0.121 0.121 0.122
FGSM 0.252 0.249 0.244 0.236 0.223 0.203 0.176 0.145 0.125 0.121 0.121 0.121
PGD 0.252 0.248 0.243 0.234 0.221 0.2 0.172 0.143 0.124 0.121 0.121 0.121
SAWAR 0.126 0.124 0.124 0.124 0.123 0.123 0.123 0.123 0.123 0.123 0.123 0.124
Framingham DRAFT 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
Noise 0.818 0.818 0.816 0.809 0.778 0.681 0.495 0.294 0.174 0.128 0.119 0.116
FGSM 0.818 0.816 0.81 0.791 0.728 0.58 0.369 0.208 0.14 0.12 0.116 0.115
PGD 0.818 0.818 0.817 0.811 0.784 0.683 0.461 0.235 0.138 0.117 0.115 0.115
SAWAR 0.165 0.152 0.141 0.133 0.127 0.122 0.118 0.116 0.114 0.114 0.113 0.113
LeukSurv DRAFT 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
Noise 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.157 0.134 0.12 0.114
FGSM 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.161 0.148 0.122 0.114 0.111
PGD 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.16 0.145 0.12 0.114 0.111
SAWAR 0.154 0.15 0.144 0.136 0.128 0.12 0.115 0.112 0.11 0.107 0.107 0.106
TRACE DRAFT 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
Noise 0.634 0.634 0.632 0.622 0.596 0.536 0.435 0.319 0.23 0.183 0.172 0.167
FGSM 0.634 0.634 0.634 0.631 0.613 0.555 0.434 0.294 0.212 0.177 0.171 0.168
PGD 0.634 0.634 0.632 0.621 0.58 0.491 0.362 0.245 0.189 0.171 0.167 0.166
SAWAR 0.282 0.254 0.231 0.213 0.198 0.186 0.177 0.17 0.166 0.163 0.163 0.163
dataDIVAT1 DRAFT 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
Noise 0.749 0.749 0.748 0.748 0.743 0.72 0.626 0.413 0.248 0.198 0.192 0.192
FGSM 0.749 0.749 0.748 0.747 0.74 0.699 0.547 0.322 0.215 0.193 0.192 0.192
PGD 0.749 0.749 0.749 0.748 0.746 0.724 0.582 0.305 0.192 0.184 0.186 0.187
SAWAR 0.266 0.241 0.224 0.212 0.202 0.193 0.187 0.183 0.18 0.178 0.178 0.178
flchain DRAFT 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
Noise 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.807 0.25 0.087 0.063 0.052
FGSM 0.831 0.831 0.831 0.83 0.83 0.828 0.707 0.146 0.094 0.06 0.054 0.051
PGD 0.831 0.831 0.831 0.83 0.83 0.797 0.298 0.114 0.079 0.055 0.051 0.05
SAWAR 0.522 0.161 0.076 0.063 0.057 0.054 0.052 0.051 0.05 0.05 0.049 0.049
prostate DRAFT 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
Noise 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.477 0.371 0.325 0.296
FGSM 0.53 0.53 0.53 0.53 0.53 0.53 0.529 0.509 0.413 0.29 0.257 0.246
PGD 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.502 0.396 0.28 0.254 0.244
SAWAR 0.508 0.482 0.441 0.393 0.348 0.309 0.278 0.251 0.229 0.214 0.209 0.205
retinopathy DRAFT 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
Noise 0.662 0.657 0.645 0.622 0.583 0.523 0.432 0.326 0.245 0.204 0.196 0.195
FGSM 0.656 0.646 0.626 0.597 0.556 0.493 0.401 0.301 0.232 0.201 0.196 0.197
PGD 0.656 0.645 0.625 0.596 0.554 0.49 0.397 0.299 0.231 0.201 0.197 0.198
SAWAR 0.526 0.484 0.436 0.385 0.334 0.29 0.254 0.228 0.211 0.203 0.202 0.202
stagec DRAFT 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
Noise 0.505 0.505 0.505 0.504 0.502 0.485 0.432 0.338 0.216 0.175 0.181 0.192
FGSM 0.505 0.505 0.505 0.503 0.494 0.458 0.382 0.268 0.174 0.161 0.169 0.178
PGD 0.505 0.505 0.505 0.503 0.493 0.454 0.377 0.261 0.169 0.159 0.167 0.176
SAWAR 0.491 0.476 0.45 0.408 0.354 0.296 0.24 0.195 0.168 0.16 0.162 0.167
zinc DRAFT 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
Noise 0.908 0.907 0.907 0.904 0.893 0.838 0.625 0.32 0.168 0.109 0.096 0.089
FGSM 0.907 0.907 0.904 0.892 0.839 0.659 0.365 0.194 0.124 0.095 0.089 0.086
PGD 0.907 0.907 0.903 0.89 0.832 0.64 0.348 0.187 0.122 0.094 0.088 0.087
SAWAR 0.72 0.609 0.476 0.352 0.26 0.198 0.159 0.132 0.113 0.1 0.096 0.094
TABLE VII: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 DRAFT 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
Noise 3.84e+05 1.61e+05 6.68e+04 2.78e+04 1.14e+04 4.42e+03 1.83e+03 8.54e+02 5.64e+02 5.20e+02 5.19e+02 5.21e+02
FGSM 6.77e+04 3.50e+04 1.78e+04 8.89e+03 4.35e+03 2.14e+03 1.10e+03 6.66e+02 5.35e+02 5.19e+02 5.19e+02 5.20e+02
PGD 5.49e+04 2.86e+04 1.49e+04 7.58e+03 3.79e+03 1.92e+03 1.02e+03 6.42e+02 5.31e+02 5.19e+02 5.19e+02 5.19e+02
SAWAR 5.29e+02 5.24e+02 5.22e+02 5.22e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.22e+02 5.22e+02 5.22e+02
Framingham DRAFT 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
Noise 1.27e+07 3.17e+06 8.03e+05 2.06e+05 5.38e+04 1.48e+04 4.92e+03 2.39e+03 1.72e+03 1.54e+03 1.51e+03 1.50e+03
FGSM 1.95e+06 5.96e+05 1.84e+05 5.70e+04 1.81e+04 6.33e+03 2.84e+03 1.85e+03 1.58e+03 1.51e+03 1.50e+03 1.49e+03
PGD 2.01e+07 4.36e+06 9.57e+05 2.13e+05 4.88e+04 1.21e+04 3.85e+03 1.97e+03 1.57e+03 1.50e+03 1.49e+03 1.49e+03
SAWAR 1.66e+03 1.62e+03 1.58e+03 1.56e+03 1.53e+03 1.52e+03 1.50e+03 1.50e+03 1.49e+03 1.49e+03 1.49e+03 1.49e+03
LeukSurv DRAFT 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
Noise 5.46e+23 1.79e+21 7.98e+18 2.52e+16 8.53e+13 1.88e+11 5.25e+08 2.35e+06 3.37e+04 2.60e+03 1.21e+03 7.21e+02
FGSM 8.34e+16 1.79e+15 3.10e+13 4.86e+11 8.13e+09 1.49e+08 2.72e+06 7.03e+04 3.45e+03 6.20e+02 4.31e+02 3.59e+02
PGD 1.28e+15 3.89e+13 1.03e+12 2.67e+10 7.96e+08 2.30e+07 7.01e+05 2.80e+04 1.98e+03 4.65e+02 3.58e+02 3.17e+02
SAWAR 2.07e+04 8.38e+03 3.50e+03 1.45e+03 6.62e+02 3.45e+02 2.91e+02 2.70e+02 2.60e+02 2.54e+02 2.52e+02 2.50e+02
TRACE DRAFT 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
Noise 9.11e+08 1.07e+08 1.34e+07 1.74e+06 2.43e+05 3.71e+04 6.63e+03 1.78e+03 8.35e+02 5.93e+02 5.51e+02 5.32e+02
FGSM 2.01e+11 9.10e+09 4.31e+08 2.24e+07 1.26e+06 8.26e+04 7.79e+03 1.61e+03 7.45e+02 5.74e+02 5.47e+02 5.35e+02
PGD 4.53e+08 4.81e+07 5.35e+06 6.12e+05 7.63e+04 1.07e+04 2.09e+03 8.37e+02 5.94e+02 5.40e+02 5.31e+02 5.26e+02
SAWAR 8.79e+02 7.60e+02 6.82e+02 6.29e+02 5.94e+02 5.69e+02 5.53e+02 5.41e+02 5.33e+02 5.28e+02 5.27e+02 5.27e+02
dataDIVAT1 DRAFT 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
Noise 5.69e+08 7.64e+07 1.05e+07 1.44e+06 2.07e+05 3.17e+04 5.75e+03 1.61e+03 9.08e+02 7.81e+02 7.67e+02 7.64e+02
FGSM 2.48e+08 3.36e+07 4.65e+06 6.56e+05 9.64e+04 1.54e+04 3.12e+03 1.14e+03 8.16e+02 7.65e+02 7.61e+02 7.60e+02
PGD 1.85e+11 8.02e+09 3.47e+08 1.75e+07 9.39e+05 6.14e+04 5.80e+03 1.27e+03 8.03e+02 7.57e+02 7.56e+02 7.57e+02
SAWAR 9.18e+02 8.55e+02 8.21e+02 8.00e+02 7.86e+02 7.75e+02 7.66e+02 7.59e+02 7.55e+02 7.52e+02 7.51e+02 7.51e+02
flchain DRAFT 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
Noise 2.18e+35 1.20e+31 7.58e+26 4.68e+22 3.48e+18 2.65e+14 2.21e+10 6.77e+06 4.45e+04 1.96e+03 1.22e+03 1.11e+03
FGSM 1.61e+20 3.47e+17 7.79e+14 1.86e+12 5.49e+09 4.02e+07 6.37e+05 1.86e+04 2.15e+03 1.15e+03 1.09e+03 1.08e+03
PGD 1.30e+16 7.91e+13 4.93e+11 3.47e+09 4.60e+07 1.33e+06 4.93e+04 4.99e+03 1.45e+03 1.10e+03 1.08e+03 1.07e+03
SAWAR 6.31e+03 2.22e+03 1.43e+03 1.19e+03 1.12e+03 1.09e+03 1.08e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03
prostate DRAFT 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
Noise 7.00e+20 5.05e+18 3.75e+16 2.85e+14 2.28e+12 2.03e+10 2.59e+08 9.04e+06 4.85e+05 6.66e+04 2.94e+04 1.37e+04
FGSM 3.16e+14 1.04e+13 3.40e+11 1.16e+10 4.07e+08 1.52e+07 6.47e+05 3.56e+04 3.75e+03 1.03e+03 7.20e+02 5.82e+02
PGD 8.84e+12 4.29e+11 2.18e+10 1.14e+09 5.94e+07 3.27e+06 1.92e+05 1.44e+04 1.91e+03 6.40e+02 5.00e+02 4.39e+02
SAWAR 9.65e+03 3.88e+03 1.66e+03 8.48e+02 5.43e+02 4.30e+02 3.83e+02 3.59e+02 3.45e+02 3.37e+02 3.35e+02 3.34e+02
retinopathy DRAFT 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
Noise 1.11e+04 5.61e+03 2.88e+03 1.51e+03 8.13e+02 4.68e+02 2.99e+02 2.19e+02 1.84e+02 1.70e+02 1.68e+02 1.68e+02
FGSM 7.43e+03 3.89e+03 2.07e+03 1.12e+03 6.33e+02 3.86e+02 2.63e+02 2.03e+02 1.78e+02 1.68e+02 1.67e+02 1.68e+02
PGD 7.30e+03 3.83e+03 2.04e+03 1.11e+03 6.25e+02 3.81e+02 2.60e+02 2.02e+02 1.77e+02 1.68e+02 1.67e+02 1.68e+02
SAWAR 4.02e+02 3.31e+02 2.79e+02 2.41e+02 2.14e+02 1.95e+02 1.82e+02 1.74e+02 1.69e+02 1.67e+02 1.67e+02 1.67e+02
stagec DRAFT 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
Noise 2.07e+06 4.93e+05 1.14e+05 2.73e+04 6.49e+03 1.61e+03 4.19e+02 1.25e+02 5.50e+01 4.10e+01 3.96e+01 3.95e+01
FGSM 1.80e+05 5.45e+04 1.67e+04 5.23e+03 1.64e+03 5.30e+02 1.77e+02 7.43e+01 4.54e+01 3.96e+01 3.91e+01 3.92e+01
PGD 1.54e+05 4.77e+04 1.47e+04 4.63e+03 1.47e+03 4.84e+02 1.69e+02 7.23e+01 4.48e+01 3.93e+01 3.89e+01 3.90e+01
SAWAR 1.03e+03 5.79e+02 3.30e+02 1.94e+02 1.21e+02 8.21e+01 6.10e+01 4.88e+01 4.22e+01 3.89e+01 3.83e+01 3.81e+01
zinc DRAFT 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
Noise 2.93e+07 4.72e+06 7.76e+05 1.32e+05 2.34e+04 4.45e+03 9.48e+02 2.56e+02 1.14e+02 8.17e+01 7.69e+01 7.53e+01
FGSM 2.50e+06 4.44e+05 9.49e+04 2.11e+04 4.89e+03 1.24e+03 3.66e+02 1.44e+02 9.01e+01 7.66e+01 7.53e+01 7.57e+01
PGD 2.55e+06 4.38e+05 8.27e+04 1.87e+04 4.43e+03 1.13e+03 3.40e+02 1.39e+02 8.88e+01 7.66e+01 7.55e+01 7.60e+01
SAWAR 1.31e+03 7.77e+02 4.72e+02 2.98e+02 1.98e+02 1.42e+02 1.10e+02 9.21e+01 8.28e+01 7.87e+01 7.80e+01 7.78e+01
TABLE VIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.