### v4

# Survival Analysis with Adversarial Regularization

Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent research indicates that Neural Networks (NNs) can effectively capture complex data patterns in SA, whereas simple generalized linear models often fall short in this regard. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade NN model performance. To address this, we leverage advances in NN verification to develop training objectives for robust, fully-parametric SA models. Specifically, we propose an adversarially robust loss function based on a Min-Max optimization problem. We employ CROWN-Interval Bound Propagation (CROWN-IBP) to tackle the computational challenges inherent in solving this Min-Max problem. Evaluated over 10 SurvSet datasets, our method, Survival Analysis with Adversarial Regularization (SAWAR), consistently outperforms baseline adversarial training methods and state-of-the-art (SOTA) deep SA models across various covariate perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI) metrics. Thus, we demonstrate that adversarial robustness enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets by up to 150% compared to baselines.
Code: https://github.com/mlpotter/SAWAR
Index Terms:
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP, Cox-Weibull
I Introduction
Survival Analysis (SA) models the time until an event occurs with extensive applications in various fields including medicine, system reliability, economics, and marketing. Medical research uses Survival Analysis (SA) to evaluate treatment efficacy and illness progression [1]. Engineers assess a system’s reliability with SA to improve design and maintenance [2]. Economists and social scientists use SA (also known as duration analysis) to investigate the length of unemployment spells and other major life events [3]. Marketing uses SA to aid in the examination of customer retention and churn rates [4]. While many of these domains have traditionally relied on linear parametric SA models [5], recent advancements have leveraged Neural Networks (NNs) to handle such complex time-to-event data [6]
A major challenge in SA is developing models that effectively handle complex time-to-event data while also ensuring robustness against dataset uncertainty, accurate calibration, and strong generalization. For instance, data uncertainty in medical applications can arise from various sources, including human factors such as illegible handwriting by physicians [7] and poor medical record-keeping [8], as well as inherent noise from medical equipment during data collection [9], termed as aleatoric uncertainty. Most of the existing approaches address aleatoric uncertainty by assuming noise in the logarithm of measured time-to-event data [10]. However, they often overlook uncertainty originating from the model’s input covariates.
SA models ought to generalize over various input covariates to sustain high accuracy in real-life applications. In medical studies, this entails high accuracy over diverse patient demographics, institutions, and data acquisition settings. While traditional SA models, such as Linear Cox-Weibull model [5], offer inductive biases specific to time-to-event analysis that aid generalization, they struggle with complex data relationships. On the other hand, while NN models can manage complex data, they may not generalize well to unseen data [11]. Combining the inductive biases of traditional SA models with NN techniques—such as modeling intensity or hazard functions with NNs —can both improve generalization and handle complex data [12]. Nonetheless, NN in SA may still suffer from poor model calibration, even with good generalization.
NNs are particularly sensitive to input perturbations [13]. Adversaries can exploit this sensitivity to manipulate data, causing models to make over- or under-confident predictions [14, 15]. Additionally, data perturbations may be intentionally introduced with techniques like differential privacy to protect sensitive information, such as personally identifiable information of patients [16]. Although better calibration does not always enhance generalization [17], integrating adversarial robustness to input covariate perturbation can jointly improve model stability, trustworthiness, and calibration when encountering unseen data [18].
Despite these challenges and the sensitivity of NNs to input uncertainties, robustness techniques for handling data perturbations in NN-based SA remain underexplored. To address these challenges, we make the following contributions:
  • We propose a novel training objective, called Survival Analysis with Adversarial Regularization (SAWAR), designed to create robust, fully-parametric NN-based SA models by incorporating adversarial covariate perturbations via convex relaxations, and thereby improve model calibration and generalization against noisy data.
  • Extensive experiments across 10 publicly-available SA datasets reveal the sensitivity of existing NN-based SA models to covariate perturbations with respect to Concordance Index (CI), Integrated Brier Score (IBS), and Negative Log Likelihood (NegLL) metrics, as well as survival curves.
  • SAWAR enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization by up to 150% compared to baseline adversarial training methods and state-of-the-art (SOTA) NN-based SA models.
II Related Work
The integration of NNs into SA began with initial models that replaced traditional linear approaches. Early works used a shallow Feedforward Neural Network (FFNN) to replace the linear estimate of the log-hazard in a Cox Proportional Hazard (CPH) model [19]. Building on this, Cox-nnet [20] substituted the input covariates of the linear log-hazard function of a CPH model [21] with a deeper NN. Subsequently, DeepSurv [22] further advanced this approach by replacing the entire linear estimate with a deep NN and adding dropout to improve generalization. The trend continued with mixture models like DeepWeiSurv [23], which used NNs to regress Weibull mixture distribution parameters, offering increased modeling flexibility while maintaining traditional statistical model biases.
As NN models evolved, the SA field saw the introduction of non-parametric approaches. For instance, DeepHit utilized NNs to learn the distribution of survival times directly without assumptions about the underlying stochastic process [24]. Even more flexible techniques emerged, such as those employing Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples non-parametrically [25, 26, 27], with CGAN serving as an implicit likelihood [28]. This approach defines the survival distribution without assuming a specific functional form [29, 30]. For a more detailed overview of deep learning methods for SA, we refer the readers to [12].
Despite these advances, integrating NNs with traditional SA models has been limited in addressing model calibration and robustness as discussed in Section I. Recent research has thus focused on improving model calibration [31, 32, 33] to ensure that predicted confidence levels reflect true likelihoods [34]. However, better calibration does not necessarily enhance generalization [17]. On the other hand, integrating adversarial robustness has been shown to improve both calibration and generalization when dealing with unexpected data [18]. Nonetheless, there remains a gap in addressing adversarial robustness in SA to ensure that where small input perturbations do not mislead model predictions [15].
Existing adversarial training methods for SA often face trade-offs between model complexity, prediction performance, and the incorporation of physical assumptions. While Generative Adversarial Network (GAN)-inspired methods incorporate adversarial components through a minimax game between the generator and discriminator [35, 36], they typically exhibit training instability and may violate Bayesian Occam’s Razor that advocates simpler models [30]. To balance complexity, interpretability, and generalization, several methods adopt time-to-event inductive biases through specific functional forms of SA models with deep NN and/or non-parametric approaches [22, 20, 23, 37].
One such inductive bias in SA is the memoryless property, satisfied by the exponential distribution [29]. This paper employs the exponential CPH model (c.f. Section III), as it is widely used in SA [38, 39, 40], as well as in related fields like queuing theory [41], reliability analysis [6], and telecommunications [42]. Crucially, this fully-parametric approach allows us to create an adversarially robust NN-based SA model that outperforms SOTA NN-based SA models including such as Deep Regularized Accelerated Failure Time (DRAFT) [25] and Adversarial Autoencoder with a Deep Cox Proportional Hazard (AAE-Cox) [37]. AAE-Cox aims to balance modeling flexibility and inductive biases by combining adversarial GAN autoencoders with Deep Cox proportional networks. AAE-Cox has been shown to outperform methods like DeepSurv, TCAP [43], and DCAP [44].
III Preliminaries on Survival Analysis
Figure 1: Visual of time-to-event dataset with exact event times and right censoring. The solid circle on each patient line indicates when the event occurred for a given patient.
SA [21] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, a study of such events happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow-up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring  [45]. Such an SA dataset consists of $N$ independent instances, each denoted by index $i\in[1,\ldots,N]$, for which the time between entry to a study and the subsequent event is recorded. Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that govern the event of interest) are measured. A visual of time-to-event data with right censoring is shown in Fig. 1. In summary, the dataset $\mathcal{D}$ has the following structure:
$\mathcal{D}=\{(\boldsymbol{x}_{i},t_{i},e_{i})\}_{i=1}^{N}.$ (1)
Being a statistical approach, we define the required probabilistic functions to model time-to-event data next.
III-A Cox Proportional Hazard Model
Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$. The survival function $S(t):=1-F(t)$ is the complement of this CDF [46, 21].
An exponential SA model has the parametric form:
$S(t)=e^{-\Lambda(t)},$ (2)
[…]
$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta},$ (3)
is the instantaneous failure rate at time $t$.
To incorporate each instance’s covariates in predicting failure rate, we use the prevalent CPH model [5]:
$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=\lambda_{0}(t)e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})},$ (4)
where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative hazard associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
The exponential CPH model assumes a time-independent baseline hazard rate $\lambda_{0}(t)=\lambda_{0}$. This allows us to “absorb” the $\lambda_{0}$ term into the bias of the NN:
$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
Further, the exponential CPH model’s PDF, CDF, and complement CDF are given by $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively. $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
The population survival curve is defined as the marginalization of the instance survival function over the covariates:
[…]
Since the distribution of covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate over a dataset $\mathcal{D}$ [31]:
$S(t)$ $\approx\frac{1}{N}\sum_{i=1}^{N}S_{\boldsymbol{\theta}}(t|\boldsymbol{x}_{i}).$ (7)
Given these definitions, the next section describes the baseline objective for estimating model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
III-B Baseline Objective Functions
We employ an objective function that combines negative log likelihood (Section III-B1) and rank correlation (Section III-B2), following [47, 25].
III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the logarithm of the time-to-event likelihood over $\mathcal{D}$ as specified by the CPH model [21]:
$\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})=$ $\sum_{i=1}^{N}e_{i}\cdot\log f_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}})\ +$
$\sum_{i=1}^{N}(1-e_{i})\cdot\log S_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}}).$ (8)
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time that occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs, where it is known that no event occurred at least up until the censoring time.
III-B2 Ranking Objective
The $\mathcal{L}_{rank}$ objective is a ranking loss that penalizes an incorrect ordering of pairs $\{F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\}$, where instance $i$ with event time $t_{i}<t_{j}$ should have a higher probability of failure than instance $j$ at time $t_{i}$:
$\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ $=\sum_{i\neq j}A_{i,j}\eta\bigg{(}F_{\theta}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\bigg{)},$ (9)
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [47, 48]. Two instances are comparable if the instance with lower observed time experienced an event, i.e., if $A_{i,j}=\mathds{1}(t_{i}<t_{j}\cap e_{i}=1)$. We set $\sigma=1$ for all experiments. This objective is a continuous differentiable proxy for the non-continuous non-differentiable concordance index [49].
III-B3 Combined Objective
The combined objective over the right censored dataset $\mathcal{D}$ is:
$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}),$ (10)
where $w$ is the hazard versus ranking objective weight trade-off. Accordingly, $\boldsymbol{\theta}$ is estimated by solving the following optimization problem:
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (11)
This combined objective provides a trade-off between maximizing the likelihood of observing $\mathcal{D}$ and maximizing a rank correlation between the instance hazard scores and the observed failure times. From here on, we term this combined loss as the baseline objective. We note that an NN-based SA model DRAFT [25] also employs a combined objective. Our baseline objective incorporates a different ranking objective than DRAFT and outperforms DRAFT as discussed in Section VI.
IV Preliminaries on Adversarial Training
Adversarial training [36, 50, 51] improves robustness and generalization of NNs to uncertain data. Rather than training on pristine data, most adversarial training approaches perturb the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [13] formulated adversarial training as solving a Min-Max robust optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}),$ (12)
where $\tilde{\boldsymbol{X}}=\{\tilde{\boldsymbol{x}}_{i}\in\mathcal{B}(\boldsymbol{x}_{i},\epsilon)\}_{i=1}^{N}$, and $\mathcal{B}(\boldsymbol{x}_{i},\epsilon)$ is the $\ell_{\infty}$-ball of radius $\epsilon$ around $\boldsymbol{x}_{i}$.
In practice, the Min-Max problem posed in [13] is intractable to solve exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over input covariates with the current NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Verifiable Robustness
One such approximation is to use techniques from NN verification during training. Neural network robustness verification algorithms determine the output neuron’s upper and lower bounds for all inputs in a set $\mathcal{B}$, typically constrained by a norm-bounded perturbation [51]. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data within the set $\mathcal{B}$. [50, 51, 52]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [53, 50] showed the linear relaxation can be extended to the entire objective function:
[…]
IV-B Adversarial Perturbations
Another way to approximate Eq. 12 is to use adversarial perturbations.
IV-B1 Fast Gradient Sign Method (FGSM)
While CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:
$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\geq\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}_{FGSM}).$ (14)
FGSM [54] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [13] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
Considering the structure of the datasets as described in Section III, we perturb only the covariates $\boldsymbol{x}$ and not the time-to-event $t$. The perturbed covariates are calculated as:
$\boldsymbol{x}^{(0)}$ $=\boldsymbol{x},$ (15)
$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),\forall k$
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)}$
where $k\in[1,\ldots,K-1]$ is the iteration step and $\alpha=\epsilon/K$ is a step size parameter.
IV-B2 Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [55]. Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$-ball:
$\boldsymbol{z}$ $\sim\mathcal{N}(0,\boldsymbol{I}),$ (16)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}+\Pi_{\mathcal{B}(0,\epsilon)}(\sqrt{\epsilon}\boldsymbol{I}\boldsymbol{z}).$ (17)
Building on this foundation, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing models robust to dataset uncertainty, generalizable, and calibrated. We propose a training objective for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and generalization on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with adversarial regularization as a Min-Max optimization problem (Eq. 12) [51, 50], as detailed in the following sections.
V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise in the logarithm of measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [21]. Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of Eq. 10:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e}).$ (18)
We use the open-source, PyTorch-based library auto_LiRPA [53] to compute the CROWN-IBP upper bound to the inner maximization term in Eq. 18, which is fully differentiable for backpropogation.
V-B Adversarial Robustness Regularization
[51] shows that tight relaxations to the inner maximization of a Min-Max problem can over-regularize the NN, leading to poor predictive performance. Therefore, we balance the objective in Eq. 10 with the adversarial robustness objective in Eq. 18 [50, 52] to combine predictive performance and adversarial robustness as:
$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e})\right],$ (19)
where $\kappa\in(0,1)$.
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [18] via the adversarial regularization term in Eq. 19 and show SAWAR’s improved performance via extensive experiments in the next section.
VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (Section VI-A) to quantify the impact of perturbation after the robust training is complete. To demonstrate that SAWAR produces SA models that are both high-performing and robust to dataset uncertainty, this section quantifies improvements in predictive accuracy, calibration, and population curve estimation compared to common adversarial training techniques (Section IV-B) including FGSM, Random Noise and PGD, as well as SOTA NN-based SA models, such as DRAFT and AAE-Cox.
VI-A Datasets
We use the API SurvSet [56] to download the benchmark medical SA datasets shown in Table I and described below:
Dataset $N$ $n_{fac}$ $n_{ohe}$ $n_{num}$
TRACE 1878 4 4 2
stagec 146 4 15 3
flchain 7874 4 26 6
Aids2 2839 3 11 1
Framingham 4699 2 12 5
dataDIVAT1 5943 3 14 2
prostate 502 6 16 9
zinc 431 11 18 2
retinopathy 394 5 9 2
LeukSurv 1043 2 24 5
TABLE I: Studied datasets from SurvSet. $n_{fac}$ is the number of categorical features, $n_{ohe}$ is number of binary features (one-hot-encoded) and $n_{num}$ is number of numerical features.
  1. 1.
    The TRACE [57] dataset is from a study on a subset of patients admitted after myocardial infarction to examine various risk factors.
  2. 2.
    The stagec [58] dataset is from a study exploring the prognostic value of flow cytometry for patients with stage C prostate cancer.
  3. 3.
    The flchain [59] dataset is from a study on the relationship between serum free light chain (FLC) and mortality of Olmsted Country residents aged 50 years or more.
  4. 4.
    The Aids2 [60] dataset is from patients diagnosed with AIDS in Australia before 1 July 1991.
  5. 5.
    The Framingham [61] dataset is from the first prospective study of cardiovascular disease with identification of risk factors and joint effects.
  6. 6.
    The dataDIVAT1 [62] dataset is from the first sample from the DIVAT Data Bank for French kidney transplant recipients from DIVAT cohort.
  7. 7.
    The prostate [63] dataset is from a randomised clinical trial comparing treatment for patients with stage 3 and stage 4 prostate cancer.
  8. 8.
    The zinc [64] is from the first study to examine association between tissue elemental zinc levels and the esophageal squamous cell carcinoma.
  9. 9.
    The retinopathy [65] is from the trial of laser coagulation as treatment to delay diabetic retinopathy.
  10. 10.
    The LeukSurv [66] is from the study on survival of acute myeloid leukemia and connection to spatial variation in survival.
We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively.
VI-B Experiment Setup
The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM. The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions. For AAE-Cox, we adopt the publicly available code and hyperparameters [37]. For DRAFT, we use the same hyperparameters as SAWAR as described below, excluding hyperparameters related to adversarial regularization.
The NN parameters are optimized by solving Eq. (19) using stochastic gradient descent with the ADAM optimizer [67]. We set $w=\frac{1}{\text{batch\_size}}$ and use early stopping [68] with respect to the validation set to prevent overfitting. We introduce the covariate perturbations after a set number of epochs that varies per dataset. A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after the aforementioned set number of epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude of $0.5$ is achieved.
As evaluation metrics, we consider the CI [49], the IBS [69] and the NegLL at varying $\epsilon$-perturbation magnitudes from 0 to 1. We use lifelines [70] and sksurv [48] for computing SA metrics.
Figure 2: Distribution of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP. As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows. However, the credible interval of SAWAR does not drastically differ from the DRAFT.
Figure 3: Adversarial robustness of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case perturbation. As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
VI-C Adversary Setup
We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B1) and a novel worst-case perturbation. The perturbed covariates $\boldsymbol{\tilde{x}}_{FGSM}$ with respect to the input $\boldsymbol{x}$ are found via Eq. 15 and the corresponding failure rate becomes:
$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\boldsymbol{\tilde{x}}_{FGSM}}.$ (20)
We define the worst-case perturbation as the maximum failure rate with respect to the covariate uncertainty:
$\tilde{\lambda}_{\boldsymbol{\theta}}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (21)
which we solve via CROWN-IBP (Section IV-B1) by optimizing a tight approximation of the upper bound of the maximization term. Accordingly, the $\epsilon$-worst-case survival curve has the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
$\tilde{S}(t|\boldsymbol{\tilde{x}})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (22)
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR
1.00 3.8 4.4 3.1 2.7 3.8 3.2 5.4 5.6 3.2 2.3 2.6 1.9 5.1 5.8 3.6 2.6 2.5 1.4
0.90 3.8 4.4 3.25 2.6 3.9 3.05 5.4 5.6 3.4 2.4 2.6 1.6 5.1 5.8 3.6 2.6 2.5 1.4
0.80 3.9 4.4 3.65 2.9 3.6 2.55 5.4 5.5 3.6 2.5 2.5 1.5 5.1 5.8 3.6 2.6 2.5 1.4
0.70 4.45 4.75 3.5 3.05 3.1 2.15 5.4 5.5 3.6 2.5 2.6 1.4 5.1 5.8 3.6 2.7 2.5 1.3
0.60 4.9 5.0 3.5 3.0 2.8 1.8 5.3 5.6 3.6 2.7 2.5 1.3 5.2 5.7 3.6 2.7 2.5 1.3
0.50 5.0 4.95 3.6 3.05 2.8 1.6 5.3 5.5 3.7 2.7 2.5 1.3 5.2 5.7 3.9 2.6 2.3 1.3
0.40 5.3 4.7 3.7 3.3 2.6 1.4 5.3 5.5 3.9 2.7 2.3 1.3 5.3 5.6 3.9 2.6 2.3 1.3
0.30 5.1 5.2 3.7 3.0 2.6 1.4 5.3 5.5 3.9 2.7 2.3 1.3 5.1 5.7 3.9 2.8 2.2 1.3
0.20 5.5 5.1 3.9 2.7 2.3 1.5 5.3 5.5 3.9 2.9 2.1 1.3 4.9 5.8 4.0 2.9 2.1 1.3
0.10 5.2 5.2 4.0 3.2 1.9 1.5 5.0 5.7 4.1 3.0 1.9 1.3 4.5 5.9 4.2 3.1 1.9 1.4
0.05 5.4 5.0 3.9 3.3 2.0 1.4 4.6 5.6 4.3 3.4 1.6 1.5 4.5 5.9 4.3 3.3 1.7 1.3
0.00 3.3 4.3 4.5 4.1 2.5 2.3 2.7 4.2 4.7 4.7 2.5 2.2 3.0 5.4 4.9 4.2 1.8 1.7
TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better. Bold and underline indicates the best and second best method, respectively.
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR
1.00 3.6 3.6 3.6 4.4 4.2 1.6 2.6 4.35 3.65 3.55 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
0.90 3.4 3.9 3.6 4.3 4.4 1.4 2.9 4.25 3.65 3.35 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
0.80 3.3 3.8 3.7 4.3 4.6 1.3 2.7 4.25 3.75 3.45 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
0.70 3.4 3.8 3.8 4.1 4.6 1.3 2.8 4.65 3.65 3.15 5.75 1.0 3.1 4.8 3.5 2.7 5.9 1.0
0.60 3.2 3.65 3.5 4.2 5.3 1.15 2.8 4.95 3.8 2.7 5.75 1.0 3.1 4.8 3.5 2.7 5.9 1.0
0.50 3.0 3.6 3.5 4.1 5.6 1.2 3.1 4.8 3.8 2.7 5.6 1.0 3.1 4.8 3.5 2.7 5.9 1.0
0.40 2.9 3.8 3.4 4.0 5.8 1.1 3.3 4.8 3.7 2.5 5.7 1.0 3.1 4.8 3.6 2.6 5.9 1.0
0.30 2.8 3.9 3.6 3.3 6.0 1.4 3.6 4.6 3.5 2.4 5.9 1.0 3.1 4.9 3.6 2.5 5.9 1.0
0.20 3.2 3.6 3.9 2.9 6.0 1.4 3.7 4.9 3.4 2.3 5.6 1.1 3.5 5.1 3.3 2.3 5.8 1.0
0.10 3.2 4.0 4.2 3.6 4.0 2.0 3.9 5.5 3.6 2.9 3.8 1.3 3.9 5.8 4.1 3.1 3.0 1.1
0.05 3.4 4.2 4.4 4.1 2.6 2.3 3.6 5.1 4.3 4.1 2.2 1.7 3.7 5.6 4.4 3.9 1.9 1.5
0.00 3.3 4.3 4.5 4.1 2.5 2.3 2.7 4.2 4.7 4.7 2.5 2.2 3.0 5.4 4.9 4.2 1.8 1.7
TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better. Bold and underline indicates the best and second best method, respectively.
For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We evaluate SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ represents no perturbation to input covariates, in order to study how sensitive different adversarial training methods are to covariate perturbations.
To simulate realistic conditions, we assume that the 10 benchmark survival analysis datasets are “clean” and contain various covariates (e.g., age, drug dosage in mg, tumor size) and survival outcomes (e.g., time until death, censored observations). By applying adversarial attack techniques, we perturb these input covariates to maximize the survival analysis error.
The perturbations are designed to be subtle yet impactful, meaning they are small in magnitude but cause significant changes in the predicted survival times. For example, since the data is standard normalized, a perturbation of $\epsilon=0.1$ would shift an input covariate by 0.1 standard deviations from its current value.
For instance, consider a scenario where one of the covariates is the body weight of a patient in kilograms. If patient $i$ has a body weight of 70 kg, an adversarial perturbation might increase this body weight to to 75 kg. While this change is minimal, it is selected because it may cause a drastic change in the predicted survival likelihood over the next 10 years for a non-robust NN-based SA model. In reality, a 5 kg change in weight should not have such a significant effect on survival, but the model may incorrectly interpret this small perturbation as a substantial change in survival hazard, highlighting its sensitivity to minor changes in input covariates.
VI-D Results
We analyze how adversarial perturbations impact the characteristics of the population survival curves.
VI-D1 Comparison to SOTA
We show the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations in Tables III and II, respectively. We empirically show that SAWAR through CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes. Thus, SAWAR has better generalization performance, better calibration, and better adversarial robustness by up to 150% compared to baseline adversarial training methods and SOTA NN-based SA models. More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case and the FGSM adversarial attacks for each dataset are available in the code repository linked in the abstract.
VI-D2 Failure Rate Distribution
Each instance $x$ in the dataset has its own survival curve $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$. All instance’s survival curves are visualized as the credible interval, where $x$ is sampled from the dataset distribution $p(x)$. Thus, let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$. Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}.$ (23)
We compute the statistical quantities of $S_{\boldsymbol{\theta}}(t)$ such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
Fig. 2 visualizes the credible interval and mean of the survival function stochastic process in Eq. 23. The 90% credible interval of the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 2). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 2). However, this trend has an exception with AAE-Cox. We hypothesize that since AAE-Cox is not truly an adversarial robust method (i.e., no perturbations to the input covariates during training), and by introducing more parameters via an encoder of a GAN, there are more model parameters to attack. The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 19, we introduce a trade-off between individualized survival curves with vulnerability to adversarial perturbations and a population survival curve that is completely resilient to adversarial perturbations. In doing so, while exhibiting significantly higher adversarial robustness, SAWAR’s credible interval remains relatively unchanged compared to the DRAFT method.
VI-D3 Population Survival Curve
The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the “closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig. 3). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 3 demonstrates the strong calibration of SAWAR model [32].
VII Conclusions
This paper introduces a robust training objective based on CROWN-IBP regularization to enhance adversarial robustness in fully-parametric NN-based SA models, improving both model calibration and generalization. We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods like Gaussian noise, FGSM, and PGD, as well as SOTA models such as DRAFT and AAE-Cox.
Our empirical results demonstrate that SAWAR consistently improves SA performance across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1. Compared to traditional adversarial regularization techniques, SAWAR shows superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1
In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration, effectively mitigating data uncertainty in SA and improving generalization on unseen datasets.
Future work will involve extending the approach to the Weibull CPH model to enable time-varying hazard rates, which will require adapting auto_LiRPA to handle power operations such as $t^{k}$, where $k$ is a learnable parameter.
Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
%$\%\Delta$ 72.8 79.69 92.54 90.82 81.56 66.21 41.01 12.83 3.19 1.88 1.72 1.6
TABLE IV: The relative percent change in the Concordance Index metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the worst-case adversarial attack. A higher relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Dataset Algorithm
Aids2 DRAFT 0.516 0.519 0.522 0.525 0.527 0.53 0.536 0.551 0.566 0.572 0.572 0.57
Noise 0.513 0.515 0.518 0.521 0.525 0.529 0.535 0.548 0.568 0.571 0.57 0.569
FGSM 0.512 0.516 0.52 0.523 0.526 0.532 0.537 0.55 0.565 0.569 0.568 0.566
PGD 0.506 0.509 0.513 0.518 0.523 0.528 0.535 0.552 0.567 0.569 0.568 0.567
AAE-Cox 0.504 0.508 0.507 0.506 0.505 0.501 0.499 0.504 0.51 0.561 0.573 0.573
SAWAR 0.565 0.568 0.568 0.569 0.569 0.57 0.57 0.57 0.57 0.57 0.569 0.569
Framingham DRAFT 0.618 0.634 0.649 0.663 0.676 0.688 0.699 0.708 0.716 0.72 0.721 0.722
Noise 0.606 0.622 0.637 0.653 0.666 0.679 0.691 0.703 0.713 0.718 0.719 0.719
FGSM 0.582 0.598 0.611 0.626 0.641 0.655 0.671 0.691 0.708 0.715 0.715 0.715
PGD 0.539 0.559 0.582 0.605 0.627 0.647 0.669 0.693 0.71 0.717 0.717 0.717
AAE-Cox 0.514 0.522 0.529 0.533 0.539 0.553 0.569 0.57 0.582 0.73 0.733 0.733
SAWAR 0.721 0.725 0.729 0.731 0.733 0.734 0.735 0.736 0.736 0.737 0.737 0.737
LeukSurv DRAFT 0.589 0.593 0.596 0.598 0.601 0.602 0.608 0.62 0.63 0.632 0.633 0.633
Noise 0.545 0.548 0.545 0.547 0.551 0.56 0.581 0.609 0.63 0.628 0.626 0.624
FGSM 0.511 0.514 0.52 0.525 0.535 0.549 0.579 0.612 0.637 0.638 0.636 0.635
PGD 0.498 0.498 0.501 0.51 0.521 0.543 0.576 0.616 0.64 0.643 0.64 0.638
AAE-Cox 0.559 0.552 0.542 0.555 0.548 0.55 0.548 0.537 0.545 0.631 0.658 0.656
SAWAR 0.495 0.51 0.524 0.536 0.552 0.57 0.588 0.609 0.633 0.658 0.669 0.673
TRACE DRAFT 0.581 0.605 0.633 0.66 0.685 0.708 0.726 0.736 0.741 0.743 0.744 0.745
Noise 0.576 0.598 0.621 0.645 0.669 0.691 0.712 0.728 0.737 0.742 0.744 0.745
FGSM 0.581 0.603 0.629 0.653 0.674 0.697 0.717 0.732 0.739 0.743 0.744 0.744
PGD 0.571 0.595 0.621 0.646 0.668 0.691 0.714 0.73 0.739 0.743 0.744 0.745
AAE-Cox 0.432 0.44 0.447 0.451 0.46 0.47 0.489 0.524 0.628 0.743 0.746 0.747
SAWAR 0.714 0.722 0.728 0.733 0.735 0.737 0.739 0.742 0.744 0.746 0.747 0.747
dataDIVAT1 DRAFT 0.573 0.585 0.598 0.611 0.625 0.636 0.646 0.653 0.657 0.657 0.656 0.655
Noise 0.532 0.541 0.552 0.562 0.575 0.589 0.603 0.619 0.631 0.636 0.635 0.635
FGSM 0.546 0.555 0.567 0.578 0.588 0.6 0.612 0.626 0.641 0.647 0.646 0.646
PGD 0.523 0.534 0.545 0.557 0.569 0.581 0.599 0.621 0.641 0.648 0.648 0.648
AAE-Cox 0.597 0.589 0.589 0.585 0.58 0.571 0.571 0.566 0.596 0.662 0.663 0.663
SAWAR 0.64 0.645 0.649 0.654 0.658 0.661 0.664 0.665 0.667 0.668 0.669 0.67
flchain DRAFT 0.109 0.111 0.115 0.122 0.133 0.158 0.239 0.566 0.905 0.917 0.92 0.921
Noise 0.166 0.131 0.111 0.123 0.149 0.224 0.433 0.792 0.911 0.917 0.919 0.918
FGSM 0.115 0.153 0.245 0.431 0.72 0.885 0.91 0.914 0.918 0.922 0.922 0.922
PGD 0.165 0.268 0.457 0.744 0.876 0.904 0.912 0.917 0.92 0.923 0.923 0.923
AAE-Cox 0.527 0.648 0.668 0.553 0.394 0.511 0.51 0.476 0.11 0.179 0.925 0.926
SAWAR 0.593 0.684 0.866 0.918 0.922 0.925 0.926 0.927 0.927 0.927 0.927 0.927
prostate DRAFT 0.402 0.416 0.428 0.444 0.469 0.505 0.542 0.585 0.627 0.653 0.661 0.668
Noise 0.433 0.44 0.448 0.451 0.463 0.48 0.492 0.511 0.537 0.558 0.561 0.562
FGSM 0.448 0.453 0.46 0.465 0.475 0.486 0.502 0.515 0.541 0.564 0.569 0.571
PGD 0.447 0.455 0.459 0.467 0.474 0.487 0.509 0.53 0.564 0.585 0.588 0.588
AAE-Cox 0.41 0.411 0.406 0.41 0.407 0.403 0.399 0.391 0.414 0.646 0.686 0.691
SAWAR 0.608 0.623 0.637 0.652 0.665 0.671 0.672 0.673 0.667 0.663 0.66 0.657
retinopathy DRAFT 0.553 0.568 0.578 0.597 0.616 0.632 0.645 0.653 0.663 0.666 0.667 0.669
Noise 0.573 0.591 0.605 0.62 0.632 0.644 0.653 0.661 0.666 0.667 0.668 0.668
FGSM 0.575 0.592 0.604 0.615 0.625 0.633 0.642 0.647 0.651 0.656 0.657 0.659
PGD 0.571 0.589 0.599 0.61 0.621 0.63 0.64 0.647 0.651 0.655 0.656 0.657
AAE-Cox 0.494 0.495 0.506 0.504 0.514 0.564 0.577 0.592 0.62 0.657 0.652 0.648
SAWAR 0.668 0.669 0.67 0.666 0.662 0.66 0.656 0.654 0.653 0.65 0.648 0.647
stagec DRAFT 0.358 0.378 0.382 0.406 0.425 0.449 0.454 0.475 0.489 0.504 0.507 0.512
Noise 0.353 0.373 0.39 0.407 0.436 0.466 0.485 0.5 0.53 0.544 0.549 0.555
FGSM 0.329 0.346 0.368 0.389 0.416 0.429 0.443 0.471 0.488 0.502 0.503 0.51
PGD 0.341 0.352 0.381 0.399 0.419 0.431 0.442 0.47 0.49 0.496 0.498 0.505
AAE-Cox 0.393 0.397 0.397 0.411 0.417 0.409 0.405 0.429 0.469 0.523 0.543 0.54
SAWAR 0.393 0.401 0.413 0.425 0.436 0.461 0.488 0.506 0.534 0.543 0.548 0.558
zinc DRAFT 0.262 0.27 0.284 0.296 0.317 0.355 0.44 0.579 0.712 0.755 0.765 0.77
Noise 0.318 0.328 0.343 0.366 0.401 0.466 0.562 0.661 0.734 0.766 0.773 0.776
FGSM 0.377 0.403 0.439 0.49 0.557 0.626 0.685 0.737 0.762 0.78 0.782 0.783
PGD 0.384 0.406 0.443 0.495 0.559 0.632 0.695 0.741 0.77 0.778 0.78 0.781
AAE-Cox 0.226 0.231 0.228 0.236 0.234 0.249 0.265 0.306 0.445 0.724 0.754 0.759
SAWAR 0.667 0.725 0.763 0.785 0.785 0.785 0.781 0.778 0.778 0.776 0.776 0.774
TABLE V: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
%$\%\Delta$37.55 -42.51 -48.08 -53.15 -55.48 -55.34 -51.51 -42.94 -28.8 -10.87 -4.78 -0.65
TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the worst-case adversarial attack. A lower relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Dataset Algorithm
Aids2 DRAFT 0.265 0.262 0.258 0.252 0.242 0.228 0.207 0.18 0.151 0.138 0.137 0.137
Noise 0.266 0.264 0.261 0.255 0.246 0.231 0.209 0.179 0.15 0.138 0.137 0.138
FGSM 0.265 0.263 0.258 0.251 0.239 0.221 0.197 0.167 0.144 0.137 0.137 0.138
PGD 0.265 0.262 0.257 0.248 0.235 0.216 0.19 0.161 0.141 0.137 0.138 0.138
AAE-Cox 0.269 0.269 0.269 0.268 0.265 0.259 0.25 0.231 0.183 0.138 0.137 0.137
SAWAR 0.14 0.139 0.138 0.138 0.137 0.137 0.137 0.137 0.137 0.137 0.137 0.137
Framingham DRAFT 0.831 0.825 0.811 0.783 0.724 0.618 0.459 0.289 0.174 0.124 0.114 0.11
Noise 0.836 0.834 0.83 0.82 0.79 0.709 0.543 0.331 0.185 0.126 0.115 0.11
FGSM 0.836 0.836 0.835 0.831 0.817 0.765 0.605 0.336 0.162 0.118 0.113 0.112
PGD 0.836 0.836 0.836 0.831 0.812 0.74 0.555 0.295 0.146 0.115 0.113 0.112
AAE-Cox 0.836 0.836 0.836 0.836 0.836 0.836 0.834 0.775 0.385 0.116 0.109 0.108
SAWAR 0.17 0.155 0.144 0.134 0.127 0.122 0.117 0.114 0.112 0.111 0.111 0.111
LeukSurv DRAFT 0.206 0.206 0.205 0.205 0.203 0.2 0.195 0.185 0.17 0.158 0.154 0.153
Noise 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.2 0.171 0.166 0.173
FGSM 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.203 0.187 0.161 0.158 0.159
PGD 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.202 0.183 0.16 0.157 0.158
AAE-Cox 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.198 0.162 0.154 0.152
SAWAR 0.194 0.19 0.184 0.178 0.171 0.165 0.16 0.155 0.151 0.148 0.147 0.146
TRACE DRAFT 0.62 0.611 0.593 0.558 0.504 0.432 0.347 0.265 0.204 0.172 0.165 0.162
Noise 0.627 0.626 0.623 0.611 0.579 0.518 0.426 0.317 0.225 0.176 0.166 0.162
FGSM 0.627 0.625 0.619 0.596 0.545 0.46 0.351 0.248 0.188 0.167 0.163 0.162
PGD 0.627 0.625 0.617 0.593 0.54 0.455 0.347 0.247 0.188 0.167 0.164 0.162
AAE-Cox 0.627 0.627 0.627 0.627 0.627 0.623 0.588 0.471 0.284 0.172 0.165 0.163
SAWAR 0.324 0.281 0.246 0.22 0.202 0.188 0.179 0.171 0.166 0.163 0.162 0.161
dataDIVAT1 DRAFT 0.702 0.689 0.663 0.614 0.527 0.405 0.285 0.21 0.177 0.17 0.172 0.177
Noise 0.72 0.72 0.719 0.718 0.713 0.689 0.599 0.4 0.245 0.197 0.193 0.195
FGSM 0.719 0.719 0.718 0.714 0.697 0.637 0.478 0.285 0.2 0.182 0.182 0.184
PGD 0.719 0.719 0.716 0.71 0.687 0.613 0.442 0.261 0.191 0.18 0.181 0.182
AAE-Cox 0.72 0.72 0.72 0.72 0.72 0.72 0.715 0.553 0.194 0.165 0.17 0.174
SAWAR 0.189 0.184 0.181 0.179 0.177 0.177 0.177 0.177 0.177 0.178 0.179 0.179
flchain DRAFT 0.816 0.816 0.816 0.816 0.816 0.816 0.814 0.797 0.46 0.093 0.069 0.054
Noise 0.816 0.816 0.816 0.816 0.816 0.816 0.814 0.77 0.227 0.088 0.067 0.057
FGSM 0.816 0.816 0.816 0.816 0.816 0.804 0.468 0.119 0.089 0.061 0.055 0.054
PGD 0.817 0.817 0.817 0.816 0.81 0.638 0.162 0.107 0.076 0.057 0.054 0.053
AAE-Cox nan nan nan nan nan nan nan nan 0.794 0.688 0.054 0.051
SAWAR 0.445 0.363 0.241 0.094 0.067 0.059 0.057 0.055 0.054 0.053 0.053 0.053
prostate DRAFT 0.517 0.517 0.516 0.513 0.508 0.495 0.466 0.408 0.322 0.234 0.202 0.181
Noise 0.518 0.518 0.519 0.519 0.52 0.52 0.519 0.508 0.457 0.321 0.267 0.249
FGSM 0.517 0.518 0.519 0.518 0.52 0.52 0.517 0.493 0.402 0.262 0.228 0.218
PGD 0.518 0.518 0.518 0.518 0.517 0.518 0.515 0.486 0.383 0.248 0.219 0.209
AAE-Cox 0.519 0.519 0.519 0.519 0.518 0.518 0.517 0.511 0.443 0.205 0.173 0.169
SAWAR 0.37 0.334 0.291 0.262 0.236 0.218 0.204 0.192 0.184 0.178 0.176 0.175
retinopathy DRAFT 0.728 0.722 0.71 0.687 0.647 0.579 0.48 0.364 0.266 0.204 0.188 0.179
Noise 0.73 0.725 0.714 0.693 0.652 0.579 0.472 0.353 0.256 0.199 0.184 0.177
FGSM 0.725 0.715 0.697 0.665 0.61 0.526 0.419 0.314 0.237 0.196 0.186 0.182
PGD 0.724 0.714 0.695 0.662 0.606 0.521 0.413 0.309 0.235 0.195 0.186 0.182
AAE-Cox 0.733 0.733 0.733 0.733 0.733 0.732 0.716 0.596 0.281 0.183 0.181 0.181
SAWAR 0.588 0.548 0.499 0.44 0.378 0.317 0.265 0.224 0.197 0.184 0.181 0.181
stagec DRAFT 0.556 0.549 0.544 0.543 0.539 0.517 0.468 0.404 0.332 0.274 0.256 0.245
Noise 0.559 0.553 0.547 0.548 0.551 0.541 0.501 0.437 0.357 0.282 0.258 0.244
FGSM 0.547 0.545 0.549 0.553 0.545 0.511 0.457 0.391 0.323 0.278 0.265 0.258
PGD 0.547 0.546 0.549 0.553 0.542 0.507 0.453 0.388 0.322 0.278 0.267 0.26
AAE-Cox 0.568 0.568 0.568 0.568 0.567 0.559 0.545 0.544 0.474 0.291 0.268 0.263
SAWAR 0.521 0.497 0.469 0.436 0.401 0.365 0.334 0.303 0.276 0.256 0.249 0.243
zinc DRAFT 0.847 0.847 0.846 0.843 0.831 0.787 0.641 0.376 0.184 0.119 0.109 0.107
Noise 0.847 0.847 0.847 0.845 0.838 0.795 0.627 0.351 0.179 0.128 0.118 0.113
FGSM 0.847 0.847 0.845 0.835 0.782 0.611 0.356 0.189 0.133 0.116 0.113 0.112
PGD 0.847 0.847 0.845 0.834 0.777 0.6 0.341 0.182 0.132 0.116 0.113 0.112
AAE-Cox 0.847 0.847 0.847 0.847 0.847 0.847 0.846 0.823 0.507 0.119 0.109 0.108
SAWAR 0.654 0.532 0.401 0.289 0.21 0.161 0.133 0.118 0.111 0.109 0.109 0.109
TABLE VII: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Dataset Algorithm
Aids2 DRAFT 7.88e+05 2.75e+05 9.79e+04 3.51e+04 1.31e+04 5.06e+03 2.06e+03 9.59e+02 6.16e+02 5.45e+02 5.41e+02 5.41e+02
Noise 3.15e+05 1.30e+05 5.45e+04 2.30e+04 9.53e+03 4.06e+03 1.79e+03 9.02e+02 6.06e+02 5.44e+02 5.41e+02 5.41e+02
FGSM 1.29e+05 5.99e+04 2.79e+04 1.29e+04 5.92e+03 2.72e+03 1.32e+03 7.49e+02 5.69e+02 5.42e+02 5.41e+02 5.41e+02
PGD 7.29e+04 3.62e+04 1.80e+04 8.83e+03 4.29e+03 2.10e+03 1.09e+03 6.82e+02 5.57e+02 5.42e+02 5.41e+02 5.41e+02
AAE-Cox 6.04e+12 2.43e+11 1.30e+10 6.94e+08 4.08e+07 2.64e+06 1.76e+05 1.27e+04 1.18e+03 5.41e+02 5.38e+02 5.39e+02
SAWAR 5.53e+02 5.48e+02 5.45e+02 5.43e+02 5.42e+02 5.42e+02 5.41e+02 5.41e+02 5.40e+02 5.40e+02 5.40e+02 5.40e+02
Framingham DRAFT 2.24e+07 4.61e+06 9.70e+05 2.13e+05 4.98e+04 1.33e+04 4.62e+03 2.35e+03 1.71e+03 1.52e+03 1.49e+03 1.48e+03
Noise 3.13e+07 6.71e+06 1.48e+06 3.35e+05 7.98e+04 2.05e+04 6.23e+03 2.67e+03 1.76e+03 1.53e+03 1.49e+03 1.48e+03
FGSM 8.34e+09 7.44e+08 6.79e+07 6.46e+06 6.50e+05 7.31e+04 1.05e+04 2.77e+03 1.67e+03 1.50e+03 1.49e+03 1.48e+03
PGD 1.53e+11 7.60e+09 3.88e+08 2.07e+07 1.20e+06 8.59e+04 9.43e+03 2.46e+03 1.61e+03 1.49e+03 1.48e+03 1.48e+03
AAE-Cox 1.38e+26 1.22e+23 1.20e+20 1.10e+17 1.21e+14 1.26e+11 1.63e+08 3.73e+05 3.68e+03 1.49e+03 1.47e+03 1.47e+03
SAWAR 1.67e+03 1.62e+03 1.58e+03 1.55e+03 1.53e+03 1.51e+03 1.50e+03 1.49e+03 1.48e+03 1.48e+03 1.48e+03 1.48e+03
LeukSurv DRAFT 2.10e+08 3.80e+07 7.18e+06 1.29e+06 2.41e+05 4.59e+04 9.21e+03 2.13e+03 6.86e+02 3.75e+02 3.28e+02 3.07e+02
Noise 1.40e+20 1.48e+18 1.45e+16 1.63e+14 1.29e+12 9.53e+09 8.92e+07 1.13e+06 2.90e+04 2.62e+03 1.25e+03 7.75e+02
FGSM 6.24e+12 4.01e+11 2.30e+10 1.27e+09 7.06e+07 4.34e+06 2.73e+05 2.04e+04 2.18e+03 5.11e+02 3.72e+02 3.25e+02
PGD 4.11e+11 3.30e+10 2.53e+09 1.99e+08 1.58e+07 1.30e+06 1.12e+05 1.10e+04 1.50e+03 4.34e+02 3.39e+02 3.07e+02
AAE-Cox 4.24e+28 2.71e+25 1.61e+34 7.46e+29 2.45e+25 7.92e+19 7.26e+14 3.55e+09 1.15e+05 4.18e+02 2.99e+02 2.84e+02
SAWAR 3.30e+03 2.00e+03 1.23e+03 7.95e+02 5.43e+02 3.97e+02 3.38e+02 3.06e+02 2.87e+02 2.75e+02 2.72e+02 2.70e+02
TRACE DRAFT 5.09e+05 1.75e+05 6.15e+04 2.23e+04 8.34e+03 3.31e+03 1.48e+03 8.34e+02 6.09e+02 5.38e+02 5.27e+02 5.24e+02
Noise 2.14e+09 2.12e+08 2.24e+07 2.58e+06 3.25e+05 4.62e+04 7.86e+03 1.89e+03 7.98e+02 5.76e+02 5.46e+02 5.35e+02
FGSM 4.68e+08 4.01e+07 3.85e+06 4.31e+05 5.82e+04 9.91e+03 2.30e+03 8.91e+02 6.03e+02 5.39e+02 5.31e+02 5.28e+02
PGD 4.42e+07 6.51e+06 1.04e+06 1.82e+05 3.49e+04 7.47e+03 2.01e+03 8.54e+02 5.99e+02 5.39e+02 5.31e+02 5.27e+02
AAE-Cox 1.06e+24 1.53e+21 2.08e+18 2.78e+15 2.40e+12 2.92e+09 5.29e+06 2.86e+04 9.75e+02 5.41e+02 5.31e+02 5.29e+02
SAWAR 1.49e+03 1.01e+03 7.74e+02 6.57e+02 5.98e+02 5.67e+02 5.48e+02 5.35e+02 5.28e+02 5.24e+02 5.24e+02 5.24e+02
dataDIVAT1 DRAFT 2.97e+04 1.56e+04 8.36e+03 4.56e+03 2.59e+03 1.59e+03 1.11e+03 8.83e+02 7.91e+02 7.56e+02 7.51e+02 7.50e+02
Noise 1.60e+10 1.12e+09 8.30e+07 6.62e+06 5.83e+05 5.99e+04 8.13e+03 1.87e+03 9.38e+02 7.78e+02 7.61e+02 7.59e+02
FGSM 3.25e+08 3.59e+07 4.13e+06 5.10e+05 7.07e+04 1.17e+04 2.66e+03 1.10e+03 8.11e+02 7.58e+02 7.53e+02 7.52e+02
PGD 1.22e+08 1.55e+07 2.06e+06 2.92e+05 4.56e+04 8.39e+03 2.15e+03 1.01e+03 7.94e+02 7.55e+02 7.52e+02 7.51e+02
AAE-Cox 9.96e+18 6.79e+16 3.12e+14 1.88e+12 1.37e+10 1.14e+08 1.05e+06 1.57e+04 1.04e+03 7.50e+02 7.45e+02 7.45e+02
SAWAR 7.89e+02 7.79e+02 7.70e+02 7.63e+02 7.58e+02 7.54e+02 7.51e+02 7.49e+02 7.47e+02 7.46e+02 7.46e+02 7.46e+02
flchain DRAFT 1.57e+22 5.59e+19 2.03e+17 7.97e+14 3.37e+12 1.52e+10 7.49e+07 5.25e+05 1.45e+04 1.95e+03 1.25e+03 1.10e+03
Noise 2.69e+32 2.90e+28 3.20e+24 5.94e+33 2.08e+27 7.50e+20 3.25e+14 2.31e+08 6.25e+04 2.40e+03 1.47e+03 1.24e+03
FGSM 5.19e+19 7.59e+16 1.21e+14 2.16e+11 7.32e+08 9.31e+06 2.85e+05 1.58e+04 2.16e+03 1.18e+03 1.12e+03 1.10e+03
PGD 3.10e+15 1.18e+13 6.11e+10 6.20e+08 1.63e+07 6.61e+05 3.81e+04 4.41e+03 1.43e+03 1.12e+03 1.09e+03 1.09e+03
AAE-Cox nan nan nan nan nan nan nan nan 4.02e+33 2.23e+10 1.11e+03 1.08e+03
SAWAR 5.84e+05 3.37e+04 4.04e+03 1.74e+03 1.31e+03 1.18e+03 1.13e+03 1.11e+03 1.10e+03 1.09e+03 1.09e+03 1.09e+03
prostate DRAFT 5.92e+05 2.18e+05 8.11e+04 3.04e+04 1.16e+04 4.49e+03 1.83e+03 8.39e+02 4.86e+02 3.71e+02 3.48e+02 3.37e+02
Noise 6.55e+23 1.63e+21 4.14e+18 1.02e+16 2.79e+13 8.22e+10 2.79e+08 1.36e+06 3.02e+04 3.29e+03 1.53e+03 9.01e+02
FGSM 5.24e+15 9.63e+13 1.84e+12 3.69e+10 7.62e+08 1.69e+07 4.35e+05 1.56e+04 1.19e+03 4.22e+02 3.76e+02 3.63e+02
PGD 1.20e+13 4.77e+11 1.96e+10 8.28e+08 3.66e+07 1.75e+06 9.45e+04 6.42e+03 8.39e+02 3.99e+02 3.67e+02 3.57e+02
AAE-Cox 1.53e+20 9.61e+17 6.16e+15 4.00e+13 2.65e+11 1.87e+09 1.33e+07 1.16e+05 1.81e+03 3.47e+02 3.33e+02 3.31e+02
SAWAR 6.21e+02 5.06e+02 4.28e+02 3.91e+02 3.67e+02 3.55e+02 3.47e+02 3.41e+02 3.37e+02 3.35e+02 3.34e+02 3.33e+02
retinopathy DRAFT 1.21e+04 6.19e+03 3.21e+03 1.68e+03 9.15e+02 5.24e+02 3.31e+02 2.39e+02 1.96e+02 1.76e+02 1.71e+02 1.69e+02
Noise 1.65e+04 8.10e+03 4.01e+03 2.03e+03 1.05e+03 5.83e+02 3.54e+02 2.47e+02 1.99e+02 1.77e+02 1.72e+02 1.69e+02
FGSM 7.62e+03 4.03e+03 2.15e+03 1.18e+03 6.72e+02 4.10e+02 2.80e+02 2.16e+02 1.86e+02 1.73e+02 1.70e+02 1.69e+02
PGD 7.32e+03 3.87e+03 2.08e+03 1.14e+03 6.55e+02 4.03e+02 2.75e+02 2.14e+02 1.85e+02 1.73e+02 1.70e+02 1.69e+02
AAE-Cox 1.32e+13 3.68e+11 1.03e+10 2.98e+08 8.67e+06 2.82e+05 1.17e+04 8.14e+02 1.98e+02 1.68e+02 1.67e+02 1.67e+02
SAWAR 6.10e+02 4.69e+02 3.69e+02 3.00e+02 2.51e+02 2.18e+02 1.96e+02 1.82e+02 1.74e+02 1.69e+02 1.68e+02 1.68e+02
stagec DRAFT 1.17e+04 5.05e+03 2.20e+03 9.79e+02 4.44e+02 2.10e+02 1.07e+02 6.47e+01 4.80e+01 4.24e+01 4.15e+01 4.13e+01
Noise 6.21e+04 2.17e+04 7.71e+03 2.72e+03 1.01e+03 3.86e+02 1.63e+02 8.11e+01 5.29e+01 4.44e+01 4.31e+01 4.27e+01
FGSM 1.02e+04 4.31e+03 1.80e+03 7.90e+02 3.54e+02 1.69e+02 9.18e+01 5.99e+01 4.77e+01 4.40e+01 4.35e+01 4.36e+01
PGD 8.55e+03 3.69e+03 1.60e+03 7.18e+02 3.32e+02 1.62e+02 8.89e+01 5.88e+01 4.73e+01 4.38e+01 4.34e+01 4.35e+01
AAE-Cox 1.32e+20 6.14e+17 2.58e+15 1.21e+13 5.37e+10 2.81e+08 1.53e+06 1.01e+04 1.53e+02 4.20e+01 4.10e+01 4.09e+01
SAWAR 2.52e+02 1.75e+02 1.26e+02 9.40e+01 7.32e+01 5.96e+01 5.14e+01 4.61e+01 4.28e+01 4.09e+01 4.04e+01 4.01e+01
zinc DRAFT 1.34e+06 3.43e+05 8.95e+04 2.36e+04 6.49e+03 1.84e+03 5.72e+02 2.12e+02 1.11e+02 8.33e+01 7.84e+01 7.64e+01
Noise 1.01e+07 2.02e+06 4.01e+05 8.52e+04 1.86e+04 4.26e+03 1.05e+03 3.06e+02 1.28e+02 8.63e+01 8.05e+01 7.86e+01
FGSM 3.49e+05 9.39e+04 2.69e+04 7.98e+03 2.42e+03 8.00e+02 3.01e+02 1.41e+02 9.24e+01 7.95e+01 7.82e+01 7.82e+01
PGD 4.28e+05 1.09e+05 2.89e+04 7.84e+03 2.37e+03 7.72e+02 2.87e+02 1.36e+02 9.09e+01 7.94e+01 7.83e+01 7.85e+01
AAE-Cox 1.15e+24 1.54e+21 4.97e+18 8.90e+15 1.88e+13 2.84e+10 5.64e+07 1.63e+05 7.39e+02 8.21e+01 7.79e+01 7.76e+01
SAWAR 5.06e+02 3.35e+02 2.31e+02 1.67e+02 1.28e+02 1.04e+02 9.02e+01 8.23e+01 7.82e+01 7.68e+01 7.68e+01 7.72e+01
TABLE VIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the FGSM adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Concordance Index metric from the DRAFT training objective to SAWAR training objective is shown in Table IX (where higher percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
%$\%\Delta$ 39.82 59.68 72.41 86.66 105.06 136.98 168.31 178.41 116.94 26.54 12.48 1.6
TABLE IX: The relative percent change in Concordance Index metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the FGSM adversarial attack. A lower relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Dataset Algorithm
Aids2 DRAFT 0.232 0.233 0.235 0.238 0.243 0.25 0.259 0.274 0.3 0.376 0.458 0.57
Noise 0.23 0.231 0.233 0.236 0.24 0.249 0.26 0.277 0.306 0.386 0.465 0.569
FGSM 0.356 0.363 0.375 0.389 0.403 0.421 0.439 0.459 0.477 0.508 0.531 0.566
PGD 0.326 0.336 0.35 0.367 0.385 0.409 0.434 0.46 0.486 0.515 0.534 0.567
AAE-Cox 0.241 0.246 0.253 0.261 0.271 0.282 0.299 0.332 0.39 0.474 0.523 0.573
SAWAR 0.259 0.265 0.276 0.289 0.306 0.327 0.352 0.385 0.43 0.493 0.53 0.569
Framingham DRAFT 0.142 0.143 0.143 0.144 0.144 0.145 0.148 0.168 0.257 0.468 0.6 0.722
Noise 0.143 0.144 0.144 0.145 0.145 0.146 0.15 0.173 0.265 0.473 0.601 0.719
FGSM 0.149 0.152 0.156 0.164 0.18 0.208 0.257 0.332 0.437 0.57 0.643 0.715
PGD 0.181 0.189 0.2 0.218 0.245 0.285 0.34 0.411 0.5 0.603 0.66 0.717
AAE-Cox 0.148 0.152 0.162 0.181 0.216 0.269 0.34 0.427 0.53 0.636 0.687 0.733
SAWAR 0.149 0.158 0.178 0.213 0.268 0.34 0.417 0.5 0.584 0.665 0.702 0.737
LeukSurv DRAFT 0.378 0.38 0.381 0.383 0.387 0.394 0.407 0.432 0.475 0.544 0.587 0.633
Noise 0.39 0.393 0.397 0.401 0.406 0.411 0.417 0.428 0.453 0.507 0.556 0.624
FGSM 0.397 0.401 0.405 0.41 0.418 0.427 0.443 0.466 0.499 0.554 0.59 0.635
PGD 0.399 0.402 0.407 0.411 0.419 0.431 0.447 0.473 0.508 0.562 0.597 0.638
AAE-Cox 0.376 0.377 0.379 0.382 0.388 0.4 0.42 0.453 0.506 0.575 0.615 0.656
SAWAR 0.392 0.399 0.409 0.423 0.442 0.465 0.494 0.529 0.574 0.623 0.649 0.673
TRACE DRAFT 0.219 0.221 0.225 0.232 0.247 0.279 0.334 0.415 0.524 0.642 0.696 0.745
Noise 0.234 0.236 0.24 0.246 0.259 0.286 0.336 0.414 0.52 0.639 0.695 0.745
FGSM 0.29 0.311 0.336 0.369 0.408 0.454 0.507 0.565 0.625 0.685 0.715 0.744
PGD 0.304 0.326 0.353 0.388 0.428 0.475 0.525 0.58 0.635 0.691 0.718 0.745
AAE-Cox 0.251 0.273 0.302 0.339 0.384 0.438 0.498 0.563 0.63 0.692 0.721 0.747
SAWAR 0.287 0.32 0.363 0.411 0.464 0.519 0.57 0.62 0.665 0.708 0.729 0.747
dataDIVAT1 DRAFT 0.097 0.098 0.099 0.1 0.101 0.103 0.105 0.122 0.214 0.414 0.53 0.655
Noise 0.094 0.094 0.095 0.096 0.097 0.099 0.101 0.111 0.178 0.366 0.494 0.635
FGSM 0.114 0.116 0.12 0.126 0.138 0.161 0.203 0.27 0.362 0.482 0.559 0.646
PGD 0.137 0.143 0.152 0.167 0.188 0.22 0.265 0.33 0.409 0.511 0.574 0.648
AAE-Cox 0.111 0.114 0.122 0.139 0.167 0.209 0.268 0.347 0.443 0.551 0.607 0.663
SAWAR 0.156 0.201 0.253 0.311 0.368 0.422 0.461 0.506 0.555 0.609 0.637 0.67
flchain DRAFT 0.152 0.153 0.154 0.155 0.156 0.157 0.159 0.164 0.197 0.895 0.903 0.921
Noise 0.166 0.167 0.169 0.171 0.175 0.182 0.213 0.333 0.809 0.9 0.904 0.918
FGSM 0.507 0.628 0.727 0.8 0.854 0.883 0.898 0.902 0.906 0.911 0.915 0.922
PGD 0.539 0.717 0.824 0.876 0.896 0.9 0.903 0.905 0.908 0.913 0.917 0.923
AAE-Cox 0.172 0.188 0.218 0.275 0.372 0.508 0.665 0.81 0.899 0.918 0.922 0.926
SAWAR 0.628 0.831 0.894 0.9 0.904 0.907 0.909 0.912 0.916 0.921 0.924 0.927
prostate DRAFT 0.308 0.311 0.314 0.318 0.321 0.326 0.333 0.339 0.357 0.435 0.531 0.668
Noise 0.305 0.305 0.306 0.307 0.309 0.314 0.32 0.328 0.344 0.394 0.457 0.562
FGSM 0.293 0.295 0.297 0.3 0.304 0.308 0.312 0.32 0.343 0.407 0.474 0.571
PGD 0.299 0.301 0.303 0.305 0.308 0.312 0.318 0.331 0.36 0.431 0.502 0.588
AAE-Cox 0.308 0.308 0.311 0.315 0.325 0.342 0.375 0.426 0.499 0.597 0.645 0.691
SAWAR 0.288 0.292 0.299 0.31 0.326 0.349 0.384 0.432 0.499 0.579 0.618 0.657
retinopathy DRAFT 0.139 0.141 0.145 0.15 0.152 0.152 0.153 0.158 0.209 0.425 0.554 0.669
Noise 0.134 0.137 0.138 0.138 0.138 0.138 0.138 0.15 0.245 0.456 0.57 0.668
FGSM 0.131 0.133 0.135 0.135 0.135 0.135 0.137 0.155 0.255 0.456 0.561 0.659
PGD 0.13 0.13 0.131 0.131 0.131 0.131 0.134 0.154 0.254 0.456 0.56 0.657
AAE-Cox 0.129 0.129 0.13 0.137 0.157 0.196 0.247 0.337 0.444 0.541 0.595 0.648
SAWAR 0.131 0.13 0.13 0.131 0.135 0.149 0.178 0.251 0.371 0.51 0.584 0.647
stagec DRAFT 0.137 0.136 0.136 0.132 0.136 0.135 0.151 0.174 0.23 0.34 0.407 0.512
Noise 0.133 0.133 0.132 0.132 0.136 0.144 0.16 0.187 0.261 0.362 0.442 0.555
FGSM 0.12 0.124 0.127 0.134 0.146 0.147 0.157 0.184 0.255 0.33 0.417 0.51
PGD 0.115 0.117 0.122 0.128 0.141 0.144 0.154 0.181 0.259 0.327 0.412 0.505
AAE-Cox 0.125 0.121 0.131 0.141 0.163 0.202 0.241 0.296 0.352 0.426 0.485 0.54
SAWAR 0.123 0.124 0.127 0.139 0.149 0.179 0.239 0.281 0.347 0.425 0.486 0.558
zinc DRAFT 0.057 0.057 0.057 0.057 0.057 0.057 0.057 0.068 0.155 0.485 0.646 0.77
Noise 0.053 0.053 0.054 0.053 0.054 0.057 0.062 0.095 0.248 0.547 0.668 0.776
FGSM 0.055 0.056 0.056 0.058 0.06 0.07 0.095 0.199 0.404 0.61 0.705 0.783
PGD 0.056 0.056 0.057 0.058 0.064 0.078 0.113 0.22 0.42 0.615 0.706 0.781
AAE-Cox 0.062 0.065 0.079 0.101 0.162 0.227 0.291 0.409 0.551 0.655 0.708 0.759
SAWAR 0.055 0.056 0.058 0.073 0.111 0.213 0.325 0.446 0.556 0.679 0.729 0.774
TABLE X: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the FGSM adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the FGSM adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Scores metric from the DRAFT training objective to SAWAR training objective is shown in Table XI (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
%$\%\Delta$44.57 -46.37 -47.93 -49.06 -49.43 -48.69 -46.35 -41.82 -33.56 -20.63 -11.37 -0.65
TABLE XI: The relative percent change in Integrated Brier Score metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the FGSM adversarial attack. A lower relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Dataset Algorithm
Aids2 DRAFT 0.232 0.233 0.235 0.238 0.243 0.25 0.259 0.274 0.3 0.376 0.458 0.57
Noise 0.23 0.231 0.233 0.236 0.24 0.249 0.26 0.277 0.306 0.386 0.465 0.569
FGSM 0.356 0.363 0.375 0.389 0.403 0.421 0.439 0.459 0.477 0.508 0.531 0.566
PGD 0.326 0.336 0.35 0.367 0.385 0.409 0.434 0.46 0.486 0.515 0.534 0.567
AAE-Cox 0.241 0.246 0.253 0.261 0.271 0.282 0.299 0.332 0.39 0.474 0.523 0.573
SAWAR 0.259 0.265 0.276 0.289 0.306 0.327 0.352 0.385 0.43 0.493 0.53 0.569
Framingham DRAFT 0.142 0.143 0.143 0.144 0.144 0.145 0.148 0.168 0.257 0.468 0.6 0.722
Noise 0.143 0.144 0.144 0.145 0.145 0.146 0.15 0.173 0.265 0.473 0.601 0.719
FGSM 0.149 0.152 0.156 0.164 0.18 0.208 0.257 0.332 0.437 0.57 0.643 0.715
PGD 0.181 0.189 0.2 0.218 0.245 0.285 0.34 0.411 0.5 0.603 0.66 0.717
AAE-Cox 0.148 0.152 0.162 0.181 0.216 0.269 0.34 0.427 0.53 0.636 0.687 0.733
SAWAR 0.149 0.158 0.178 0.213 0.268 0.34 0.417 0.5 0.584 0.665 0.702 0.737
LeukSurv DRAFT 0.378 0.38 0.381 0.383 0.387 0.394 0.407 0.432 0.475 0.544 0.587 0.633
Noise 0.39 0.393 0.397 0.401 0.406 0.411 0.417 0.428 0.453 0.507 0.556 0.624
FGSM 0.397 0.401 0.405 0.41 0.418 0.427 0.443 0.466 0.499 0.554 0.59 0.635
PGD 0.399 0.402 0.407 0.411 0.419 0.431 0.447 0.473 0.508 0.562 0.597 0.638
AAE-Cox 0.376 0.377 0.379 0.382 0.388 0.4 0.42 0.453 0.506 0.575 0.615 0.656
SAWAR 0.392 0.399 0.409 0.423 0.442 0.465 0.494 0.529 0.574 0.623 0.649 0.673
TRACE DRAFT 0.219 0.221 0.225 0.232 0.247 0.279 0.334 0.415 0.524 0.642 0.696 0.745
Noise 0.234 0.236 0.24 0.246 0.259 0.286 0.336 0.414 0.52 0.639 0.695 0.745
FGSM 0.29 0.311 0.336 0.369 0.408 0.454 0.507 0.565 0.625 0.685 0.715 0.744
PGD 0.304 0.326 0.353 0.388 0.428 0.475 0.525 0.58 0.635 0.691 0.718 0.745
AAE-Cox 0.251 0.273 0.302 0.339 0.384 0.438 0.498 0.563 0.63 0.692 0.721 0.747
SAWAR 0.287 0.32 0.363 0.411 0.464 0.519 0.57 0.62 0.665 0.708 0.729 0.747
dataDIVAT1 DRAFT 0.097 0.098 0.099 0.1 0.101 0.103 0.105 0.122 0.214 0.414 0.53 0.655
Noise 0.094 0.094 0.095 0.096 0.097 0.099 0.101 0.111 0.178 0.366 0.494 0.635
FGSM 0.114 0.116 0.12 0.126 0.138 0.161 0.203 0.27 0.362 0.482 0.559 0.646
PGD 0.137 0.143 0.152 0.167 0.188 0.22 0.265 0.33 0.409 0.511 0.574 0.648
AAE-Cox 0.111 0.114 0.122 0.139 0.167 0.209 0.268 0.347 0.443 0.551 0.607 0.663
SAWAR 0.156 0.201 0.253 0.311 0.368 0.422 0.461 0.506 0.555 0.609 0.637 0.67
flchain DRAFT 0.152 0.153 0.154 0.155 0.156 0.157 0.159 0.164 0.197 0.895 0.903 0.921
Noise 0.166 0.167 0.169 0.171 0.175 0.182 0.213 0.333 0.809 0.9 0.904 0.918
FGSM 0.507 0.628 0.727 0.8 0.854 0.883 0.898 0.902 0.906 0.911 0.915 0.922
PGD 0.539 0.717 0.824 0.876 0.896 0.9 0.903 0.905 0.908 0.913 0.917 0.923
AAE-Cox 0.172 0.188 0.218 0.275 0.372 0.508 0.665 0.81 0.899 0.918 0.922 0.926
SAWAR 0.628 0.831 0.894 0.9 0.904 0.907 0.909 0.912 0.916 0.921 0.924 0.927
prostate DRAFT 0.308 0.311 0.314 0.318 0.321 0.326 0.333 0.339 0.357 0.435 0.531 0.668
Noise 0.305 0.305 0.306 0.307 0.309 0.314 0.32 0.328 0.344 0.394 0.457 0.562
FGSM 0.293 0.295 0.297 0.3 0.304 0.308 0.312 0.32 0.343 0.407 0.474 0.571
PGD 0.299 0.301 0.303 0.305 0.308 0.312 0.318 0.331 0.36 0.431 0.502 0.588
AAE-Cox 0.308 0.308 0.311 0.315 0.325 0.342 0.375 0.426 0.499 0.597 0.645 0.691
SAWAR 0.288 0.292 0.299 0.31 0.326 0.349 0.384 0.432 0.499 0.579 0.618 0.657
retinopathy DRAFT 0.139 0.141 0.145 0.15 0.152 0.152 0.153 0.158 0.209 0.425 0.554 0.669
Noise 0.134 0.137 0.138 0.138 0.138 0.138 0.138 0.15 0.245 0.456 0.57 0.668
FGSM 0.131 0.133 0.135 0.135 0.135 0.135 0.137 0.155 0.255 0.456 0.561 0.659
PGD 0.13 0.13 0.131 0.131 0.131 0.131 0.134 0.154 0.254 0.456 0.56 0.657
AAE-Cox 0.129 0.129 0.13 0.137 0.157 0.196 0.247 0.337 0.444 0.541 0.595 0.648
SAWAR 0.131 0.13 0.13 0.131 0.135 0.149 0.178 0.251 0.371 0.51 0.584 0.647
stagec DRAFT 0.137 0.136 0.136 0.132 0.136 0.135 0.151 0.174 0.23 0.34 0.407 0.512
Noise 0.133 0.133 0.132 0.132 0.136 0.144 0.16 0.187 0.261 0.362 0.442 0.555
FGSM 0.12 0.124 0.127 0.134 0.146 0.147 0.157 0.184 0.255 0.33 0.417 0.51
PGD 0.115 0.117 0.122 0.128 0.141 0.144 0.154 0.181 0.259 0.327 0.412 0.505
AAE-Cox 0.125 0.121 0.131 0.141 0.163 0.202 0.241 0.296 0.352 0.426 0.485 0.54
SAWAR 0.123 0.124 0.127 0.139 0.149 0.179 0.239 0.281 0.347 0.425 0.486 0.558
zinc DRAFT 0.057 0.057 0.057 0.057 0.057 0.057 0.057 0.068 0.155 0.485 0.646 0.77
Noise 0.053 0.053 0.054 0.053 0.054 0.057 0.062 0.095 0.248 0.547 0.668 0.776
FGSM 0.055 0.056 0.056 0.058 0.06 0.07 0.095 0.199 0.404 0.61 0.705 0.783
PGD 0.056 0.056 0.057 0.058 0.064 0.078 0.113 0.22 0.42 0.615 0.706 0.781
AAE-Cox 0.062 0.065 0.079 0.101 0.162 0.227 0.291 0.409 0.551 0.655 0.708 0.759
SAWAR 0.055 0.056 0.058 0.073 0.111 0.213 0.325 0.446 0.556 0.679 0.729 0.774
TABLE XII: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the FGSM adversarial attack.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Dataset Algorithm
Aids2 DRAFT 1.04e+03 9.95e+02 9.48e+02 9.01e+02 8.54e+02 8.06e+02 7.56e+02 7.04e+02 6.49e+02 5.92e+02 5.66e+02 5.41e+02
Noise 9.94e+02 9.53e+02 9.11e+02 8.69e+02 8.26e+02 7.82e+02 7.37e+02 6.89e+02 6.39e+02 5.88e+02 5.64e+02 5.41e+02
FGSM 6.16e+02 6.10e+02 6.04e+02 5.97e+02 5.91e+02 5.84e+02 5.77e+02 5.70e+02 5.63e+02 5.54e+02 5.48e+02 5.41e+02
PGD 6.20e+02 6.13e+02 6.05e+02 5.98e+02 5.91e+02 5.83e+02 5.76e+02 5.68e+02 5.60e+02 5.52e+02 5.47e+02 5.41e+02
AAE-Cox 7.00e+02 6.83e+02 6.67e+02 6.51e+02 6.35e+02 6.19e+02 6.03e+02 5.87e+02 5.71e+02 5.55e+02 5.47e+02 5.39e+02
SAWAR 5.84e+02 5.80e+02 5.76e+02 5.72e+02 5.68e+02 5.63e+02 5.59e+02 5.55e+02 5.50e+02 5.45e+02 5.43e+02 5.40e+02
Framingham DRAFT 4.23e+03 4.10e+03 3.91e+03 3.65e+03 3.33e+03 2.97e+03 2.61e+03 2.26e+03 1.95e+03 1.68e+03 1.57e+03 1.48e+03
Noise 4.00e+03 3.89e+03 3.73e+03 3.51e+03 3.23e+03 2.91e+03 2.57e+03 2.24e+03 1.94e+03 1.69e+03 1.58e+03 1.48e+03
FGSM 2.11e+03 2.07e+03 2.03e+03 1.98e+03 1.93e+03 1.88e+03 1.82e+03 1.75e+03 1.67e+03 1.58e+03 1.53e+03 1.48e+03
PGD 1.99e+03 1.95e+03 1.91e+03 1.87e+03 1.83e+03 1.79e+03 1.74e+03 1.68e+03 1.62e+03 1.56e+03 1.52e+03 1.48e+03
AAE-Cox 2.38e+03 2.27e+03 2.16e+03 2.06e+03 1.96e+03 1.87e+03 1.78e+03 1.70e+03 1.62e+03 1.54e+03 1.51e+03 1.47e+03
SAWAR 1.96e+03 1.90e+03 1.84e+03 1.79e+03 1.73e+03 1.68e+03 1.64e+03 1.59e+03 1.55e+03 1.51e+03 1.49e+03 1.48e+03
LeukSurv DRAFT 2.66e+03 2.16e+03 1.77e+03 1.45e+03 1.19e+03 9.73e+02 7.93e+02 6.41e+02 5.10e+02 3.99e+02 3.50e+02 3.07e+02
Noise 1.13e+06 5.84e+05 2.96e+05 1.52e+05 7.88e+04 4.05e+04 1.94e+04 9.21e+03 4.25e+03 1.89e+03 1.23e+03 7.75e+02
FGSM 2.76e+03 2.33e+03 1.96e+03 1.65e+03 1.39e+03 1.16e+03 9.64e+02 7.82e+02 6.18e+02 4.64e+02 3.92e+02 3.25e+02
PGD 1.79e+03 1.57e+03 1.38e+03 1.21e+03 1.05e+03 9.10e+02 7.76e+02 6.49e+02 5.30e+02 4.15e+02 3.60e+02 3.07e+02
AAE-Cox 1.09e+03 9.83e+02 8.87e+02 7.97e+02 7.12e+02 6.32e+02 5.57e+02 4.84e+02 4.12e+02 3.43e+02 3.12e+02 2.84e+02
SAWAR 5.40e+02 5.12e+02 4.85e+02 4.57e+02 4.29e+02 4.01e+02 3.73e+02 3.46e+02 3.20e+02 2.94e+02 2.82e+02 2.70e+02
TRACE DRAFT 2.31e+03 2.09e+03 1.87e+03 1.64e+03 1.43e+03 1.22e+03 1.03e+03 8.69e+02 7.29e+02 6.14e+02 5.66e+02 5.24e+02
Noise 4.55e+03 3.85e+03 3.20e+03 2.63e+03 2.14e+03 1.72e+03 1.36e+03 1.06e+03 8.29e+02 6.60e+02 5.93e+02 5.35e+02
FGSM 1.23e+03 1.15e+03 1.07e+03 9.98e+02 9.26e+02 8.54e+02 7.82e+02 7.13e+02 6.47e+02 5.85e+02 5.56e+02 5.28e+02
PGD 1.17e+03 1.10e+03 1.03e+03 9.59e+02 8.90e+02 8.21e+02 7.54e+02 6.93e+02 6.35e+02 5.80e+02 5.53e+02 5.27e+02
AAE-Cox 1.07e+03 1.00e+03 9.38e+02 8.77e+02 8.19e+02 7.63e+02 7.09e+02 6.59e+02 6.11e+02 5.68e+02 5.48e+02 5.29e+02
SAWAR 1.01e+03 9.32e+02 8.63e+02 8.01e+02 7.47e+02 6.99e+02 6.57e+02 6.19e+02 5.84e+02 5.53e+02 5.38e+02 5.24e+02
dataDIVAT1 DRAFT 1.78e+03 1.72e+03 1.64e+03 1.55e+03 1.44e+03 1.32e+03 1.20e+03 1.07e+03 9.55e+02 8.46e+02 7.96e+02 7.50e+02
Noise 2.02e+03 1.96e+03 1.88e+03 1.78e+03 1.65e+03 1.50e+03 1.35e+03 1.19e+03 1.03e+03 8.87e+02 8.20e+02 7.59e+02
FGSM 1.09e+03 1.07e+03 1.05e+03 1.02e+03 1.00e+03 9.72e+02 9.40e+02 9.03e+02 8.59e+02 8.09e+02 7.81e+02 7.52e+02
PGD 1.01e+03 9.95e+02 9.78e+02 9.60e+02 9.41e+02 9.19e+02 8.94e+02 8.65e+02 8.32e+02 7.95e+02 7.74e+02 7.51e+02
AAE-Cox 1.25e+03 1.18e+03 1.12e+03 1.06e+03 1.01e+03 9.55e+02 9.09e+02 8.65e+02 8.23e+02 7.83e+02 7.64e+02 7.45e+02
SAWAR 9.24e+02 9.01e+02 8.79e+02 8.59e+02 8.39e+02 8.20e+02 8.04e+02 7.89e+02 7.74e+02 7.60e+02 7.53e+02 7.46e+02
flchain DRAFT 2.56e+04 2.36e+04 2.15e+04 1.95e+04 1.73e+04 1.47e+04 1.11e+04 6.38e+03 2.91e+03 1.59e+03 1.27e+03 1.10e+03
Noise 1.62e+05 1.27e+05 8.24e+04 4.85e+04 2.97e+04 1.81e+04 9.84e+03 5.14e+03 2.95e+03 1.80e+03 1.45e+03 1.24e+03
FGSM 2.49e+03 2.17e+03 1.87e+03 1.62e+03 1.44e+03 1.34e+03 1.28e+03 1.23e+03 1.19e+03 1.14e+03 1.12e+03 1.10e+03
PGD 2.07e+03 1.76e+03 1.53e+03 1.37e+03 1.28e+03 1.22e+03 1.19e+03 1.16e+03 1.14e+03 1.11e+03 1.10e+03 1.09e+03
AAE-Cox 2.98e+03 2.85e+03 2.73e+03 2.62e+03 2.50e+03 2.38e+03 2.23e+03 1.97e+03 1.47e+03 1.12e+03 1.09e+03 1.08e+03
SAWAR 2.18e+03 1.83e+03 1.55e+03 1.33e+03 1.20e+03 1.14e+03 1.13e+03 1.11e+03 1.10e+03 1.09e+03 1.09e+03 1.09e+03
prostate DRAFT 1.50e+03 1.44e+03 1.36e+03 1.24e+03 1.10e+03 9.38e+02 7.70e+02 6.16e+02 4.90e+02 3.97e+02 3.63e+02 3.37e+02
Noise 1.88e+06 1.21e+06 6.72e+05 3.64e+05 1.97e+05 9.86e+04 4.32e+04 1.69e+04 6.34e+03 2.38e+03 1.46e+03 9.01e+02
FGSM 8.51e+02 8.48e+02 8.41e+02 8.25e+02 7.97e+02 7.47e+02 6.78e+02 5.92e+02 5.05e+02 4.25e+02 3.92e+02 3.63e+02
PGD 7.46e+02 7.32e+02 7.17e+02 6.97e+02 6.69e+02 6.30e+02 5.79e+02 5.22e+02 4.63e+02 4.05e+02 3.79e+02 3.57e+02
AAE-Cox 4.64e+02 4.47e+02 4.31e+02 4.16e+02 4.02e+02 3.89e+02 3.77e+02 3.64e+02 3.52e+02 3.41e+02 3.36e+02 3.31e+02
SAWAR 4.40e+02 4.28e+02 4.15e+02 4.04e+02 3.92e+02 3.81e+02 3.70e+02 3.60e+02 3.51e+02 3.42e+02 3.37e+02 3.33e+02
retinopathy DRAFT 5.04e+02 4.93e+02 4.70e+02 4.35e+02 3.93e+02 3.48e+02 3.02e+02 2.59e+02 2.22e+02 1.92e+02 1.80e+02 1.69e+02
Noise 5.47e+02 5.29e+02 4.96e+02 4.54e+02 4.04e+02 3.54e+02 3.05e+02 2.62e+02 2.24e+02 1.94e+02 1.81e+02 1.69e+02
FGSM 4.16e+02 4.09e+02 3.93e+02 3.69e+02 3.39e+02 3.06e+02 2.73e+02 2.41e+02 2.13e+02 1.89e+02 1.79e+02 1.69e+02
PGD 4.11e+02 4.03e+02 3.88e+02 3.65e+02 3.37e+02 3.05e+02 2.71e+02 2.40e+02 2.12e+02 1.89e+02 1.78e+02 1.69e+02
AAE-Cox 2.27e+02 2.21e+02 2.14e+02 2.09e+02 2.03e+02 1.97e+02 1.91e+02 1.85e+02 1.79e+02 1.73e+02 1.70e+02 1.67e+02
SAWAR 3.27e+02 3.14e+02 2.99e+02 2.82e+02 2.65e+02 2.47e+02 2.29e+02 2.12e+02 1.96e+02 1.81e+02 1.74e+02 1.68e+02
stagec DRAFT 1.38e+02 1.32e+02 1.24e+02 1.15e+02 1.04e+02 9.29e+01 8.09e+01 6.91e+01 5.82e+01 4.90e+01 4.49e+01 4.13e+01
Noise 1.52e+02 1.44e+02 1.35e+02 1.24e+02 1.12e+02 9.97e+01 8.67e+01 7.39e+01 6.18e+01 5.13e+01 4.68e+01 4.27e+01
FGSM 1.04e+02 1.01e+02 9.63e+01 9.11e+01 8.52e+01 7.85e+01 7.12e+01 6.37e+01 5.64e+01 4.96e+01 4.65e+01 4.36e+01
PGD 1.01e+02 9.73e+01 9.31e+01 8.82e+01 8.27e+01 7.65e+01 6.97e+01 6.27e+01 5.58e+01 4.93e+01 4.63e+01 4.35e+01
AAE-Cox 8.00e+01 7.47e+01 6.99e+01 6.55e+01 6.15e+01 5.77e+01 5.42e+01 5.07e+01 4.74e+01 4.41e+01 4.25e+01 4.09e+01
SAWAR 9.46e+01 8.98e+01 8.41e+01 7.80e+01 7.17e+01 6.57e+01 6.03e+01 5.48e+01 4.95e+01 4.45e+01 4.23e+01 4.01e+01
zinc DRAFT 5.01e+02 4.94e+02 4.77e+02 4.45e+02 3.93e+02 3.19e+02 2.39e+02 1.73e+02 1.27e+02 9.62e+01 8.52e+01 7.64e+01
Noise 5.54e+02 5.34e+02 5.00e+02 4.52e+02 3.87e+02 3.10e+02 2.35e+02 1.75e+02 1.32e+02 1.01e+02 8.85e+01 7.86e+01
FGSM 2.89e+02 2.79e+02 2.63e+02 2.39e+02 2.10e+02 1.79e+02 1.52e+02 1.28e+02 1.09e+02 9.21e+01 8.48e+01 7.82e+01
PGD 2.77e+02 2.67e+02 2.51e+02 2.29e+02 2.02e+02 1.74e+02 1.48e+02 1.26e+02 1.07e+02 9.18e+01 8.48e+01 7.85e+01
AAE-Cox 1.66e+02 1.57e+02 1.48e+02 1.38e+02 1.27e+02 1.17e+02 1.08e+02 9.94e+01 9.16e+01 8.44e+01 8.10e+01 7.76e+01
SAWAR 2.11e+02 1.92e+02 1.73e+02 1.57e+02 1.42e+02 1.28e+02 1.15e+02 1.04e+02 9.42e+01 8.52e+01 8.11e+01 7.72e+01
TABLE XIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the FGSM adversarial attack.