### v5

# Survival Analysis with Adversarial Regularization

Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent research indicates that Neural Networks (NNs) can effectively capture complex data patterns in SA, whereas simple generalized linear models often fall short in this regard. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade NN model performance. To address this, we leverage advances in NN verification to develop training objectives for robust, fully-parametric SA models. Specifically, we propose an adversarially robust loss function based on a Min-Max optimization problem. We employ CROWN-Interval Bound Propagation (CROWN-IBP) to tackle the computational challenges inherent in solving this Min-Max problem. Evaluated on 10 SurvSet datasets, Survival Analysis with Adversarial Regularization (SAWAR) consistently outperforms baseline adversarial training methods and state-of-the-art (SOTA) deep SA models on negative log-likelihood (Negative Log Likelihood (NegLL)), integrated Brier score (Integrated Brier Score (IBS)), and concordance index (Concordance Index (CI)). These improvements persist across a range of input covariate perturbations. Thus, we demonstrate that adversarial robustness enhances SA predictive performance and calibration, mitigating input covariate uncertainty and improving generalization across diverse datasets by up to 150% compared to baselines.
https://github.com/mlpotter/SAWAR
Index Terms:
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP, Cox-WeibullI Introduction
Survival Analysis (SA) models the time until an event occurs with extensive applications in various fields including medicine, system reliability, economics, and marketing. Medical research uses Survival Analysis (SA) to evaluate treatment efficacy and illness progression [1]. Engineers assess a system’s reliability with SA to improve design and maintenance [2]. Economists and social scientists use SA (also known as duration analysis) to investigate the length of unemployment spells and other major life events [3]. Marketing uses SA to aid in the examination of customer retention and churn rates [4]. While many of these domains have traditionally relied on linear parametric SA models [5], recent advancements have leveraged Neural Networks (NNs) to handle such complex time-to-event data [6]
A major challenge in SA is developing models that effectively handle complex time-to-event data while also ensuring robustness against dataset uncertainty, accurate calibration, and strong generalization. For instance, input covariate uncertainty in medical applications can arise from various sources, including human factors such as illegible handwriting by physicians [7] and poor medical record-keeping [8], as well as inherent noise from medical equipment during data collection [9], termed as aleatoric uncertainty. Most of the existing approaches address aleatoric uncertainty by assuming noise in the logarithm of measured time-to-event data [10]. However, they often overlook uncertainty originating from the model’s input covariates.
SA models should generalize over various input covariates to maintain high accuracy in real-life applications. In medical studies, this entails high accuracy over diverse patient demographics, institutions, and data acquisition settings. While traditional SA models, such as Linear Cox-Weibull model [5], offer inductive biases specific to time-to-event analysis that aid generalization, they struggle with complex data relationships. On the other hand, while NN models can manage complex data, they may not generalize well to unseen data [11]. Combining the inductive biases of traditional SA models with NN techniques, such as modeling intensity or hazard functions with NNs, can both improve generalization and handle complex data [12]. Nonetheless, NNs in SA may still suffer from poor model calibration, even with good generalization.
NNs are particularly sensitive to input perturbations [13]. Adversaries can exploit this sensitivity to manipulate data, causing models to make over- or under-confident predictions [14, 15]. Additionally, data perturbations may be intentionally introduced with techniques like differential privacy to protect sensitive information, such as personally identifiable information of patients [16]. Although better calibration does not always enhance generalization [17], integrating adversarial robustness to input covariate perturbation can jointly improve model stability, trustworthiness, and calibration when encountering unseen data [18].
Despite these challenges and the sensitivity of NNs to input uncertainties, robustness techniques for handling input covariate perturbations in NN-based SA remain underexplored. To address these challenges, we make the following contributions:
We propose a novel training objective, called Survival Analysis with Adversarial Regularization (SAWAR), designed to create robust, fully-parametric NN-based SA models. We achieve this by adding an adversarial regularization term formulated as a Min-Max objective. The inner maximization is approximated with convex relaxations (e.g., CROWN-IBP), producing a tractable upper bound used during training to improve calibration and generalization under covariate uncertainty.
Extensive experiments on 10 publicly available SA datasets demonstrate that current adversarial training methods for NN-based SA models are highly sensitive to covariate perturbations. We quantify this sensitivity using Concordance Index (CI), Integrated Brier Score (IBS), and Negative Log Likelihood (NegLL) metrics and by comparing the resulting population survival curves under various levels of input covariate perturbation.
SAWAR enhances SA predictive performance and calibration, mitigating input covariate uncertainty and improving generalization by up to 150% compared to baseline adversarial training methods and state-of-the-art (SOTA) NN-based SA models.
II Related Work
The integration of NNs into SA began with initial models that replaced traditional linear approaches. Early works used a shallow Feedforward Neural Network (FFNN) to replace the linear estimate of the log-hazard in a Cox Proportional Hazard (CPH) model [19]. Building on this, Cox-nnet [20] substituted the input covariates of the linear log-hazard function of a CPH model [21] with a deeper NN. Subsequently, DeepSurv [22] further advanced this approach by replacing the entire linear estimate with a deep NN and adding dropout to improve generalization. The trend continued with mixture models like DeepWeiSurv [23], which used NNs to regress Weibull mixture distribution parameters, offering increased modeling flexibility while maintaining traditional statistical model biases.
As NN models evolved, the SA field saw the introduction of non-parametric approaches. For instance, DeepHit utilized NNs to learn the distribution of survival times directly without assumptions about the underlying stochastic process [24]. Even more flexible techniques emerged, such as those employing Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples non-parametrically [25, 26, 27], with CGAN serving as an implicit likelihood [28]. This approach defines the survival distribution without assuming a specific functional form [29, 30]. For a more detailed overview of deep learning methods for SA, we refer the readers to [12].
Despite these advances, integrating NNs with traditional SA models has been limited in addressing model calibration and robustness as discussed in Section I. Recent research has thus focused on improving model calibration [31, 32, 33] to ensure that predicted confidence levels reflect true likelihoods [34]. However, better calibration does not necessarily enhance generalization [17]. On the other hand, integrating adversarial robustness has been shown to improve both calibration and generalization when dealing with unexpected data [18]. Nonetheless, there remains a gap in addressing adversarial robustness in SA to ensure that where small input perturbations do not mislead model predictions [15].
Existing adversarial training methods for SA often face trade-offs between model complexity, prediction performance, and the incorporation of physical assumptions. While Generative Adversarial Network (GAN)-inspired methods incorporate adversarial components through a minimax game between the generator and discriminator [35, 36], they typically exhibit training instability and may violate Bayesian Occam’s Razor that advocates simpler models [30]. To balance complexity, interpretability, and generalization, several methods adopt time-to-event inductive biases through specific functional forms of SA models with deep NN and/or non-parametric approaches [22, 20, 23, 37].
One such inductive bias in SA is the memoryless property, satisfied by the exponential distribution [29]. This paper employs the exponential CPH model (c.f. Section III), as it is widely used in SA [38, 39, 40], as well as in related fields like queuing theory [41], reliability analysis [6], and telecommunications [42]. Crucially, this fully-parametric approach allows us to create an adversarially robust NN-based SA model that outperforms SOTA NN-based SA models including such as Deep Regularized Accelerated Failure Time (DRAFT) [25] and Adversarial Autoencoder with a Deep Cox Proportional Hazard (AAE-Cox) [37]. AAE-Cox aims to balance modeling flexibility and inductive biases by combining adversarial GAN autoencoders with Deep Cox proportional networks. AAE-Cox has been shown to outperform methods like DeepSurv, TCAP [43], and DCAP [44].
III Preliminaries on Survival Analysis
SA [21] is a statistical approach used to analyze and model the time until an event occurs (e.g., a disease onset or device failure). Instances/individuals may leave a study before an event is observed; this is recorded as right censoring [45]. Each of the independent instances/individuals (indexed by has: covariates measured at study entry, a recorded time , and an event indicator . We use the convention:
: the event was observed for instance ; then is the event time.
: instance is right-censored; then is the censoring time (the last time the instance was known to be event-free).
A visual of time-to-event data with right censoring is shown in Fig. 1. In summary, the dataset is written as
We next define the required probabilistic functions to model time-to-event data next.
III-A Cox Proportional Hazard Model
Let us denote as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) and Cumulative Distribution Function (CDF) . The survival function is the complement of this CDF [46, 21]. A SA model has the parametric form:
where is the cumulative hazard and:
is the instantaneous risk at time . Informally, is the probability that an event occurs within a small time window , given that the event has not yet occurred.
To incorporate each instance’s covariates in predicting hazard, we use the prevalent CPH SA model formulation [5]:
where is a baseline hazard, is the relative hazard (or relative risk) associated with covariates , and is an NN with parameters in our setting. The exponential CPH model assumes a time-independent baseline hazard rate , which allows us to “absorb” the term into the bias of the NN:
Further, the exponential CPH model’s PDF, CDF, and complement CDF are given by , , , respectively. is also known as the instance survival function. The population survival curve is defined as the marginalization of the instance survival function over the input covariates:
Since the distribution of input covariates is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate over a dataset [31]:
Given these definitions, the next section describes the baseline objective for estimating SA model parameters from dataset .
III-B Baseline Objective Functions
We employ an objective function that combines negative log likelihood (Section III-B1) and rank correlation (Section III-B2), following [47, 25].
III-B1 Right-Censored Log Likelihood Objective
The objective is the logarithm of the time-to-event likelihood over as specified by the CPH model [21]:
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time that occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs, where it is known that no event occurred at least up until the censoring time. A higher right-censored log likelihood indicates a better SA model fit to the data .
III-B2 Ranking Objective
The objective is a ranking loss that penalizes an incorrect ordering of pairs , where instance with event time should have a higher probability of event occurrence than instance at time (:
where and indicates an “acceptable” comparison pair [47, 48]. Two instances are comparable if the instance with lower time experienced an event, i.e., if . We set for all experiments. This objective is a continuous differentiable proxy for the non-continuous non-differentiable concordance index [49].
III-B3 Combined Objective
The combined objective over the right-censored dataset is:
where controls a trade-off between maximizing the right-censored likelihood of observing and maximizing a rank correlation Accordingly, is estimated by solving the following optimization problem:
IV Preliminaries on Adversarial Training
Adversarial training [36, 50, 51] improves robustness and generalization of NNs to uncertain input covariates. Rather than training on “clean” data, most adversarial training approaches perturb the input covariates throughout the training process to encourage the NN model to optimize for parameters that are resilient against such perturbations. For example, [13] formulated adversarial training as solving a Min-Max robust optimization problem:
where , and is the -ball of radius around .
In practice, the Min-Max problem posed in Eq. 12 [13] is intractable to solve exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over input covariates with the current NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different input covariate uncertainties. To understand the robustness properties of SA models to a variety of input covariate uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Verifiable Robustness
One such approximation to the Min-Max objective in Eq. 12 is to use techniques from NN verification during training. NN robustness verification algorithms determine the output neuron’s upper and lower bounds for all input covariates in a set , typically constrained by a norm-bounded perturbation [51]. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input covariates within the set [50, 51, 52]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [53, 50] showed the linear relaxation can be extended to the entire objective function:
where is a convex relaxation that upper-bounds . This relaxation is convex with respect to the NN weights, enabling efficient optimization. By optimizing an upper bound on the inner maximization, the procedure accounts for a strong adversary during training.
IV-B Adversarial Perturbations
Another way to approximate the inner maximization of Eq. 12 is to use adversarial perturbations to the input covariates.
IV-B1 Fast Gradient Sign Method (FGSM)
While CROWN-IBP provides an upper bound on the inner maximization (i.e., ), adversarial perturbation algorithms typically provide a lower bound:
FGSM [54] uses the gradient of the objective function with respect to the input to perturb in the direction that maximizes the objective . Projected Gradient Descent (PGD) [13] iteratively applies that same type of first-order perturbation times (or until convergence is reached), but constrains the perturbation of to remain within a specified set (e.g., -ball of ). FGSM is a special case of PGD, with and the -ball.
We only perturb the input covariates and not the time-to-event . The perturbed input covariates are calculated as:
where is the iteration step and is a step size parameter.
IV-B2 Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [55]. Therefore, this paper also considers Gaussian Noise perturbations with variance bounded within an -ball of :
Explicitly modeling input covariate uncertainty via a Min–Max formulation enables principled propagation of aleatoric uncertainty. Techniques such as CROWN-IBP provide formal analytical upper bounds for predefined perturbation sets, whereas gradient-based methods (FGSM, PGD) produce empirical, sensitivity-driven perturbations without formal guarantees. Approaches that do not specify uncertainty sets (e.g., GAN-based methods or AAE-Cox) rely on learned latent representations and therefore lack formal coverage of uncertainty and direct control over the strength of adversarial regularization.
Building on this foundation of incorportating aleatoric uncertainty via a Min-Max objective, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing SA models robust to input covariate uncertainty, generalizable, and calibrated. We propose a training objective for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and generalization on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with a Min-Max adversarial regularization term (Eq. 12) [51, 50], as detailed in the following sections.
V-A Min-Max Formulation
When solving for in SA, many approaches treat each measurement in dataset as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise in the logarithm of measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [21]. Instead, the proposed SAWAR approach assumes that each input covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we add an adversarial regularization term which optimizes the NN parameters for the worst-case realizations of Eq. 10:
We use the open-source, PyTorch-based library auto_LiRPA [53] to compute the CROWN-IBP upper bound to the inner maximization term in Eq. 18, which is fully differentiable for backpropogation.
V-B Adversarial Robustness Regularization
[51] shows that tight relaxations to the inner maximization of a Min-Max problem can over-regularize the NN, leading to poor predictive performance. Therefore, we balance the objective in Eq. 10 with the adversarial robustness objective in Eq. 18 [50, 52] to combine predictive performance and adversarial robustness as:
where .
In initial experiments, we conducted a hyper parameter search over to assess its impact on the survival curves. At , the objective function reduces to the CROWN-IBP Min-Max formulation, resulting in narrower credible intervals. This is expected, as minimizing deviations in the survival curve due to input covariate perturbations may be achieved by reducing reliance on the covariates. Thus, may lead to over regularization hurting predictive performance of individual patients, but should still have a good estimate of the global population survival curve. At , the objective aligned with the standard log-likelihood in traditional SA for exponential-Cox proportional hazard models. Therefore, we may view as a hyperparameter which controls the tradeoff between adversarial regularization and predictive performance. We chose to equally balance adversarial regularization and predictive performance.
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating input covariate uncertainty in our SAWAR objective [18] via the adversarial regularization term in Eq. 19 and show SAWAR’s improved performance via extensive experiments in the next section.
VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (Section VI-A) to quantify the impact of input covariate perturbation after the robust training is complete.
To evaluate SAWAR, we train one neural network per adversarial training method. Standard approaches approximate the inner maximization in Eq. 18 using FGSM (Eq. 15), PGD, or Gaussian noise (Section IV-B2). We also include DRAFT and AAE-Cox as baselines, which alter the training objective or NN architecture. For each trained model, we then perturb test covariates by ( means no perturbation) and evaluate CI, IBS, and NegLL at each . Metrics are computed using lifelines [56] and sksurv [48].
VI-A Datasets
We use the API SurvSet [57] to download the benchmark medical SA datasets shown in Table I and described below:
| Dataset | ||||
| TRACE | 1878 | 4 | 4 | 2 |
| stagec | 146 | 4 | 15 | 3 |
| flchain | 7874 | 4 | 26 | 6 |
| Aids2 | 2839 | 3 | 11 | 1 |
| Framingham | 4699 | 2 | 12 | 5 |
| dataDIVAT1 | 5943 | 3 | 14 | 2 |
| prostate | 502 | 6 | 16 | 9 |
| zinc | 431 | 11 | 18 | 2 |
| retinopathy | 394 | 5 | 9 | 2 |
| LeukSurv | 1043 | 2 | 24 | 5 |
1.
The TRACE [58] dataset is from a study on a subset of patients admitted after myocardial infarction to examine various risk factors.
2.
The stagec [59] dataset is from a study exploring the prognostic value of flow cytometry for patients with stage C prostate cancer.
3.
The flchain [60] dataset is from a study on the relationship between serum free light chain (FLC) and mortality of Olmsted Country residents aged 50 years or more.
4.
The Aids2 [61] dataset is from patients diagnosed with AIDS in Australia before 1 July 1991.
5.
The Framingham [62] dataset is from the first prospective study of cardiovascular disease with identification of risk factors and joint effects.
6.
The dataDIVAT1 [63] dataset is from the first sample from the DIVAT Data Bank for French kidney transplant recipients from DIVAT cohort.
7.
The prostate [64] dataset is from a randomised clinical trial comparing treatment for patients with stage 3 and stage 4 prostate cancer.
8.
The zinc [65] is from the first study to examine association between tissue elemental zinc levels and the esophageal squamous cell carcinoma.
9.
The retinopathy [66] is from the trial of laser coagulation as treatment to delay diabetic retinopathy.
10.
The LeukSurv [67] is from the study on survival of acute myeloid leukemia and connection to spatial variation in survival.
We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively.
VI-B Experiment Setup
The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM. The NN has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions. For AAE-Cox, we adopt the publicly available code and hyperparameters [37]. For DRAFT, we use the same hyperparameters as SAWAR as described below, excluding hyperparameters related to adversarial regularization.
The NN parameters are optimized by solving Eq. (19) using stochastic gradient descent with the ADAM optimizer [68]. We set and use early stopping [69] with respect to the validation set to prevent overfitting. We introduce the covariate perturbations after a set number of epochs that varies per dataset. A smooth -scheduler linearly increases the -perturbation magnitude from to within a 30 epoch window after the aforementioned set number of epochs. We do not allow early stopping until after the maximum -perturbation magnitude of is achieved.
VI-C Adversary Setup
We apply two adversarial perturbation methods on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B1) and a novel worst-case perturbation. The perturbed covariates with respect to the input are found via Eq. 15 and the corresponding hazard becomes:
We define the worst-case perturbation as the maximum hazard with respect to the input covariate uncertainty set:
which we solve via CROWN-IBP (Section IV-B1) by optimizing a tight approximation of the upper bound of the maximization term.
For the exponential CPH distribution, as the hazard increases with the -perturbation, the probability of survival decreases monotonically. Accordingly, the -worst-case survival curve has the lowest survival probability at a given time with respect to the -perturbed input covariates:
We evaluate SA metrics for , where represents no perturbation to input covariates, in order to study how sensitive different adversarial training methods are to covariate perturbations.
| Concordance Index | Integrated Brier Score | Negative Log Likelihood | ||||||||||||||||
| DRAFT | Noise | FGSM | PGD | AAE-Cox | SAWAR | DRAFT | Noise | FGSM | PGD | AAE-Cox | SAWAR | DRAFT | Noise | FGSM | PGD | AAE-Cox | SAWAR | |
| 1.00 | 3.8 | 4.4 | 3.1 | 2.7 | 3.8 | 3.2 | 5.4 | 5.6 | 3.2 | 2.3 | 2.6 | 1.9 | 5.1 | 5.8 | 3.6 | 2.6 | 2.5 | 1.4 |
| 0.90 | 3.8 | 4.4 | 3.25 | 2.6 | 3.9 | 3.05 | 5.4 | 5.6 | 3.4 | 2.4 | 2.6 | 1.6 | 5.1 | 5.8 | 3.6 | 2.6 | 2.5 | 1.4 |
| 0.80 | 3.9 | 4.4 | 3.65 | 2.9 | 3.6 | 2.55 | 5.4 | 5.5 | 3.6 | 2.5 | 2.5 | 1.5 | 5.1 | 5.8 | 3.6 | 2.6 | 2.5 | 1.4 |
| 0.70 | 4.45 | 4.75 | 3.5 | 3.05 | 3.1 | 2.15 | 5.4 | 5.5 | 3.6 | 2.5 | 2.6 | 1.4 | 5.1 | 5.8 | 3.6 | 2.7 | 2.5 | 1.3 |
| 0.60 | 4.9 | 5.0 | 3.5 | 3.0 | 2.8 | 1.8 | 5.3 | 5.6 | 3.6 | 2.7 | 2.5 | 1.3 | 5.2 | 5.7 | 3.6 | 2.7 | 2.5 | 1.3 |
| 0.50 | 5.0 | 4.95 | 3.6 | 3.05 | 2.8 | 1.6 | 5.3 | 5.5 | 3.7 | 2.7 | 2.5 | 1.3 | 5.2 | 5.7 | 3.9 | 2.6 | 2.3 | 1.3 |
| 0.40 | 5.3 | 4.7 | 3.7 | 3.3 | 2.6 | 1.4 | 5.3 | 5.5 | 3.9 | 2.7 | 2.3 | 1.3 | 5.3 | 5.6 | 3.9 | 2.6 | 2.3 | 1.3 |
| 0.30 | 5.1 | 5.2 | 3.7 | 3.0 | 2.6 | 1.4 | 5.3 | 5.5 | 3.9 | 2.7 | 2.3 | 1.3 | 5.1 | 5.7 | 3.9 | 2.8 | 2.2 | 1.3 |
| 0.20 | 5.5 | 5.1 | 3.9 | 2.7 | 2.3 | 1.5 | 5.3 | 5.5 | 3.9 | 2.9 | 2.1 | 1.3 | 4.9 | 5.8 | 4.0 | 2.9 | 2.1 | 1.3 |
| 0.10 | 5.2 | 5.2 | 4.0 | 3.2 | 1.9 | 1.5 | 5.0 | 5.7 | 4.1 | 3.0 | 1.9 | 1.3 | 4.5 | 5.9 | 4.2 | 3.1 | 1.9 | 1.4 |
| 0.05 | 5.4 | 5.0 | 3.9 | 3.3 | 2.0 | 1.4 | 4.6 | 5.6 | 4.3 | 3.4 | 1.6 | 1.5 | 4.5 | 5.9 | 4.3 | 3.3 | 1.7 | 1.3 |
| 0.00 | 3.3 | 4.3 | 4.5 | 4.1 | 2.5 | 2.3 | 2.7 | 4.2 | 4.7 | 4.7 | 2.5 | 2.2 | 3.0 | 5.4 | 4.9 | 4.2 | 1.8 | 1.7 |
| Concordance Index | Integrated Brier Score | Negative Log Likelihood | ||||||||||||||||
| DRAFT | Noise | FGSM | PGD | AAE-Cox | SAWAR | DRAFT | Noise | FGSM | PGD | AAE-Cox | SAWAR | DRAFT | Noise | FGSM | PGD | AAE-Cox | SAWAR | |
| 1.00 | 3.6 | 3.6 | 3.6 | 4.4 | 4.2 | 1.6 | 2.6 | 4.35 | 3.65 | 3.55 | 5.85 | 1.0 | 3.1 | 4.8 | 3.4 | 2.8 | 5.9 | 1.0 |
| 0.90 | 3.4 | 3.9 | 3.6 | 4.3 | 4.4 | 1.4 | 2.9 | 4.25 | 3.65 | 3.35 | 5.85 | 1.0 | 3.1 | 4.8 | 3.4 | 2.8 | 5.9 | 1.0 |
| 0.80 | 3.3 | 3.8 | 3.7 | 4.3 | 4.6 | 1.3 | 2.7 | 4.25 | 3.75 | 3.45 | 5.85 | 1.0 | 3.1 | 4.8 | 3.4 | 2.8 | 5.9 | 1.0 |
| 0.70 | 3.4 | 3.8 | 3.8 | 4.1 | 4.6 | 1.3 | 2.8 | 4.65 | 3.65 | 3.15 | 5.75 | 1.0 | 3.1 | 4.8 | 3.5 | 2.7 | 5.9 | 1.0 |
| 0.60 | 3.2 | 3.65 | 3.5 | 4.2 | 5.3 | 1.15 | 2.8 | 4.95 | 3.8 | 2.7 | 5.75 | 1.0 | 3.1 | 4.8 | 3.5 | 2.7 | 5.9 | 1.0 |
| 0.50 | 3.0 | 3.6 | 3.5 | 4.1 | 5.6 | 1.2 | 3.1 | 4.8 | 3.8 | 2.7 | 5.6 | 1.0 | 3.1 | 4.8 | 3.5 | 2.7 | 5.9 | 1.0 |
| 0.40 | 2.9 | 3.8 | 3.4 | 4.0 | 5.8 | 1.1 | 3.3 | 4.8 | 3.7 | 2.5 | 5.7 | 1.0 | 3.1 | 4.8 | 3.6 | 2.6 | 5.9 | 1.0 |
| 0.30 | 2.8 | 3.9 | 3.6 | 3.3 | 6.0 | 1.4 | 3.6 | 4.6 | 3.5 | 2.4 | 5.9 | 1.0 | 3.1 | 4.9 | 3.6 | 2.5 | 5.9 | 1.0 |
| 0.20 | 3.2 | 3.6 | 3.9 | 2.9 | 6.0 | 1.4 | 3.7 | 4.9 | 3.4 | 2.3 | 5.6 | 1.1 | 3.5 | 5.1 | 3.3 | 2.3 | 5.8 | 1.0 |
| 0.10 | 3.2 | 4.0 | 4.2 | 3.6 | 4.0 | 2.0 | 3.9 | 5.5 | 3.6 | 2.9 | 3.8 | 1.3 | 3.9 | 5.8 | 4.1 | 3.1 | 3.0 | 1.1 |
| 0.05 | 3.4 | 4.2 | 4.4 | 4.1 | 2.6 | 2.3 | 3.6 | 5.1 | 4.3 | 4.1 | 2.2 | 1.7 | 3.7 | 5.6 | 4.4 | 3.9 | 1.9 | 1.5 |
| 0.00 | 3.3 | 4.3 | 4.5 | 4.1 | 2.5 | 2.3 | 2.7 | 4.2 | 4.7 | 4.7 | 2.5 | 2.2 | 3.0 | 5.4 | 4.9 | 4.2 | 1.8 | 1.7 |
To simulate realistic conditions, we assume that the 10 benchmark SA datasets are “clean” and contain various input covariates (e.g., age, drug dosage in mg, tumor size) and survival outcomes (e.g., time until death, censored observations). By applying adversarial attack techniques, we perturb these input covariates to maximize the SA error.
The perturbations are designed to be subtle yet impactful, meaning they are small in magnitude but cause significant changes in the predicted survival times. Since the data is standard normalized, a perturbation of would shift an input covariate by 0.1 standard deviations from its current value. For example, increasing a patient weight from 70 kg to 72 kg (a small real‑world change that would not alter clinical categorization of “Normal weight” according to National Institute of Health (NIH) [70]) may still produce a disproportionate change in the patient’s predicted survival for a sensitive model, illustrating the need for adversarial regularization.
VI-D Results
We analyze how input covariate perturbations impact the characteristics of the population survival curves, and SA metrics CI,IBS, and NegLL.
VI-D1 Comparison to SOTA
We show the average ranking of each adversarial training method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and -worst-case perturbations in Tables III and II, respectively. We empirically show SAWAR leads to increased performance, consistently ranking the best across various -perturbation magnitudes. Thus, SAWAR has better generalization performance, better calibration, and better adversarial robustness by up to 150% compared to baseline adversarial training methods and SOTA NN-based SA models.
In order to confirm that SAWAR showed statistically significant improved performance, we performed a Friedman hypothesis test () for each metric (CI, IBS, and negative log-likelihood), using adversarial training methods as “treatments” and combinations of dataset, perturbation magnitude, and perturbation method as “blocks”. This test, suited for repeated measures, assessed whether our adversarial training method yielded more robust, well-fitted, and calibrated models across perturbation attacks compared to other adversarial training methods. All Friedman hypothesis tests had statistically signifcant results, which prompted post-hoc Conover-Iman test for pairwise comparisons amongst adversarial training methods, with p-values adjusted via the Benjamini-Hochberg procedure (critical difference diagrams shown in Fig. 2).
More detailed results, along with the exact metric values for each of the worst-case and the FGSM adversarial attacks for each dataset are available in Appendices D, E, F, A, B and C.
VI-D2 Survival Curve Distribution
Each instance in the dataset has its own survival curve . Thus, let us assume that the input covariates are random variables . Then the hazard becomes a random variable that occurs from the stochasticity of the input covariates . Moreover, we treat as a stochastic process :
Then, instance’s survival curves are visualized as the credible interval, where is sampled from the dataset distribution . We compute the statistical quantities of such as expectation, lower 5th quantile , and upper 95th quantile .
Fig. 3 visualizes the credible interval and mean of the survival function stochastic process in Eq. 23. The 90% credible interval of the stochastic process changes with respect to the strength of the adversarial training method (Fig. 3). We observe the trend that the credible interval slightly narrows down as the strength (regularization) of the adversarial training method increases (Fig. 3). However, this trend has an exception with AAE-Cox. We hypothesize that since AAE-Cox is not truly an adversarial robust method (i.e., no perturbations to the input covariates during training), and by introducing more parameters via an encoder of a GAN, there are more model parameters to attack. The variance of the hazard decreasing leads to a more robust model against input covariate perturbations. For example, when , then there is a mode collapse such that . This is the case for point-estimates of the parameter derived via maximum likelihood estimation of a Poisson process without input covariates. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 19, we introduce a trade-off between individualized survival curves with vulnerability to input covariate perturbations and a population survival curve that is completely resilient to input covariate perturbations. In doing so, while exhibiting significantly higher adversarial robustness, SAWAR’s credible interval remains relatively unchanged compared to the DRAFT method.
VI-D3 Population Survival Curve
The -perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger (regularization) the adversarial training method, the “closer” the -worst-case population curve becomes to the unperturbed population survival curve for a given (Fig. 4). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples , to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 4 demonstrates the strong calibration of SAWAR model [32].
Overall, robust and well-calibrated SA models are essential for making reliable clinical decisions, including evaluating treatment efficacy, analyzing dose-response relationships, determining optimal treatment timing, performing subgroup analysis, and assessing comparative effectiveness. In contrast, SA models prone to overfitting or sensitive to small perturbations in input covariates can yield unreliable or even harmful recommendations, potentially leading to fatal outcomes.
VII Conclusions
This paper introduces SAWAR, a training objective that approximates a Min–Max optimization with CROWN-IBP to improve adversarial robustness in fully‑parametric NN-based SA models. We evaluated SAWAR on 10 benchmark medical time‑to‑event datasets and compared it to standard adversarial training methods (Gaussian noise, FGSM, PGD) and SOTA models (DRAFT, AAE-Cox). Our results show that SAWAR consistently improves performance across CI, IBS, and NegLL for various input covariate perturbation magnitudes . Compared with traditional adversarial training method, SAWAR offers superior resilience to input covariate perturbations, with relative gains in CI and IBS ranging from 1% to 150%. Overall, SAWAR enhances predictive accuracy, robustness, and calibration, effectively mitigating input covariate uncertainty and improving generalization to unseen datasets.
Future work will involve extending the approach to the Weibull CPH model to enable time-varying hazard rates, which will require adapting auto_LiRPA to handle power operations such as , where is a learnable parameter.
Appendix A Concordance Index Worst-Case Perturbation
We find that as the -perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better). We note that for very large , since our data is standard normalized all methods begin to fail.
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | |
| 72.8 | 79.69 | 92.54 | 90.82 | 81.56 | 66.21 | 41.01 | 12.83 | 3.19 | 1.88 | 1.72 | 1.6 |
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | ||
| Dataset | Algorithm | ||||||||||||
| Aids2 | DRAFT | 0.516 | 0.519 | 0.522 | 0.525 | 0.527 | 0.53 | 0.536 | 0.551 | 0.566 | 0.572 | 0.572 | 0.57 |
| Noise | 0.513 | 0.515 | 0.518 | 0.521 | 0.525 | 0.529 | 0.535 | 0.548 | 0.568 | 0.571 | 0.57 | 0.569 | |
| FGSM | 0.512 | 0.516 | 0.52 | 0.523 | 0.526 | 0.532 | 0.537 | 0.55 | 0.565 | 0.569 | 0.568 | 0.566 | |
| PGD | 0.506 | 0.509 | 0.513 | 0.518 | 0.523 | 0.528 | 0.535 | 0.552 | 0.567 | 0.569 | 0.568 | 0.567 | |
| AAE-Cox | 0.504 | 0.508 | 0.507 | 0.506 | 0.505 | 0.501 | 0.499 | 0.504 | 0.51 | 0.561 | 0.573 | 0.573 | |
| SAWAR | 0.565 | 0.568 | 0.568 | 0.569 | 0.569 | 0.57 | 0.57 | 0.57 | 0.57 | 0.57 | 0.569 | 0.569 | |
| Framingham | DRAFT | 0.618 | 0.634 | 0.649 | 0.663 | 0.676 | 0.688 | 0.699 | 0.708 | 0.716 | 0.72 | 0.721 | 0.722 |
| Noise | 0.606 | 0.622 | 0.637 | 0.653 | 0.666 | 0.679 | 0.691 | 0.703 | 0.713 | 0.718 | 0.719 | 0.719 | |
| FGSM | 0.582 | 0.598 | 0.611 | 0.626 | 0.641 | 0.655 | 0.671 | 0.691 | 0.708 | 0.715 | 0.715 | 0.715 | |
| PGD | 0.539 | 0.559 | 0.582 | 0.605 | 0.627 | 0.647 | 0.669 | 0.693 | 0.71 | 0.717 | 0.717 | 0.717 | |
| AAE-Cox | 0.514 | 0.522 | 0.529 | 0.533 | 0.539 | 0.553 | 0.569 | 0.57 | 0.582 | 0.73 | 0.733 | 0.733 | |
| SAWAR | 0.721 | 0.725 | 0.729 | 0.731 | 0.733 | 0.734 | 0.735 | 0.736 | 0.736 | 0.737 | 0.737 | 0.737 | |
| LeukSurv | DRAFT | 0.589 | 0.593 | 0.596 | 0.598 | 0.601 | 0.602 | 0.608 | 0.62 | 0.63 | 0.632 | 0.633 | 0.633 |
| Noise | 0.545 | 0.548 | 0.545 | 0.547 | 0.551 | 0.56 | 0.581 | 0.609 | 0.63 | 0.628 | 0.626 | 0.624 | |
| FGSM | 0.511 | 0.514 | 0.52 | 0.525 | 0.535 | 0.549 | 0.579 | 0.612 | 0.637 | 0.638 | 0.636 | 0.635 | |
| PGD | 0.498 | 0.498 | 0.501 | 0.51 | 0.521 | 0.543 | 0.576 | 0.616 | 0.64 | 0.643 | 0.64 | 0.638 | |
| AAE-Cox | 0.559 | 0.552 | 0.542 | 0.555 | 0.548 | 0.55 | 0.548 | 0.537 | 0.545 | 0.631 | 0.658 | 0.656 | |
| SAWAR | 0.495 | 0.51 | 0.524 | 0.536 | 0.552 | 0.57 | 0.588 | 0.609 | 0.633 | 0.658 | 0.669 | 0.673 | |
| TRACE | DRAFT | 0.581 | 0.605 | 0.633 | 0.66 | 0.685 | 0.708 | 0.726 | 0.736 | 0.741 | 0.743 | 0.744 | 0.745 |
| Noise | 0.576 | 0.598 | 0.621 | 0.645 | 0.669 | 0.691 | 0.712 | 0.728 | 0.737 | 0.742 | 0.744 | 0.745 | |
| FGSM | 0.581 | 0.603 | 0.629 | 0.653 | 0.674 | 0.697 | 0.717 | 0.732 | 0.739 | 0.743 | 0.744 | 0.744 | |
| PGD | 0.571 | 0.595 | 0.621 | 0.646 | 0.668 | 0.691 | 0.714 | 0.73 | 0.739 | 0.743 | 0.744 | 0.745 | |
| AAE-Cox | 0.432 | 0.44 | 0.447 | 0.451 | 0.46 | 0.47 | 0.489 | 0.524 | 0.628 | 0.743 | 0.746 | 0.747 | |
| SAWAR | 0.714 | 0.722 | 0.728 | 0.733 | 0.735 | 0.737 | 0.739 | 0.742 | 0.744 | 0.746 | 0.747 | 0.747 | |
| dataDIVAT1 | DRAFT | 0.573 | 0.585 | 0.598 | 0.611 | 0.625 | 0.636 | 0.646 | 0.653 | 0.657 | 0.657 | 0.656 | 0.655 |
| Noise | 0.532 | 0.541 | 0.552 | 0.562 | 0.575 | 0.589 | 0.603 | 0.619 | 0.631 | 0.636 | 0.635 | 0.635 | |
| FGSM | 0.546 | 0.555 | 0.567 | 0.578 | 0.588 | 0.6 | 0.612 | 0.626 | 0.641 | 0.647 | 0.646 | 0.646 | |
| PGD | 0.523 | 0.534 | 0.545 | 0.557 | 0.569 | 0.581 | 0.599 | 0.621 | 0.641 | 0.648 | 0.648 | 0.648 | |
| AAE-Cox | 0.597 | 0.589 | 0.589 | 0.585 | 0.58 | 0.571 | 0.571 | 0.566 | 0.596 | 0.662 | 0.663 | 0.663 | |
| SAWAR | 0.64 | 0.645 | 0.649 | 0.654 | 0.658 | 0.661 | 0.664 | 0.665 | 0.667 | 0.668 | 0.669 | 0.67 | |
| flchain | DRAFT | 0.109 | 0.111 | 0.115 | 0.122 | 0.133 | 0.158 | 0.239 | 0.566 | 0.905 | 0.917 | 0.92 | 0.921 |
| Noise | 0.166 | 0.131 | 0.111 | 0.123 | 0.149 | 0.224 | 0.433 | 0.792 | 0.911 | 0.917 | 0.919 | 0.918 | |
| FGSM | 0.115 | 0.153 | 0.245 | 0.431 | 0.72 | 0.885 | 0.91 | 0.914 | 0.918 | 0.922 | 0.922 | 0.922 | |
| PGD | 0.165 | 0.268 | 0.457 | 0.744 | 0.876 | 0.904 | 0.912 | 0.917 | 0.92 | 0.923 | 0.923 | 0.923 | |
| AAE-Cox | 0.527 | 0.648 | 0.668 | 0.553 | 0.394 | 0.511 | 0.51 | 0.476 | 0.11 | 0.179 | 0.925 | 0.926 | |
| SAWAR | 0.593 | 0.684 | 0.866 | 0.918 | 0.922 | 0.925 | 0.926 | 0.927 | 0.927 | 0.927 | 0.927 | 0.927 | |
| prostate | DRAFT | 0.402 | 0.416 | 0.428 | 0.444 | 0.469 | 0.505 | 0.542 | 0.585 | 0.627 | 0.653 | 0.661 | 0.668 |
| Noise | 0.433 | 0.44 | 0.448 | 0.451 | 0.463 | 0.48 | 0.492 | 0.511 | 0.537 | 0.558 | 0.561 | 0.562 | |
| FGSM | 0.448 | 0.453 | 0.46 | 0.465 | 0.475 | 0.486 | 0.502 | 0.515 | 0.541 | 0.564 | 0.569 | 0.571 | |
| PGD | 0.447 | 0.455 | 0.459 | 0.467 | 0.474 | 0.487 | 0.509 | 0.53 | 0.564 | 0.585 | 0.588 | 0.588 | |
| AAE-Cox | 0.41 | 0.411 | 0.406 | 0.41 | 0.407 | 0.403 | 0.399 | 0.391 | 0.414 | 0.646 | 0.686 | 0.691 | |
| SAWAR | 0.608 | 0.623 | 0.637 | 0.652 | 0.665 | 0.671 | 0.672 | 0.673 | 0.667 | 0.663 | 0.66 | 0.657 | |
| retinopathy | DRAFT | 0.553 | 0.568 | 0.578 | 0.597 | 0.616 | 0.632 | 0.645 | 0.653 | 0.663 | 0.666 | 0.667 | 0.669 |
| Noise | 0.573 | 0.591 | 0.605 | 0.62 | 0.632 | 0.644 | 0.653 | 0.661 | 0.666 | 0.667 | 0.668 | 0.668 | |
| FGSM | 0.575 | 0.592 | 0.604 | 0.615 | 0.625 | 0.633 | 0.642 | 0.647 | 0.651 | 0.656 | 0.657 | 0.659 | |
| PGD | 0.571 | 0.589 | 0.599 | 0.61 | 0.621 | 0.63 | 0.64 | 0.647 | 0.651 | 0.655 | 0.656 | 0.657 | |
| AAE-Cox | 0.494 | 0.495 | 0.506 | 0.504 | 0.514 | 0.564 | 0.577 | 0.592 | 0.62 | 0.657 | 0.652 | 0.648 | |
| SAWAR | 0.668 | 0.669 | 0.67 | 0.666 | 0.662 | 0.66 | 0.656 | 0.654 | 0.653 | 0.65 | 0.648 | 0.647 | |
| stagec | DRAFT | 0.358 | 0.378 | 0.382 | 0.406 | 0.425 | 0.449 | 0.454 | 0.475 | 0.489 | 0.504 | 0.507 | 0.512 |
| Noise | 0.353 | 0.373 | 0.39 | 0.407 | 0.436 | 0.466 | 0.485 | 0.5 | 0.53 | 0.544 | 0.549 | 0.555 | |
| FGSM | 0.329 | 0.346 | 0.368 | 0.389 | 0.416 | 0.429 | 0.443 | 0.471 | 0.488 | 0.502 | 0.503 | 0.51 | |
| PGD | 0.341 | 0.352 | 0.381 | 0.399 | 0.419 | 0.431 | 0.442 | 0.47 | 0.49 | 0.496 | 0.498 | 0.505 | |
| AAE-Cox | 0.393 | 0.397 | 0.397 | 0.411 | 0.417 | 0.409 | 0.405 | 0.429 | 0.469 | 0.523 | 0.543 | 0.54 | |
| SAWAR | 0.393 | 0.401 | 0.413 | 0.425 | 0.436 | 0.461 | 0.488 | 0.506 | 0.534 | 0.543 | 0.548 | 0.558 | |
| zinc | DRAFT | 0.262 | 0.27 | 0.284 | 0.296 | 0.317 | 0.355 | 0.44 | 0.579 | 0.712 | 0.755 | 0.765 | 0.77 |
| Noise | 0.318 | 0.328 | 0.343 | 0.366 | 0.401 | 0.466 | 0.562 | 0.661 | 0.734 | 0.766 | 0.773 | 0.776 | |
| FGSM | 0.377 | 0.403 | 0.439 | 0.49 | 0.557 | 0.626 | 0.685 | 0.737 | 0.762 | 0.78 | 0.782 | 0.783 | |
| PGD | 0.384 | 0.406 | 0.443 | 0.495 | 0.559 | 0.632 | 0.695 | 0.741 | 0.77 | 0.778 | 0.78 | 0.781 | |
| AAE-Cox | 0.226 | 0.231 | 0.228 | 0.236 | 0.234 | 0.249 | 0.265 | 0.306 | 0.445 | 0.724 | 0.754 | 0.759 | |
| SAWAR | 0.667 | 0.725 | 0.763 | 0.785 | 0.785 | 0.785 | 0.781 | 0.778 | 0.778 | 0.776 | 0.776 | 0.774 |
Appendix B Integrated Brier Score Worst-Case Perturbation
We find that as the -perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better). We note that for very large , since our data is standard normalized all methods begin to fail.
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | |
| -37.55 | -42.51 | -48.08 | -53.15 | -55.48 | -55.34 | -51.51 | -42.94 | -28.8 | -10.87 | -4.78 | -0.65 |
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | ||
| Dataset | Algorithm | ||||||||||||
| Aids2 | DRAFT | 0.265 | 0.262 | 0.258 | 0.252 | 0.242 | 0.228 | 0.207 | 0.18 | 0.151 | 0.138 | 0.137 | 0.137 |
| Noise | 0.266 | 0.264 | 0.261 | 0.255 | 0.246 | 0.231 | 0.209 | 0.179 | 0.15 | 0.138 | 0.137 | 0.138 | |
| FGSM | 0.265 | 0.263 | 0.258 | 0.251 | 0.239 | 0.221 | 0.197 | 0.167 | 0.144 | 0.137 | 0.137 | 0.138 | |
| PGD | 0.265 | 0.262 | 0.257 | 0.248 | 0.235 | 0.216 | 0.19 | 0.161 | 0.141 | 0.137 | 0.138 | 0.138 | |
| AAE-Cox | 0.269 | 0.269 | 0.269 | 0.268 | 0.265 | 0.259 | 0.25 | 0.231 | 0.183 | 0.138 | 0.137 | 0.137 | |
| SAWAR | 0.14 | 0.139 | 0.138 | 0.138 | 0.137 | 0.137 | 0.137 | 0.137 | 0.137 | 0.137 | 0.137 | 0.137 | |
| Framingham | DRAFT | 0.831 | 0.825 | 0.811 | 0.783 | 0.724 | 0.618 | 0.459 | 0.289 | 0.174 | 0.124 | 0.114 | 0.11 |
| Noise | 0.836 | 0.834 | 0.83 | 0.82 | 0.79 | 0.709 | 0.543 | 0.331 | 0.185 | 0.126 | 0.115 | 0.11 | |
| FGSM | 0.836 | 0.836 | 0.835 | 0.831 | 0.817 | 0.765 | 0.605 | 0.336 | 0.162 | 0.118 | 0.113 | 0.112 | |
| PGD | 0.836 | 0.836 | 0.836 | 0.831 | 0.812 | 0.74 | 0.555 | 0.295 | 0.146 | 0.115 | 0.113 | 0.112 | |
| AAE-Cox | 0.836 | 0.836 | 0.836 | 0.836 | 0.836 | 0.836 | 0.834 | 0.775 | 0.385 | 0.116 | 0.109 | 0.108 | |
| SAWAR | 0.17 | 0.155 | 0.144 | 0.134 | 0.127 | 0.122 | 0.117 | 0.114 | 0.112 | 0.111 | 0.111 | 0.111 | |
| LeukSurv | DRAFT | 0.206 | 0.206 | 0.205 | 0.205 | 0.203 | 0.2 | 0.195 | 0.185 | 0.17 | 0.158 | 0.154 | 0.153 |
| Noise | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.2 | 0.171 | 0.166 | 0.173 | |
| FGSM | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.203 | 0.187 | 0.161 | 0.158 | 0.159 | |
| PGD | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.202 | 0.183 | 0.16 | 0.157 | 0.158 | |
| AAE-Cox | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.206 | 0.198 | 0.162 | 0.154 | 0.152 | |
| SAWAR | 0.194 | 0.19 | 0.184 | 0.178 | 0.171 | 0.165 | 0.16 | 0.155 | 0.151 | 0.148 | 0.147 | 0.146 | |
| TRACE | DRAFT | 0.62 | 0.611 | 0.593 | 0.558 | 0.504 | 0.432 | 0.347 | 0.265 | 0.204 | 0.172 | 0.165 | 0.162 |
| Noise | 0.627 | 0.626 | 0.623 | 0.611 | 0.579 | 0.518 | 0.426 | 0.317 | 0.225 | 0.176 | 0.166 | 0.162 | |
| FGSM | 0.627 | 0.625 | 0.619 | 0.596 | 0.545 | 0.46 | 0.351 | 0.248 | 0.188 | 0.167 | 0.163 | 0.162 | |
| PGD | 0.627 | 0.625 | 0.617 | 0.593 | 0.54 | 0.455 | 0.347 | 0.247 | 0.188 | 0.167 | 0.164 | 0.162 | |
| AAE-Cox | 0.627 | 0.627 | 0.627 | 0.627 | 0.627 | 0.623 | 0.588 | 0.471 | 0.284 | 0.172 | 0.165 | 0.163 | |
| SAWAR | 0.324 | 0.281 | 0.246 | 0.22 | 0.202 | 0.188 | 0.179 | 0.171 | 0.166 | 0.163 | 0.162 | 0.161 | |
| dataDIVAT1 | DRAFT | 0.702 | 0.689 | 0.663 | 0.614 | 0.527 | 0.405 | 0.285 | 0.21 | 0.177 | 0.17 | 0.172 | 0.177 |
| Noise | 0.72 | 0.72 | 0.719 | 0.718 | 0.713 | 0.689 | 0.599 | 0.4 | 0.245 | 0.197 | 0.193 | 0.195 | |
| FGSM | 0.719 | 0.719 | 0.718 | 0.714 | 0.697 | 0.637 | 0.478 | 0.285 | 0.2 | 0.182 | 0.182 | 0.184 | |
| PGD | 0.719 | 0.719 | 0.716 | 0.71 | 0.687 | 0.613 | 0.442 | 0.261 | 0.191 | 0.18 | 0.181 | 0.182 | |
| AAE-Cox | 0.72 | 0.72 | 0.72 | 0.72 | 0.72 | 0.72 | 0.715 | 0.553 | 0.194 | 0.165 | 0.17 | 0.174 | |
| SAWAR | 0.189 | 0.184 | 0.181 | 0.179 | 0.177 | 0.177 | 0.177 | 0.177 | 0.177 | 0.178 | 0.179 | 0.179 | |
| flchain | DRAFT | 0.816 | 0.816 | 0.816 | 0.816 | 0.816 | 0.816 | 0.814 | 0.797 | 0.46 | 0.093 | 0.069 | 0.054 |
| Noise | 0.816 | 0.816 | 0.816 | 0.816 | 0.816 | 0.816 | 0.814 | 0.77 | 0.227 | 0.088 | 0.067 | 0.057 | |
| FGSM | 0.816 | 0.816 | 0.816 | 0.816 | 0.816 | 0.804 | 0.468 | 0.119 | 0.089 | 0.061 | 0.055 | 0.054 | |
| PGD | 0.817 | 0.817 | 0.817 | 0.816 | 0.81 | 0.638 | 0.162 | 0.107 | 0.076 | 0.057 | 0.054 | 0.053 | |
| AAE-Cox | nan | nan | nan | nan | nan | nan | nan | nan | 0.794 | 0.688 | 0.054 | 0.051 | |
| SAWAR | 0.445 | 0.363 | 0.241 | 0.094 | 0.067 | 0.059 | 0.057 | 0.055 | 0.054 | 0.053 | 0.053 | 0.053 | |
| prostate | DRAFT | 0.517 | 0.517 | 0.516 | 0.513 | 0.508 | 0.495 | 0.466 | 0.408 | 0.322 | 0.234 | 0.202 | 0.181 |
| Noise | 0.518 | 0.518 | 0.519 | 0.519 | 0.52 | 0.52 | 0.519 | 0.508 | 0.457 | 0.321 | 0.267 | 0.249 | |
| FGSM | 0.517 | 0.518 | 0.519 | 0.518 | 0.52 | 0.52 | 0.517 | 0.493 | 0.402 | 0.262 | 0.228 | 0.218 | |
| PGD | 0.518 | 0.518 | 0.518 | 0.518 | 0.517 | 0.518 | 0.515 | 0.486 | 0.383 | 0.248 | 0.219 | 0.209 | |
| AAE-Cox | 0.519 | 0.519 | 0.519 | 0.519 | 0.518 | 0.518 | 0.517 | 0.511 | 0.443 | 0.205 | 0.173 | 0.169 | |
| SAWAR | 0.37 | 0.334 | 0.291 | 0.262 | 0.236 | 0.218 | 0.204 | 0.192 | 0.184 | 0.178 | 0.176 | 0.175 | |
| retinopathy | DRAFT | 0.728 | 0.722 | 0.71 | 0.687 | 0.647 | 0.579 | 0.48 | 0.364 | 0.266 | 0.204 | 0.188 | 0.179 |
| Noise | 0.73 | 0.725 | 0.714 | 0.693 | 0.652 | 0.579 | 0.472 | 0.353 | 0.256 | 0.199 | 0.184 | 0.177 | |
| FGSM | 0.725 | 0.715 | 0.697 | 0.665 | 0.61 | 0.526 | 0.419 | 0.314 | 0.237 | 0.196 | 0.186 | 0.182 | |
| PGD | 0.724 | 0.714 | 0.695 | 0.662 | 0.606 | 0.521 | 0.413 | 0.309 | 0.235 | 0.195 | 0.186 | 0.182 | |
| AAE-Cox | 0.733 | 0.733 | 0.733 | 0.733 | 0.733 | 0.732 | 0.716 | 0.596 | 0.281 | 0.183 | 0.181 | 0.181 | |
| SAWAR | 0.588 | 0.548 | 0.499 | 0.44 | 0.378 | 0.317 | 0.265 | 0.224 | 0.197 | 0.184 | 0.181 | 0.181 | |
| stagec | DRAFT | 0.556 | 0.549 | 0.544 | 0.543 | 0.539 | 0.517 | 0.468 | 0.404 | 0.332 | 0.274 | 0.256 | 0.245 |
| Noise | 0.559 | 0.553 | 0.547 | 0.548 | 0.551 | 0.541 | 0.501 | 0.437 | 0.357 | 0.282 | 0.258 | 0.244 | |
| FGSM | 0.547 | 0.545 | 0.549 | 0.553 | 0.545 | 0.511 | 0.457 | 0.391 | 0.323 | 0.278 | 0.265 | 0.258 | |
| PGD | 0.547 | 0.546 | 0.549 | 0.553 | 0.542 | 0.507 | 0.453 | 0.388 | 0.322 | 0.278 | 0.267 | 0.26 | |
| AAE-Cox | 0.568 | 0.568 | 0.568 | 0.568 | 0.567 | 0.559 | 0.545 | 0.544 | 0.474 | 0.291 | 0.268 | 0.263 | |
| SAWAR | 0.521 | 0.497 | 0.469 | 0.436 | 0.401 | 0.365 | 0.334 | 0.303 | 0.276 | 0.256 | 0.249 | 0.243 | |
| zinc | DRAFT | 0.847 | 0.847 | 0.846 | 0.843 | 0.831 | 0.787 | 0.641 | 0.376 | 0.184 | 0.119 | 0.109 | 0.107 |
| Noise | 0.847 | 0.847 | 0.847 | 0.845 | 0.838 | 0.795 | 0.627 | 0.351 | 0.179 | 0.128 | 0.118 | 0.113 | |
| FGSM | 0.847 | 0.847 | 0.845 | 0.835 | 0.782 | 0.611 | 0.356 | 0.189 | 0.133 | 0.116 | 0.113 | 0.112 | |
| PGD | 0.847 | 0.847 | 0.845 | 0.834 | 0.777 | 0.6 | 0.341 | 0.182 | 0.132 | 0.116 | 0.113 | 0.112 | |
| AAE-Cox | 0.847 | 0.847 | 0.847 | 0.847 | 0.847 | 0.847 | 0.846 | 0.823 | 0.507 | 0.119 | 0.109 | 0.108 | |
| SAWAR | 0.654 | 0.532 | 0.401 | 0.289 | 0.21 | 0.161 | 0.133 | 0.118 | 0.111 | 0.109 | 0.109 | 0.109 |
Appendix C Negative Log Likelihood Worst-Case Perturbation
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | ||
| Dataset | Algorithm | ||||||||||||
| Aids2 | DRAFT | 7.88e+05 | 2.75e+05 | 9.79e+04 | 3.51e+04 | 1.31e+04 | 5.06e+03 | 2.06e+03 | 9.59e+02 | 6.16e+02 | 5.45e+02 | 5.41e+02 | 5.41e+02 |
| Noise | 3.15e+05 | 1.30e+05 | 5.45e+04 | 2.30e+04 | 9.53e+03 | 4.06e+03 | 1.79e+03 | 9.02e+02 | 6.06e+02 | 5.44e+02 | 5.41e+02 | 5.41e+02 | |
| FGSM | 1.29e+05 | 5.99e+04 | 2.79e+04 | 1.29e+04 | 5.92e+03 | 2.72e+03 | 1.32e+03 | 7.49e+02 | 5.69e+02 | 5.42e+02 | 5.41e+02 | 5.41e+02 | |
| PGD | 7.29e+04 | 3.62e+04 | 1.80e+04 | 8.83e+03 | 4.29e+03 | 2.10e+03 | 1.09e+03 | 6.82e+02 | 5.57e+02 | 5.42e+02 | 5.41e+02 | 5.41e+02 | |
| AAE-Cox | 6.04e+12 | 2.43e+11 | 1.30e+10 | 6.94e+08 | 4.08e+07 | 2.64e+06 | 1.76e+05 | 1.27e+04 | 1.18e+03 | 5.41e+02 | 5.38e+02 | 5.39e+02 | |
| SAWAR | 5.53e+02 | 5.48e+02 | 5.45e+02 | 5.43e+02 | 5.42e+02 | 5.42e+02 | 5.41e+02 | 5.41e+02 | 5.40e+02 | 5.40e+02 | 5.40e+02 | 5.40e+02 | |
| Framingham | DRAFT | 2.24e+07 | 4.61e+06 | 9.70e+05 | 2.13e+05 | 4.98e+04 | 1.33e+04 | 4.62e+03 | 2.35e+03 | 1.71e+03 | 1.52e+03 | 1.49e+03 | 1.48e+03 |
| Noise | 3.13e+07 | 6.71e+06 | 1.48e+06 | 3.35e+05 | 7.98e+04 | 2.05e+04 | 6.23e+03 | 2.67e+03 | 1.76e+03 | 1.53e+03 | 1.49e+03 | 1.48e+03 | |
| FGSM | 8.34e+09 | 7.44e+08 | 6.79e+07 | 6.46e+06 | 6.50e+05 | 7.31e+04 | 1.05e+04 | 2.77e+03 | 1.67e+03 | 1.50e+03 | 1.49e+03 | 1.48e+03 | |
| PGD | 1.53e+11 | 7.60e+09 | 3.88e+08 | 2.07e+07 | 1.20e+06 | 8.59e+04 | 9.43e+03 | 2.46e+03 | 1.61e+03 | 1.49e+03 | 1.48e+03 | 1.48e+03 | |
| AAE-Cox | 1.38e+26 | 1.22e+23 | 1.20e+20 | 1.10e+17 | 1.21e+14 | 1.26e+11 | 1.63e+08 | 3.73e+05 | 3.68e+03 | 1.49e+03 | 1.47e+03 | 1.47e+03 | |
| SAWAR | 1.67e+03 | 1.62e+03 | 1.58e+03 | 1.55e+03 | 1.53e+03 | 1.51e+03 | 1.50e+03 | 1.49e+03 | 1.48e+03 | 1.48e+03 | 1.48e+03 | 1.48e+03 | |
| LeukSurv | DRAFT | 2.10e+08 | 3.80e+07 | 7.18e+06 | 1.29e+06 | 2.41e+05 | 4.59e+04 | 9.21e+03 | 2.13e+03 | 6.86e+02 | 3.75e+02 | 3.28e+02 | 3.07e+02 |
| Noise | 1.40e+20 | 1.48e+18 | 1.45e+16 | 1.63e+14 | 1.29e+12 | 9.53e+09 | 8.92e+07 | 1.13e+06 | 2.90e+04 | 2.62e+03 | 1.25e+03 | 7.75e+02 | |
| FGSM | 6.24e+12 | 4.01e+11 | 2.30e+10 | 1.27e+09 | 7.06e+07 | 4.34e+06 | 2.73e+05 | 2.04e+04 | 2.18e+03 | 5.11e+02 | 3.72e+02 | 3.25e+02 | |
| PGD | 4.11e+11 | 3.30e+10 | 2.53e+09 | 1.99e+08 | 1.58e+07 | 1.30e+06 | 1.12e+05 | 1.10e+04 | 1.50e+03 | 4.34e+02 | 3.39e+02 | 3.07e+02 | |
| AAE-Cox | 4.24e+28 | 2.71e+25 | 1.61e+34 | 7.46e+29 | 2.45e+25 | 7.92e+19 | 7.26e+14 | 3.55e+09 | 1.15e+05 | 4.18e+02 | 2.99e+02 | 2.84e+02 | |
| SAWAR | 3.30e+03 | 2.00e+03 | 1.23e+03 | 7.95e+02 | 5.43e+02 | 3.97e+02 | 3.38e+02 | 3.06e+02 | 2.87e+02 | 2.75e+02 | 2.72e+02 | 2.70e+02 | |
| TRACE | DRAFT | 5.09e+05 | 1.75e+05 | 6.15e+04 | 2.23e+04 | 8.34e+03 | 3.31e+03 | 1.48e+03 | 8.34e+02 | 6.09e+02 | 5.38e+02 | 5.27e+02 | 5.24e+02 |
| Noise | 2.14e+09 | 2.12e+08 | 2.24e+07 | 2.58e+06 | 3.25e+05 | 4.62e+04 | 7.86e+03 | 1.89e+03 | 7.98e+02 | 5.76e+02 | 5.46e+02 | 5.35e+02 | |
| FGSM | 4.68e+08 | 4.01e+07 | 3.85e+06 | 4.31e+05 | 5.82e+04 | 9.91e+03 | 2.30e+03 | 8.91e+02 | 6.03e+02 | 5.39e+02 | 5.31e+02 | 5.28e+02 | |
| PGD | 4.42e+07 | 6.51e+06 | 1.04e+06 | 1.82e+05 | 3.49e+04 | 7.47e+03 | 2.01e+03 | 8.54e+02 | 5.99e+02 | 5.39e+02 | 5.31e+02 | 5.27e+02 | |
| AAE-Cox | 1.06e+24 | 1.53e+21 | 2.08e+18 | 2.78e+15 | 2.40e+12 | 2.92e+09 | 5.29e+06 | 2.86e+04 | 9.75e+02 | 5.41e+02 | 5.31e+02 | 5.29e+02 | |
| SAWAR | 1.49e+03 | 1.01e+03 | 7.74e+02 | 6.57e+02 | 5.98e+02 | 5.67e+02 | 5.48e+02 | 5.35e+02 | 5.28e+02 | 5.24e+02 | 5.24e+02 | 5.24e+02 | |
| dataDIVAT1 | DRAFT | 2.97e+04 | 1.56e+04 | 8.36e+03 | 4.56e+03 | 2.59e+03 | 1.59e+03 | 1.11e+03 | 8.83e+02 | 7.91e+02 | 7.56e+02 | 7.51e+02 | 7.50e+02 |
| Noise | 1.60e+10 | 1.12e+09 | 8.30e+07 | 6.62e+06 | 5.83e+05 | 5.99e+04 | 8.13e+03 | 1.87e+03 | 9.38e+02 | 7.78e+02 | 7.61e+02 | 7.59e+02 | |
| FGSM | 3.25e+08 | 3.59e+07 | 4.13e+06 | 5.10e+05 | 7.07e+04 | 1.17e+04 | 2.66e+03 | 1.10e+03 | 8.11e+02 | 7.58e+02 | 7.53e+02 | 7.52e+02 | |
| PGD | 1.22e+08 | 1.55e+07 | 2.06e+06 | 2.92e+05 | 4.56e+04 | 8.39e+03 | 2.15e+03 | 1.01e+03 | 7.94e+02 | 7.55e+02 | 7.52e+02 | 7.51e+02 | |
| AAE-Cox | 9.96e+18 | 6.79e+16 | 3.12e+14 | 1.88e+12 | 1.37e+10 | 1.14e+08 | 1.05e+06 | 1.57e+04 | 1.04e+03 | 7.50e+02 | 7.45e+02 | 7.45e+02 | |
| SAWAR | 7.89e+02 | 7.79e+02 | 7.70e+02 | 7.63e+02 | 7.58e+02 | 7.54e+02 | 7.51e+02 | 7.49e+02 | 7.47e+02 | 7.46e+02 | 7.46e+02 | 7.46e+02 | |
| flchain | DRAFT | 1.57e+22 | 5.59e+19 | 2.03e+17 | 7.97e+14 | 3.37e+12 | 1.52e+10 | 7.49e+07 | 5.25e+05 | 1.45e+04 | 1.95e+03 | 1.25e+03 | 1.10e+03 |
| Noise | 2.69e+32 | 2.90e+28 | 3.20e+24 | 5.94e+33 | 2.08e+27 | 7.50e+20 | 3.25e+14 | 2.31e+08 | 6.25e+04 | 2.40e+03 | 1.47e+03 | 1.24e+03 | |
| FGSM | 5.19e+19 | 7.59e+16 | 1.21e+14 | 2.16e+11 | 7.32e+08 | 9.31e+06 | 2.85e+05 | 1.58e+04 | 2.16e+03 | 1.18e+03 | 1.12e+03 | 1.10e+03 | |
| PGD | 3.10e+15 | 1.18e+13 | 6.11e+10 | 6.20e+08 | 1.63e+07 | 6.61e+05 | 3.81e+04 | 4.41e+03 | 1.43e+03 | 1.12e+03 | 1.09e+03 | 1.09e+03 | |
| AAE-Cox | nan | nan | nan | nan | nan | nan | nan | nan | 4.02e+33 | 2.23e+10 | 1.11e+03 | 1.08e+03 | |
| SAWAR | 5.84e+05 | 3.37e+04 | 4.04e+03 | 1.74e+03 | 1.31e+03 | 1.18e+03 | 1.13e+03 | 1.11e+03 | 1.10e+03 | 1.09e+03 | 1.09e+03 | 1.09e+03 | |
| prostate | DRAFT | 5.92e+05 | 2.18e+05 | 8.11e+04 | 3.04e+04 | 1.16e+04 | 4.49e+03 | 1.83e+03 | 8.39e+02 | 4.86e+02 | 3.71e+02 | 3.48e+02 | 3.37e+02 |
| Noise | 6.55e+23 | 1.63e+21 | 4.14e+18 | 1.02e+16 | 2.79e+13 | 8.22e+10 | 2.79e+08 | 1.36e+06 | 3.02e+04 | 3.29e+03 | 1.53e+03 | 9.01e+02 | |
| FGSM | 5.24e+15 | 9.63e+13 | 1.84e+12 | 3.69e+10 | 7.62e+08 | 1.69e+07 | 4.35e+05 | 1.56e+04 | 1.19e+03 | 4.22e+02 | 3.76e+02 | 3.63e+02 | |
| PGD | 1.20e+13 | 4.77e+11 | 1.96e+10 | 8.28e+08 | 3.66e+07 | 1.75e+06 | 9.45e+04 | 6.42e+03 | 8.39e+02 | 3.99e+02 | 3.67e+02 | 3.57e+02 | |
| AAE-Cox | 1.53e+20 | 9.61e+17 | 6.16e+15 | 4.00e+13 | 2.65e+11 | 1.87e+09 | 1.33e+07 | 1.16e+05 | 1.81e+03 | 3.47e+02 | 3.33e+02 | 3.31e+02 | |
| SAWAR | 6.21e+02 | 5.06e+02 | 4.28e+02 | 3.91e+02 | 3.67e+02 | 3.55e+02 | 3.47e+02 | 3.41e+02 | 3.37e+02 | 3.35e+02 | 3.34e+02 | 3.33e+02 | |
| retinopathy | DRAFT | 1.21e+04 | 6.19e+03 | 3.21e+03 | 1.68e+03 | 9.15e+02 | 5.24e+02 | 3.31e+02 | 2.39e+02 | 1.96e+02 | 1.76e+02 | 1.71e+02 | 1.69e+02 |
| Noise | 1.65e+04 | 8.10e+03 | 4.01e+03 | 2.03e+03 | 1.05e+03 | 5.83e+02 | 3.54e+02 | 2.47e+02 | 1.99e+02 | 1.77e+02 | 1.72e+02 | 1.69e+02 | |
| FGSM | 7.62e+03 | 4.03e+03 | 2.15e+03 | 1.18e+03 | 6.72e+02 | 4.10e+02 | 2.80e+02 | 2.16e+02 | 1.86e+02 | 1.73e+02 | 1.70e+02 | 1.69e+02 | |
| PGD | 7.32e+03 | 3.87e+03 | 2.08e+03 | 1.14e+03 | 6.55e+02 | 4.03e+02 | 2.75e+02 | 2.14e+02 | 1.85e+02 | 1.73e+02 | 1.70e+02 | 1.69e+02 | |
| AAE-Cox | 1.32e+13 | 3.68e+11 | 1.03e+10 | 2.98e+08 | 8.67e+06 | 2.82e+05 | 1.17e+04 | 8.14e+02 | 1.98e+02 | 1.68e+02 | 1.67e+02 | 1.67e+02 | |
| SAWAR | 6.10e+02 | 4.69e+02 | 3.69e+02 | 3.00e+02 | 2.51e+02 | 2.18e+02 | 1.96e+02 | 1.82e+02 | 1.74e+02 | 1.69e+02 | 1.68e+02 | 1.68e+02 | |
| stagec | DRAFT | 1.17e+04 | 5.05e+03 | 2.20e+03 | 9.79e+02 | 4.44e+02 | 2.10e+02 | 1.07e+02 | 6.47e+01 | 4.80e+01 | 4.24e+01 | 4.15e+01 | 4.13e+01 |
| Noise | 6.21e+04 | 2.17e+04 | 7.71e+03 | 2.72e+03 | 1.01e+03 | 3.86e+02 | 1.63e+02 | 8.11e+01 | 5.29e+01 | 4.44e+01 | 4.31e+01 | 4.27e+01 | |
| FGSM | 1.02e+04 | 4.31e+03 | 1.80e+03 | 7.90e+02 | 3.54e+02 | 1.69e+02 | 9.18e+01 | 5.99e+01 | 4.77e+01 | 4.40e+01 | 4.35e+01 | 4.36e+01 | |
| PGD | 8.55e+03 | 3.69e+03 | 1.60e+03 | 7.18e+02 | 3.32e+02 | 1.62e+02 | 8.89e+01 | 5.88e+01 | 4.73e+01 | 4.38e+01 | 4.34e+01 | 4.35e+01 | |
| AAE-Cox | 1.32e+20 | 6.14e+17 | 2.58e+15 | 1.21e+13 | 5.37e+10 | 2.81e+08 | 1.53e+06 | 1.01e+04 | 1.53e+02 | 4.20e+01 | 4.10e+01 | 4.09e+01 | |
| SAWAR | 2.52e+02 | 1.75e+02 | 1.26e+02 | 9.40e+01 | 7.32e+01 | 5.96e+01 | 5.14e+01 | 4.61e+01 | 4.28e+01 | 4.09e+01 | 4.04e+01 | 4.01e+01 | |
| zinc | DRAFT | 1.34e+06 | 3.43e+05 | 8.95e+04 | 2.36e+04 | 6.49e+03 | 1.84e+03 | 5.72e+02 | 2.12e+02 | 1.11e+02 | 8.33e+01 | 7.84e+01 | 7.64e+01 |
| Noise | 1.01e+07 | 2.02e+06 | 4.01e+05 | 8.52e+04 | 1.86e+04 | 4.26e+03 | 1.05e+03 | 3.06e+02 | 1.28e+02 | 8.63e+01 | 8.05e+01 | 7.86e+01 | |
| FGSM | 3.49e+05 | 9.39e+04 | 2.69e+04 | 7.98e+03 | 2.42e+03 | 8.00e+02 | 3.01e+02 | 1.41e+02 | 9.24e+01 | 7.95e+01 | 7.82e+01 | 7.82e+01 | |
| PGD | 4.28e+05 | 1.09e+05 | 2.89e+04 | 7.84e+03 | 2.37e+03 | 7.72e+02 | 2.87e+02 | 1.36e+02 | 9.09e+01 | 7.94e+01 | 7.83e+01 | 7.85e+01 | |
| AAE-Cox | 1.15e+24 | 1.54e+21 | 4.97e+18 | 8.90e+15 | 1.88e+13 | 2.84e+10 | 5.64e+07 | 1.63e+05 | 7.39e+02 | 8.21e+01 | 7.79e+01 | 7.76e+01 | |
| SAWAR | 5.06e+02 | 3.35e+02 | 2.31e+02 | 1.67e+02 | 1.28e+02 | 1.04e+02 | 9.02e+01 | 8.23e+01 | 7.82e+01 | 7.68e+01 | 7.68e+01 | 7.72e+01 |
Appendix D Concordance Index FGSM Perturbation
We find that as the -perturbation magnitude increases from 0 to 1 for the FGSM adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Concordance Index metric from the DRAFT training objective to SAWAR training objective is shown in Table IX (where higher percentage change is better). We note that for very large , since our data is standard normalized all methods begin to fail.
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | |
| 39.82 | 59.68 | 72.41 | 86.66 | 105.06 | 136.98 | 168.31 | 178.41 | 116.94 | 26.54 | 12.48 | 1.6 |
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | ||
| Dataset | Algorithm | ||||||||||||
| Aids2 | DRAFT | 0.232 | 0.233 | 0.235 | 0.238 | 0.243 | 0.25 | 0.259 | 0.274 | 0.3 | 0.376 | 0.458 | 0.57 |
| Noise | 0.23 | 0.231 | 0.233 | 0.236 | 0.24 | 0.249 | 0.26 | 0.277 | 0.306 | 0.386 | 0.465 | 0.569 | |
| FGSM | 0.356 | 0.363 | 0.375 | 0.389 | 0.403 | 0.421 | 0.439 | 0.459 | 0.477 | 0.508 | 0.531 | 0.566 | |
| PGD | 0.326 | 0.336 | 0.35 | 0.367 | 0.385 | 0.409 | 0.434 | 0.46 | 0.486 | 0.515 | 0.534 | 0.567 | |
| AAE-Cox | 0.241 | 0.246 | 0.253 | 0.261 | 0.271 | 0.282 | 0.299 | 0.332 | 0.39 | 0.474 | 0.523 | 0.573 | |
| SAWAR | 0.259 | 0.265 | 0.276 | 0.289 | 0.306 | 0.327 | 0.352 | 0.385 | 0.43 | 0.493 | 0.53 | 0.569 | |
| Framingham | DRAFT | 0.142 | 0.143 | 0.143 | 0.144 | 0.144 | 0.145 | 0.148 | 0.168 | 0.257 | 0.468 | 0.6 | 0.722 |
| Noise | 0.143 | 0.144 | 0.144 | 0.145 | 0.145 | 0.146 | 0.15 | 0.173 | 0.265 | 0.473 | 0.601 | 0.719 | |
| FGSM | 0.149 | 0.152 | 0.156 | 0.164 | 0.18 | 0.208 | 0.257 | 0.332 | 0.437 | 0.57 | 0.643 | 0.715 | |
| PGD | 0.181 | 0.189 | 0.2 | 0.218 | 0.245 | 0.285 | 0.34 | 0.411 | 0.5 | 0.603 | 0.66 | 0.717 | |
| AAE-Cox | 0.148 | 0.152 | 0.162 | 0.181 | 0.216 | 0.269 | 0.34 | 0.427 | 0.53 | 0.636 | 0.687 | 0.733 | |
| SAWAR | 0.149 | 0.158 | 0.178 | 0.213 | 0.268 | 0.34 | 0.417 | 0.5 | 0.584 | 0.665 | 0.702 | 0.737 | |
| LeukSurv | DRAFT | 0.378 | 0.38 | 0.381 | 0.383 | 0.387 | 0.394 | 0.407 | 0.432 | 0.475 | 0.544 | 0.587 | 0.633 |
| Noise | 0.39 | 0.393 | 0.397 | 0.401 | 0.406 | 0.411 | 0.417 | 0.428 | 0.453 | 0.507 | 0.556 | 0.624 | |
| FGSM | 0.397 | 0.401 | 0.405 | 0.41 | 0.418 | 0.427 | 0.443 | 0.466 | 0.499 | 0.554 | 0.59 | 0.635 | |
| PGD | 0.399 | 0.402 | 0.407 | 0.411 | 0.419 | 0.431 | 0.447 | 0.473 | 0.508 | 0.562 | 0.597 | 0.638 | |
| AAE-Cox | 0.376 | 0.377 | 0.379 | 0.382 | 0.388 | 0.4 | 0.42 | 0.453 | 0.506 | 0.575 | 0.615 | 0.656 | |
| SAWAR | 0.392 | 0.399 | 0.409 | 0.423 | 0.442 | 0.465 | 0.494 | 0.529 | 0.574 | 0.623 | 0.649 | 0.673 | |
| TRACE | DRAFT | 0.219 | 0.221 | 0.225 | 0.232 | 0.247 | 0.279 | 0.334 | 0.415 | 0.524 | 0.642 | 0.696 | 0.745 |
| Noise | 0.234 | 0.236 | 0.24 | 0.246 | 0.259 | 0.286 | 0.336 | 0.414 | 0.52 | 0.639 | 0.695 | 0.745 | |
| FGSM | 0.29 | 0.311 | 0.336 | 0.369 | 0.408 | 0.454 | 0.507 | 0.565 | 0.625 | 0.685 | 0.715 | 0.744 | |
| PGD | 0.304 | 0.326 | 0.353 | 0.388 | 0.428 | 0.475 | 0.525 | 0.58 | 0.635 | 0.691 | 0.718 | 0.745 | |
| AAE-Cox | 0.251 | 0.273 | 0.302 | 0.339 | 0.384 | 0.438 | 0.498 | 0.563 | 0.63 | 0.692 | 0.721 | 0.747 | |
| SAWAR | 0.287 | 0.32 | 0.363 | 0.411 | 0.464 | 0.519 | 0.57 | 0.62 | 0.665 | 0.708 | 0.729 | 0.747 | |
| dataDIVAT1 | DRAFT | 0.097 | 0.098 | 0.099 | 0.1 | 0.101 | 0.103 | 0.105 | 0.122 | 0.214 | 0.414 | 0.53 | 0.655 |
| Noise | 0.094 | 0.094 | 0.095 | 0.096 | 0.097 | 0.099 | 0.101 | 0.111 | 0.178 | 0.366 | 0.494 | 0.635 | |
| FGSM | 0.114 | 0.116 | 0.12 | 0.126 | 0.138 | 0.161 | 0.203 | 0.27 | 0.362 | 0.482 | 0.559 | 0.646 | |
| PGD | 0.137 | 0.143 | 0.152 | 0.167 | 0.188 | 0.22 | 0.265 | 0.33 | 0.409 | 0.511 | 0.574 | 0.648 | |
| AAE-Cox | 0.111 | 0.114 | 0.122 | 0.139 | 0.167 | 0.209 | 0.268 | 0.347 | 0.443 | 0.551 | 0.607 | 0.663 | |
| SAWAR | 0.156 | 0.201 | 0.253 | 0.311 | 0.368 | 0.422 | 0.461 | 0.506 | 0.555 | 0.609 | 0.637 | 0.67 | |
| flchain | DRAFT | 0.152 | 0.153 | 0.154 | 0.155 | 0.156 | 0.157 | 0.159 | 0.164 | 0.197 | 0.895 | 0.903 | 0.921 |
| Noise | 0.166 | 0.167 | 0.169 | 0.171 | 0.175 | 0.182 | 0.213 | 0.333 | 0.809 | 0.9 | 0.904 | 0.918 | |
| FGSM | 0.507 | 0.628 | 0.727 | 0.8 | 0.854 | 0.883 | 0.898 | 0.902 | 0.906 | 0.911 | 0.915 | 0.922 | |
| PGD | 0.539 | 0.717 | 0.824 | 0.876 | 0.896 | 0.9 | 0.903 | 0.905 | 0.908 | 0.913 | 0.917 | 0.923 | |
| AAE-Cox | 0.172 | 0.188 | 0.218 | 0.275 | 0.372 | 0.508 | 0.665 | 0.81 | 0.899 | 0.918 | 0.922 | 0.926 | |
| SAWAR | 0.628 | 0.831 | 0.894 | 0.9 | 0.904 | 0.907 | 0.909 | 0.912 | 0.916 | 0.921 | 0.924 | 0.927 | |
| prostate | DRAFT | 0.308 | 0.311 | 0.314 | 0.318 | 0.321 | 0.326 | 0.333 | 0.339 | 0.357 | 0.435 | 0.531 | 0.668 |
| Noise | 0.305 | 0.305 | 0.306 | 0.307 | 0.309 | 0.314 | 0.32 | 0.328 | 0.344 | 0.394 | 0.457 | 0.562 | |
| FGSM | 0.293 | 0.295 | 0.297 | 0.3 | 0.304 | 0.308 | 0.312 | 0.32 | 0.343 | 0.407 | 0.474 | 0.571 | |
| PGD | 0.299 | 0.301 | 0.303 | 0.305 | 0.308 | 0.312 | 0.318 | 0.331 | 0.36 | 0.431 | 0.502 | 0.588 | |
| AAE-Cox | 0.308 | 0.308 | 0.311 | 0.315 | 0.325 | 0.342 | 0.375 | 0.426 | 0.499 | 0.597 | 0.645 | 0.691 | |
| SAWAR | 0.288 | 0.292 | 0.299 | 0.31 | 0.326 | 0.349 | 0.384 | 0.432 | 0.499 | 0.579 | 0.618 | 0.657 | |
| retinopathy | DRAFT | 0.139 | 0.141 | 0.145 | 0.15 | 0.152 | 0.152 | 0.153 | 0.158 | 0.209 | 0.425 | 0.554 | 0.669 |
| Noise | 0.134 | 0.137 | 0.138 | 0.138 | 0.138 | 0.138 | 0.138 | 0.15 | 0.245 | 0.456 | 0.57 | 0.668 | |
| FGSM | 0.131 | 0.133 | 0.135 | 0.135 | 0.135 | 0.135 | 0.137 | 0.155 | 0.255 | 0.456 | 0.561 | 0.659 | |
| PGD | 0.13 | 0.13 | 0.131 | 0.131 | 0.131 | 0.131 | 0.134 | 0.154 | 0.254 | 0.456 | 0.56 | 0.657 | |
| AAE-Cox | 0.129 | 0.129 | 0.13 | 0.137 | 0.157 | 0.196 | 0.247 | 0.337 | 0.444 | 0.541 | 0.595 | 0.648 | |
| SAWAR | 0.131 | 0.13 | 0.13 | 0.131 | 0.135 | 0.149 | 0.178 | 0.251 | 0.371 | 0.51 | 0.584 | 0.647 | |
| stagec | DRAFT | 0.137 | 0.136 | 0.136 | 0.132 | 0.136 | 0.135 | 0.151 | 0.174 | 0.23 | 0.34 | 0.407 | 0.512 |
| Noise | 0.133 | 0.133 | 0.132 | 0.132 | 0.136 | 0.144 | 0.16 | 0.187 | 0.261 | 0.362 | 0.442 | 0.555 | |
| FGSM | 0.12 | 0.124 | 0.127 | 0.134 | 0.146 | 0.147 | 0.157 | 0.184 | 0.255 | 0.33 | 0.417 | 0.51 | |
| PGD | 0.115 | 0.117 | 0.122 | 0.128 | 0.141 | 0.144 | 0.154 | 0.181 | 0.259 | 0.327 | 0.412 | 0.505 | |
| AAE-Cox | 0.125 | 0.121 | 0.131 | 0.141 | 0.163 | 0.202 | 0.241 | 0.296 | 0.352 | 0.426 | 0.485 | 0.54 | |
| SAWAR | 0.123 | 0.124 | 0.127 | 0.139 | 0.149 | 0.179 | 0.239 | 0.281 | 0.347 | 0.425 | 0.486 | 0.558 | |
| zinc | DRAFT | 0.057 | 0.057 | 0.057 | 0.057 | 0.057 | 0.057 | 0.057 | 0.068 | 0.155 | 0.485 | 0.646 | 0.77 |
| Noise | 0.053 | 0.053 | 0.054 | 0.053 | 0.054 | 0.057 | 0.062 | 0.095 | 0.248 | 0.547 | 0.668 | 0.776 | |
| FGSM | 0.055 | 0.056 | 0.056 | 0.058 | 0.06 | 0.07 | 0.095 | 0.199 | 0.404 | 0.61 | 0.705 | 0.783 | |
| PGD | 0.056 | 0.056 | 0.057 | 0.058 | 0.064 | 0.078 | 0.113 | 0.22 | 0.42 | 0.615 | 0.706 | 0.781 | |
| AAE-Cox | 0.062 | 0.065 | 0.079 | 0.101 | 0.162 | 0.227 | 0.291 | 0.409 | 0.551 | 0.655 | 0.708 | 0.759 | |
| SAWAR | 0.055 | 0.056 | 0.058 | 0.073 | 0.111 | 0.213 | 0.325 | 0.446 | 0.556 | 0.679 | 0.729 | 0.774 |
Appendix E Integrated Brier Score FGSM Perturbation
We find that as the -perturbation magnitude increases from 0 to 1 for the FGSM adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Scores metric from the DRAFT training objective to SAWAR training objective is shown in Table XI (where lower percentage change is better). We note that for very large , since our data is standard normalized all methods begin to fail.
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | |
| -44.57 | -46.37 | -47.93 | -49.06 | -49.43 | -48.69 | -46.35 | -41.82 | -33.56 | -20.63 | -11.37 | -0.65 |
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | ||
| Dataset | Algorithm | ||||||||||||
| Aids2 | DRAFT | 0.232 | 0.233 | 0.235 | 0.238 | 0.243 | 0.25 | 0.259 | 0.274 | 0.3 | 0.376 | 0.458 | 0.57 |
| Noise | 0.23 | 0.231 | 0.233 | 0.236 | 0.24 | 0.249 | 0.26 | 0.277 | 0.306 | 0.386 | 0.465 | 0.569 | |
| FGSM | 0.356 | 0.363 | 0.375 | 0.389 | 0.403 | 0.421 | 0.439 | 0.459 | 0.477 | 0.508 | 0.531 | 0.566 | |
| PGD | 0.326 | 0.336 | 0.35 | 0.367 | 0.385 | 0.409 | 0.434 | 0.46 | 0.486 | 0.515 | 0.534 | 0.567 | |
| AAE-Cox | 0.241 | 0.246 | 0.253 | 0.261 | 0.271 | 0.282 | 0.299 | 0.332 | 0.39 | 0.474 | 0.523 | 0.573 | |
| SAWAR | 0.259 | 0.265 | 0.276 | 0.289 | 0.306 | 0.327 | 0.352 | 0.385 | 0.43 | 0.493 | 0.53 | 0.569 | |
| Framingham | DRAFT | 0.142 | 0.143 | 0.143 | 0.144 | 0.144 | 0.145 | 0.148 | 0.168 | 0.257 | 0.468 | 0.6 | 0.722 |
| Noise | 0.143 | 0.144 | 0.144 | 0.145 | 0.145 | 0.146 | 0.15 | 0.173 | 0.265 | 0.473 | 0.601 | 0.719 | |
| FGSM | 0.149 | 0.152 | 0.156 | 0.164 | 0.18 | 0.208 | 0.257 | 0.332 | 0.437 | 0.57 | 0.643 | 0.715 | |
| PGD | 0.181 | 0.189 | 0.2 | 0.218 | 0.245 | 0.285 | 0.34 | 0.411 | 0.5 | 0.603 | 0.66 | 0.717 | |
| AAE-Cox | 0.148 | 0.152 | 0.162 | 0.181 | 0.216 | 0.269 | 0.34 | 0.427 | 0.53 | 0.636 | 0.687 | 0.733 | |
| SAWAR | 0.149 | 0.158 | 0.178 | 0.213 | 0.268 | 0.34 | 0.417 | 0.5 | 0.584 | 0.665 | 0.702 | 0.737 | |
| LeukSurv | DRAFT | 0.378 | 0.38 | 0.381 | 0.383 | 0.387 | 0.394 | 0.407 | 0.432 | 0.475 | 0.544 | 0.587 | 0.633 |
| Noise | 0.39 | 0.393 | 0.397 | 0.401 | 0.406 | 0.411 | 0.417 | 0.428 | 0.453 | 0.507 | 0.556 | 0.624 | |
| FGSM | 0.397 | 0.401 | 0.405 | 0.41 | 0.418 | 0.427 | 0.443 | 0.466 | 0.499 | 0.554 | 0.59 | 0.635 | |
| PGD | 0.399 | 0.402 | 0.407 | 0.411 | 0.419 | 0.431 | 0.447 | 0.473 | 0.508 | 0.562 | 0.597 | 0.638 | |
| AAE-Cox | 0.376 | 0.377 | 0.379 | 0.382 | 0.388 | 0.4 | 0.42 | 0.453 | 0.506 | 0.575 | 0.615 | 0.656 | |
| SAWAR | 0.392 | 0.399 | 0.409 | 0.423 | 0.442 | 0.465 | 0.494 | 0.529 | 0.574 | 0.623 | 0.649 | 0.673 | |
| TRACE | DRAFT | 0.219 | 0.221 | 0.225 | 0.232 | 0.247 | 0.279 | 0.334 | 0.415 | 0.524 | 0.642 | 0.696 | 0.745 |
| Noise | 0.234 | 0.236 | 0.24 | 0.246 | 0.259 | 0.286 | 0.336 | 0.414 | 0.52 | 0.639 | 0.695 | 0.745 | |
| FGSM | 0.29 | 0.311 | 0.336 | 0.369 | 0.408 | 0.454 | 0.507 | 0.565 | 0.625 | 0.685 | 0.715 | 0.744 | |
| PGD | 0.304 | 0.326 | 0.353 | 0.388 | 0.428 | 0.475 | 0.525 | 0.58 | 0.635 | 0.691 | 0.718 | 0.745 | |
| AAE-Cox | 0.251 | 0.273 | 0.302 | 0.339 | 0.384 | 0.438 | 0.498 | 0.563 | 0.63 | 0.692 | 0.721 | 0.747 | |
| SAWAR | 0.287 | 0.32 | 0.363 | 0.411 | 0.464 | 0.519 | 0.57 | 0.62 | 0.665 | 0.708 | 0.729 | 0.747 | |
| dataDIVAT1 | DRAFT | 0.097 | 0.098 | 0.099 | 0.1 | 0.101 | 0.103 | 0.105 | 0.122 | 0.214 | 0.414 | 0.53 | 0.655 |
| Noise | 0.094 | 0.094 | 0.095 | 0.096 | 0.097 | 0.099 | 0.101 | 0.111 | 0.178 | 0.366 | 0.494 | 0.635 | |
| FGSM | 0.114 | 0.116 | 0.12 | 0.126 | 0.138 | 0.161 | 0.203 | 0.27 | 0.362 | 0.482 | 0.559 | 0.646 | |
| PGD | 0.137 | 0.143 | 0.152 | 0.167 | 0.188 | 0.22 | 0.265 | 0.33 | 0.409 | 0.511 | 0.574 | 0.648 | |
| AAE-Cox | 0.111 | 0.114 | 0.122 | 0.139 | 0.167 | 0.209 | 0.268 | 0.347 | 0.443 | 0.551 | 0.607 | 0.663 | |
| SAWAR | 0.156 | 0.201 | 0.253 | 0.311 | 0.368 | 0.422 | 0.461 | 0.506 | 0.555 | 0.609 | 0.637 | 0.67 | |
| flchain | DRAFT | 0.152 | 0.153 | 0.154 | 0.155 | 0.156 | 0.157 | 0.159 | 0.164 | 0.197 | 0.895 | 0.903 | 0.921 |
| Noise | 0.166 | 0.167 | 0.169 | 0.171 | 0.175 | 0.182 | 0.213 | 0.333 | 0.809 | 0.9 | 0.904 | 0.918 | |
| FGSM | 0.507 | 0.628 | 0.727 | 0.8 | 0.854 | 0.883 | 0.898 | 0.902 | 0.906 | 0.911 | 0.915 | 0.922 | |
| PGD | 0.539 | 0.717 | 0.824 | 0.876 | 0.896 | 0.9 | 0.903 | 0.905 | 0.908 | 0.913 | 0.917 | 0.923 | |
| AAE-Cox | 0.172 | 0.188 | 0.218 | 0.275 | 0.372 | 0.508 | 0.665 | 0.81 | 0.899 | 0.918 | 0.922 | 0.926 | |
| SAWAR | 0.628 | 0.831 | 0.894 | 0.9 | 0.904 | 0.907 | 0.909 | 0.912 | 0.916 | 0.921 | 0.924 | 0.927 | |
| prostate | DRAFT | 0.308 | 0.311 | 0.314 | 0.318 | 0.321 | 0.326 | 0.333 | 0.339 | 0.357 | 0.435 | 0.531 | 0.668 |
| Noise | 0.305 | 0.305 | 0.306 | 0.307 | 0.309 | 0.314 | 0.32 | 0.328 | 0.344 | 0.394 | 0.457 | 0.562 | |
| FGSM | 0.293 | 0.295 | 0.297 | 0.3 | 0.304 | 0.308 | 0.312 | 0.32 | 0.343 | 0.407 | 0.474 | 0.571 | |
| PGD | 0.299 | 0.301 | 0.303 | 0.305 | 0.308 | 0.312 | 0.318 | 0.331 | 0.36 | 0.431 | 0.502 | 0.588 | |
| AAE-Cox | 0.308 | 0.308 | 0.311 | 0.315 | 0.325 | 0.342 | 0.375 | 0.426 | 0.499 | 0.597 | 0.645 | 0.691 | |
| SAWAR | 0.288 | 0.292 | 0.299 | 0.31 | 0.326 | 0.349 | 0.384 | 0.432 | 0.499 | 0.579 | 0.618 | 0.657 | |
| retinopathy | DRAFT | 0.139 | 0.141 | 0.145 | 0.15 | 0.152 | 0.152 | 0.153 | 0.158 | 0.209 | 0.425 | 0.554 | 0.669 |
| Noise | 0.134 | 0.137 | 0.138 | 0.138 | 0.138 | 0.138 | 0.138 | 0.15 | 0.245 | 0.456 | 0.57 | 0.668 | |
| FGSM | 0.131 | 0.133 | 0.135 | 0.135 | 0.135 | 0.135 | 0.137 | 0.155 | 0.255 | 0.456 | 0.561 | 0.659 | |
| PGD | 0.13 | 0.13 | 0.131 | 0.131 | 0.131 | 0.131 | 0.134 | 0.154 | 0.254 | 0.456 | 0.56 | 0.657 | |
| AAE-Cox | 0.129 | 0.129 | 0.13 | 0.137 | 0.157 | 0.196 | 0.247 | 0.337 | 0.444 | 0.541 | 0.595 | 0.648 | |
| SAWAR | 0.131 | 0.13 | 0.13 | 0.131 | 0.135 | 0.149 | 0.178 | 0.251 | 0.371 | 0.51 | 0.584 | 0.647 | |
| stagec | DRAFT | 0.137 | 0.136 | 0.136 | 0.132 | 0.136 | 0.135 | 0.151 | 0.174 | 0.23 | 0.34 | 0.407 | 0.512 |
| Noise | 0.133 | 0.133 | 0.132 | 0.132 | 0.136 | 0.144 | 0.16 | 0.187 | 0.261 | 0.362 | 0.442 | 0.555 | |
| FGSM | 0.12 | 0.124 | 0.127 | 0.134 | 0.146 | 0.147 | 0.157 | 0.184 | 0.255 | 0.33 | 0.417 | 0.51 | |
| PGD | 0.115 | 0.117 | 0.122 | 0.128 | 0.141 | 0.144 | 0.154 | 0.181 | 0.259 | 0.327 | 0.412 | 0.505 | |
| AAE-Cox | 0.125 | 0.121 | 0.131 | 0.141 | 0.163 | 0.202 | 0.241 | 0.296 | 0.352 | 0.426 | 0.485 | 0.54 | |
| SAWAR | 0.123 | 0.124 | 0.127 | 0.139 | 0.149 | 0.179 | 0.239 | 0.281 | 0.347 | 0.425 | 0.486 | 0.558 | |
| zinc | DRAFT | 0.057 | 0.057 | 0.057 | 0.057 | 0.057 | 0.057 | 0.057 | 0.068 | 0.155 | 0.485 | 0.646 | 0.77 |
| Noise | 0.053 | 0.053 | 0.054 | 0.053 | 0.054 | 0.057 | 0.062 | 0.095 | 0.248 | 0.547 | 0.668 | 0.776 | |
| FGSM | 0.055 | 0.056 | 0.056 | 0.058 | 0.06 | 0.07 | 0.095 | 0.199 | 0.404 | 0.61 | 0.705 | 0.783 | |
| PGD | 0.056 | 0.056 | 0.057 | 0.058 | 0.064 | 0.078 | 0.113 | 0.22 | 0.42 | 0.615 | 0.706 | 0.781 | |
| AAE-Cox | 0.062 | 0.065 | 0.079 | 0.101 | 0.162 | 0.227 | 0.291 | 0.409 | 0.551 | 0.655 | 0.708 | 0.759 | |
| SAWAR | 0.055 | 0.056 | 0.058 | 0.073 | 0.111 | 0.213 | 0.325 | 0.446 | 0.556 | 0.679 | 0.729 | 0.774 |
Appendix F Negative Log Likelihood FGSM Perturbation
| 1.00 | 0.90 | 0.80 | 0.70 | 0.60 | 0.50 | 0.40 | 0.30 | 0.20 | 0.10 | 0.05 | 0.00 | ||
| Dataset | Algorithm | ||||||||||||
| Aids2 | DRAFT | 1.04e+03 | 9.95e+02 | 9.48e+02 | 9.01e+02 | 8.54e+02 | 8.06e+02 | 7.56e+02 | 7.04e+02 | 6.49e+02 | 5.92e+02 | 5.66e+02 | 5.41e+02 |
| Noise | 9.94e+02 | 9.53e+02 | 9.11e+02 | 8.69e+02 | 8.26e+02 | 7.82e+02 | 7.37e+02 | 6.89e+02 | 6.39e+02 | 5.88e+02 | 5.64e+02 | 5.41e+02 | |
| FGSM | 6.16e+02 | 6.10e+02 | 6.04e+02 | 5.97e+02 | 5.91e+02 | 5.84e+02 | 5.77e+02 | 5.70e+02 | 5.63e+02 | 5.54e+02 | 5.48e+02 | 5.41e+02 | |
| PGD | 6.20e+02 | 6.13e+02 | 6.05e+02 | 5.98e+02 | 5.91e+02 | 5.83e+02 | 5.76e+02 | 5.68e+02 | 5.60e+02 | 5.52e+02 | 5.47e+02 | 5.41e+02 | |
| AAE-Cox | 7.00e+02 | 6.83e+02 | 6.67e+02 | 6.51e+02 | 6.35e+02 | 6.19e+02 | 6.03e+02 | 5.87e+02 | 5.71e+02 | 5.55e+02 | 5.47e+02 | 5.39e+02 | |
| SAWAR | 5.84e+02 | 5.80e+02 | 5.76e+02 | 5.72e+02 | 5.68e+02 | 5.63e+02 | 5.59e+02 | 5.55e+02 | 5.50e+02 | 5.45e+02 | 5.43e+02 | 5.40e+02 | |
| Framingham | DRAFT | 4.23e+03 | 4.10e+03 | 3.91e+03 | 3.65e+03 | 3.33e+03 | 2.97e+03 | 2.61e+03 | 2.26e+03 | 1.95e+03 | 1.68e+03 | 1.57e+03 | 1.48e+03 |
| Noise | 4.00e+03 | 3.89e+03 | 3.73e+03 | 3.51e+03 | 3.23e+03 | 2.91e+03 | 2.57e+03 | 2.24e+03 | 1.94e+03 | 1.69e+03 | 1.58e+03 | 1.48e+03 | |
| FGSM | 2.11e+03 | 2.07e+03 | 2.03e+03 | 1.98e+03 | 1.93e+03 | 1.88e+03 | 1.82e+03 | 1.75e+03 | 1.67e+03 | 1.58e+03 | 1.53e+03 | 1.48e+03 | |
| PGD | 1.99e+03 | 1.95e+03 | 1.91e+03 | 1.87e+03 | 1.83e+03 | 1.79e+03 | 1.74e+03 | 1.68e+03 | 1.62e+03 | 1.56e+03 | 1.52e+03 | 1.48e+03 | |
| AAE-Cox | 2.38e+03 | 2.27e+03 | 2.16e+03 | 2.06e+03 | 1.96e+03 | 1.87e+03 | 1.78e+03 | 1.70e+03 | 1.62e+03 | 1.54e+03 | 1.51e+03 | 1.47e+03 | |
| SAWAR | 1.96e+03 | 1.90e+03 | 1.84e+03 | 1.79e+03 | 1.73e+03 | 1.68e+03 | 1.64e+03 | 1.59e+03 | 1.55e+03 | 1.51e+03 | 1.49e+03 | 1.48e+03 | |
| LeukSurv | DRAFT | 2.66e+03 | 2.16e+03 | 1.77e+03 | 1.45e+03 | 1.19e+03 | 9.73e+02 | 7.93e+02 | 6.41e+02 | 5.10e+02 | 3.99e+02 | 3.50e+02 | 3.07e+02 |
| Noise | 1.13e+06 | 5.84e+05 | 2.96e+05 | 1.52e+05 | 7.88e+04 | 4.05e+04 | 1.94e+04 | 9.21e+03 | 4.25e+03 | 1.89e+03 | 1.23e+03 | 7.75e+02 | |
| FGSM | 2.76e+03 | 2.33e+03 | 1.96e+03 | 1.65e+03 | 1.39e+03 | 1.16e+03 | 9.64e+02 | 7.82e+02 | 6.18e+02 | 4.64e+02 | 3.92e+02 | 3.25e+02 | |
| PGD | 1.79e+03 | 1.57e+03 | 1.38e+03 | 1.21e+03 | 1.05e+03 | 9.10e+02 | 7.76e+02 | 6.49e+02 | 5.30e+02 | 4.15e+02 | 3.60e+02 | 3.07e+02 | |
| AAE-Cox | 1.09e+03 | 9.83e+02 | 8.87e+02 | 7.97e+02 | 7.12e+02 | 6.32e+02 | 5.57e+02 | 4.84e+02 | 4.12e+02 | 3.43e+02 | 3.12e+02 | 2.84e+02 | |
| SAWAR | 5.40e+02 | 5.12e+02 | 4.85e+02 | 4.57e+02 | 4.29e+02 | 4.01e+02 | 3.73e+02 | 3.46e+02 | 3.20e+02 | 2.94e+02 | 2.82e+02 | 2.70e+02 | |
| TRACE | DRAFT | 2.31e+03 | 2.09e+03 | 1.87e+03 | 1.64e+03 | 1.43e+03 | 1.22e+03 | 1.03e+03 | 8.69e+02 | 7.29e+02 | 6.14e+02 | 5.66e+02 | 5.24e+02 |
| Noise | 4.55e+03 | 3.85e+03 | 3.20e+03 | 2.63e+03 | 2.14e+03 | 1.72e+03 | 1.36e+03 | 1.06e+03 | 8.29e+02 | 6.60e+02 | 5.93e+02 | 5.35e+02 | |
| FGSM | 1.23e+03 | 1.15e+03 | 1.07e+03 | 9.98e+02 | 9.26e+02 | 8.54e+02 | 7.82e+02 | 7.13e+02 | 6.47e+02 | 5.85e+02 | 5.56e+02 | 5.28e+02 | |
| PGD | 1.17e+03 | 1.10e+03 | 1.03e+03 | 9.59e+02 | 8.90e+02 | 8.21e+02 | 7.54e+02 | 6.93e+02 | 6.35e+02 | 5.80e+02 | 5.53e+02 | 5.27e+02 | |
| AAE-Cox | 1.07e+03 | 1.00e+03 | 9.38e+02 | 8.77e+02 | 8.19e+02 | 7.63e+02 | 7.09e+02 | 6.59e+02 | 6.11e+02 | 5.68e+02 | 5.48e+02 | 5.29e+02 | |
| SAWAR | 1.01e+03 | 9.32e+02 | 8.63e+02 | 8.01e+02 | 7.47e+02 | 6.99e+02 | 6.57e+02 | 6.19e+02 | 5.84e+02 | 5.53e+02 | 5.38e+02 | 5.24e+02 | |
| dataDIVAT1 | DRAFT | 1.78e+03 | 1.72e+03 | 1.64e+03 | 1.55e+03 | 1.44e+03 | 1.32e+03 | 1.20e+03 | 1.07e+03 | 9.55e+02 | 8.46e+02 | 7.96e+02 | 7.50e+02 |
| Noise | 2.02e+03 | 1.96e+03 | 1.88e+03 | 1.78e+03 | 1.65e+03 | 1.50e+03 | 1.35e+03 | 1.19e+03 | 1.03e+03 | 8.87e+02 | 8.20e+02 | 7.59e+02 | |
| FGSM | 1.09e+03 | 1.07e+03 | 1.05e+03 | 1.02e+03 | 1.00e+03 | 9.72e+02 | 9.40e+02 | 9.03e+02 | 8.59e+02 | 8.09e+02 | 7.81e+02 | 7.52e+02 | |
| PGD | 1.01e+03 | 9.95e+02 | 9.78e+02 | 9.60e+02 | 9.41e+02 | 9.19e+02 | 8.94e+02 | 8.65e+02 | 8.32e+02 | 7.95e+02 | 7.74e+02 | 7.51e+02 | |
| AAE-Cox | 1.25e+03 | 1.18e+03 | 1.12e+03 | 1.06e+03 | 1.01e+03 | 9.55e+02 | 9.09e+02 | 8.65e+02 | 8.23e+02 | 7.83e+02 | 7.64e+02 | 7.45e+02 | |
| SAWAR | 9.24e+02 | 9.01e+02 | 8.79e+02 | 8.59e+02 | 8.39e+02 | 8.20e+02 | 8.04e+02 | 7.89e+02 | 7.74e+02 | 7.60e+02 | 7.53e+02 | 7.46e+02 | |
| flchain | DRAFT | 2.56e+04 | 2.36e+04 | 2.15e+04 | 1.95e+04 | 1.73e+04 | 1.47e+04 | 1.11e+04 | 6.38e+03 | 2.91e+03 | 1.59e+03 | 1.27e+03 | 1.10e+03 |
| Noise | 1.62e+05 | 1.27e+05 | 8.24e+04 | 4.85e+04 | 2.97e+04 | 1.81e+04 | 9.84e+03 | 5.14e+03 | 2.95e+03 | 1.80e+03 | 1.45e+03 | 1.24e+03 | |
| FGSM | 2.49e+03 | 2.17e+03 | 1.87e+03 | 1.62e+03 | 1.44e+03 | 1.34e+03 | 1.28e+03 | 1.23e+03 | 1.19e+03 | 1.14e+03 | 1.12e+03 | 1.10e+03 | |
| PGD | 2.07e+03 | 1.76e+03 | 1.53e+03 | 1.37e+03 | 1.28e+03 | 1.22e+03 | 1.19e+03 | 1.16e+03 | 1.14e+03 | 1.11e+03 | 1.10e+03 | 1.09e+03 | |
| AAE-Cox | 2.98e+03 | 2.85e+03 | 2.73e+03 | 2.62e+03 | 2.50e+03 | 2.38e+03 | 2.23e+03 | 1.97e+03 | 1.47e+03 | 1.12e+03 | 1.09e+03 | 1.08e+03 | |
| SAWAR | 2.18e+03 | 1.83e+03 | 1.55e+03 | 1.33e+03 | 1.20e+03 | 1.14e+03 | 1.13e+03 | 1.11e+03 | 1.10e+03 | 1.09e+03 | 1.09e+03 | 1.09e+03 | |
| prostate | DRAFT | 1.50e+03 | 1.44e+03 | 1.36e+03 | 1.24e+03 | 1.10e+03 | 9.38e+02 | 7.70e+02 | 6.16e+02 | 4.90e+02 | 3.97e+02 | 3.63e+02 | 3.37e+02 |
| Noise | 1.88e+06 | 1.21e+06 | 6.72e+05 | 3.64e+05 | 1.97e+05 | 9.86e+04 | 4.32e+04 | 1.69e+04 | 6.34e+03 | 2.38e+03 | 1.46e+03 | 9.01e+02 | |
| FGSM | 8.51e+02 | 8.48e+02 | 8.41e+02 | 8.25e+02 | 7.97e+02 | 7.47e+02 | 6.78e+02 | 5.92e+02 | 5.05e+02 | 4.25e+02 | 3.92e+02 | 3.63e+02 | |
| PGD | 7.46e+02 | 7.32e+02 | 7.17e+02 | 6.97e+02 | 6.69e+02 | 6.30e+02 | 5.79e+02 | 5.22e+02 | 4.63e+02 | 4.05e+02 | 3.79e+02 | 3.57e+02 | |
| AAE-Cox | 4.64e+02 | 4.47e+02 | 4.31e+02 | 4.16e+02 | 4.02e+02 | 3.89e+02 | 3.77e+02 | 3.64e+02 | 3.52e+02 | 3.41e+02 | 3.36e+02 | 3.31e+02 | |
| SAWAR | 4.40e+02 | 4.28e+02 | 4.15e+02 | 4.04e+02 | 3.92e+02 | 3.81e+02 | 3.70e+02 | 3.60e+02 | 3.51e+02 | 3.42e+02 | 3.37e+02 | 3.33e+02 | |
| retinopathy | DRAFT | 5.04e+02 | 4.93e+02 | 4.70e+02 | 4.35e+02 | 3.93e+02 | 3.48e+02 | 3.02e+02 | 2.59e+02 | 2.22e+02 | 1.92e+02 | 1.80e+02 | 1.69e+02 |
| Noise | 5.47e+02 | 5.29e+02 | 4.96e+02 | 4.54e+02 | 4.04e+02 | 3.54e+02 | 3.05e+02 | 2.62e+02 | 2.24e+02 | 1.94e+02 | 1.81e+02 | 1.69e+02 | |
| FGSM | 4.16e+02 | 4.09e+02 | 3.93e+02 | 3.69e+02 | 3.39e+02 | 3.06e+02 | 2.73e+02 | 2.41e+02 | 2.13e+02 | 1.89e+02 | 1.79e+02 | 1.69e+02 | |
| PGD | 4.11e+02 | 4.03e+02 | 3.88e+02 | 3.65e+02 | 3.37e+02 | 3.05e+02 | 2.71e+02 | 2.40e+02 | 2.12e+02 | 1.89e+02 | 1.78e+02 | 1.69e+02 | |
| AAE-Cox | 2.27e+02 | 2.21e+02 | 2.14e+02 | 2.09e+02 | 2.03e+02 | 1.97e+02 | 1.91e+02 | 1.85e+02 | 1.79e+02 | 1.73e+02 | 1.70e+02 | 1.67e+02 | |
| SAWAR | 3.27e+02 | 3.14e+02 | 2.99e+02 | 2.82e+02 | 2.65e+02 | 2.47e+02 | 2.29e+02 | 2.12e+02 | 1.96e+02 | 1.81e+02 | 1.74e+02 | 1.68e+02 | |
| stagec | DRAFT | 1.38e+02 | 1.32e+02 | 1.24e+02 | 1.15e+02 | 1.04e+02 | 9.29e+01 | 8.09e+01 | 6.91e+01 | 5.82e+01 | 4.90e+01 | 4.49e+01 | 4.13e+01 |
| Noise | 1.52e+02 | 1.44e+02 | 1.35e+02 | 1.24e+02 | 1.12e+02 | 9.97e+01 | 8.67e+01 | 7.39e+01 | 6.18e+01 | 5.13e+01 | 4.68e+01 | 4.27e+01 | |
| FGSM | 1.04e+02 | 1.01e+02 | 9.63e+01 | 9.11e+01 | 8.52e+01 | 7.85e+01 | 7.12e+01 | 6.37e+01 | 5.64e+01 | 4.96e+01 | 4.65e+01 | 4.36e+01 | |
| PGD | 1.01e+02 | 9.73e+01 | 9.31e+01 | 8.82e+01 | 8.27e+01 | 7.65e+01 | 6.97e+01 | 6.27e+01 | 5.58e+01 | 4.93e+01 | 4.63e+01 | 4.35e+01 | |
| AAE-Cox | 8.00e+01 | 7.47e+01 | 6.99e+01 | 6.55e+01 | 6.15e+01 | 5.77e+01 | 5.42e+01 | 5.07e+01 | 4.74e+01 | 4.41e+01 | 4.25e+01 | 4.09e+01 | |
| SAWAR | 9.46e+01 | 8.98e+01 | 8.41e+01 | 7.80e+01 | 7.17e+01 | 6.57e+01 | 6.03e+01 | 5.48e+01 | 4.95e+01 | 4.45e+01 | 4.23e+01 | 4.01e+01 | |
| zinc | DRAFT | 5.01e+02 | 4.94e+02 | 4.77e+02 | 4.45e+02 | 3.93e+02 | 3.19e+02 | 2.39e+02 | 1.73e+02 | 1.27e+02 | 9.62e+01 | 8.52e+01 | 7.64e+01 |
| Noise | 5.54e+02 | 5.34e+02 | 5.00e+02 | 4.52e+02 | 3.87e+02 | 3.10e+02 | 2.35e+02 | 1.75e+02 | 1.32e+02 | 1.01e+02 | 8.85e+01 | 7.86e+01 | |
| FGSM | 2.89e+02 | 2.79e+02 | 2.63e+02 | 2.39e+02 | 2.10e+02 | 1.79e+02 | 1.52e+02 | 1.28e+02 | 1.09e+02 | 9.21e+01 | 8.48e+01 | 7.82e+01 | |
| PGD | 2.77e+02 | 2.67e+02 | 2.51e+02 | 2.29e+02 | 2.02e+02 | 1.74e+02 | 1.48e+02 | 1.26e+02 | 1.07e+02 | 9.18e+01 | 8.48e+01 | 7.85e+01 | |
| AAE-Cox | 1.66e+02 | 1.57e+02 | 1.48e+02 | 1.38e+02 | 1.27e+02 | 1.17e+02 | 1.08e+02 | 9.94e+01 | 9.16e+01 | 8.44e+01 | 8.10e+01 | 7.76e+01 | |
| SAWAR | 2.11e+02 | 1.92e+02 | 1.73e+02 | 1.57e+02 | 1.42e+02 | 1.28e+02 | 1.15e+02 | 1.04e+02 | 9.42e+01 | 8.52e+01 | 8.11e+01 | 7.72e+01 |
Appendix G Friedman Hypothesis Test
Instead of relying on confidence intervals—which do not definitively determine whether one adversarial training method outperforms another—we conducted a Friedman hypothesis test. While it is common in the machine learning and AI research community to repeatedly perform pairwise hypothesis tests based on confidence intervals (e.g., repeatedly for each dataset comparing two adversarial methods performance), this approach has significant drawbacks. It increases the risk of a high false discovery rate, potentially leading to incorrect rejection of the null hypothesis, and often relies on the normality assumption, which may not hold in practice. Therefore, we conducted a Friedman hypothesis test at a significance level of 0.05 for each metric (CI, IBS, and negative log-likelihood), treating the adversarial training methods as the “treatments” and the combinations of datasets, perturbation strengths, and perturbation methods as the “blocks”. This test was chosen because it is well-suited for repeated measurements and allowed us to assess whether our training method produces a more robust, better-fitted, and more calibrated model across various perturbations (perturbation method and perturbation strength). Since the Friedman test revealed statistically significant differences for each metric, we conducted a post-hoc Conover-Iman test (a rank-based approach) to analyze pairwise differences between groups. To control the false discovery rate, we applied the Benjamini-Hochberg procedure for p-value adjustment. Our findings show that our method, SAWAR, is indeed statistical significant improvement in performance (see critical difference diagram in Fig. 5).